### v1

# PRISM: Proof-Carrying Artifact Generation through LLM $\times$ MDE Synergy and Stratified Constraints

Abstract.
PRISM unifies Large Language Models with Model-Driven Engineering to generate regulator-ready artifacts and machine-checkable evidence for safety- and compliance-critical domains.
PRISM integrates three pillars: a Unified Meta-Model (UMM) reconciles heterogeneous schemas and regulatory text into a single semantic space; an Integrated Constraint Model (ICM) compiles structural and semantic requirements into enforcement artifacts including generation-time automata (GBNF, DFA) and post-generation validators (e.g., SHACL, SMT); and Constraint-Guided Verifiable Generation (CVG) applies these through two-layer enforcement—structural constraints drive prefix-safe decoding while semantic/logical validation produces machine-checkable certificates. When violations occur, PRISM performs audit-guided repair and records generation traces for compliance review.
We evaluate PRISM in automotive software engineering (AUTOSAR) and cross-border legal jurisdiction (Brussels I bis). PRISM produces structurally valid, auditable artifacts that integrate with existing tooling and substantially reduce manual remediation effort, providing a practical path toward automated artifact generation with built-in assurance.
model-driven engineering; meta-model integration; constraint-guided generation; large language models; formal verification; traceable artifacts
1. Introduction
1.1. From PIM–PSM Transformations to Unified Meta-Models
Model-driven architecture (MDA) advocated raising design effort from code to high-level platform-independent models (PIMs), then systematically refining them into platform-specific models (PSMs) and code (Object Management Group, 2003; Brambilla et al., 2017; Mardani Korani et al., 2023). Recent advances in meta-model integration aim to provide unified views across heterogeneous domain descriptions by merging multiple domain-specific languages (DSLs) into a single unified meta-model (UMM) (Amaral et al., 2023). However, substantial human curation is still required to align concepts and avoid semantic drift during integration, making the “unification power of models” an active research challenge rather than a solved problem.
1.2. Conformance and Assurance in Safety-Critical Domains
Safety-critical domains demand rigorous conformance to normative standards and bidirectional traceability across the entire development lifecycle (de la Vara et al., 2016). In automotive and avionics practice, models such as AUTOSAR (AUTomotive Open System ARchitecture) XML (ARXML) or the Architecture Analysis & Design Language (AADL) must satisfy prescribed structural and semantic rules; certification bodies additionally require that every requirement, design element, code unit, and test case be traceable in both directions (de la Vara et al., 2016).
Despite mature certification processes, a persistent verification gap remains between model-level properties and generated artifacts: most industrial generators still do not emit machine-checkable correctness certificates. Formal verification—including techniques encouraged for Design Assurance Level (DAL) A/B avionics software via DO-333—is increasingly applied to prove the absence of runtime errors or to discharge temporal-logic obligations. Qualified code generators such as Simulink/Stateflow and SCADE seek to preserve verified properties when emitting production code (Leroy, 2009). Together, these practices accumulate objective evidence for certification authorities, but they rely on a toolchain that is expensive to qualify and difficult to evolve.
1.3. Pain Points of Conventional MDE Pipelines
Industrial model-driven engineering (MDE) pipelines expose three long-standing limitations.
(i) Evolution overhead. Large, chained transformations are brittle: schema updates or new domain rules induce schema drift, invalidating hard-coded mappings and templates and forcing costly rework.
(ii) Lack of correctness-by-construction. Unlike verified research compilers such as CompCert (Leroy, 2009), which deliver machine-checkable proofs alongside the generated object code, most model compilers emit artifacts with no accompanying certificate. As a result, correctness must be established post hoc through manual review or downstream analysis (Mugnier et al., 2025).
(iii) Missing solver-based self-repair. When a generated artifact violates a constraint, engineers typically diagnose and repair the issue by hand. Automated, constraint-guided repair that spans structural, semantic, and logical layers remains largely confined to research prototypes (Mündler et al., 2025; HouXinyi et al., 2024). These gaps hinder continuous compliance and force recurring review cycles to re-establish conformance after each change.
1.4. LLMs and Knowledge-Guided Generation: A New Opportunity
Large Language Models (LLMs) such as Codex and GPT, together with open-source peers including StarCoder and DeepSeek-Coder, can already generate compilable multi-file projects directly from natural-language prompts, albeit with varying degrees of structural and semantic correctness (Liu et al., 2024; Idrisov and Schlippe, 2024; Wong et al., 2023). Recent work has begun exploring the synergy between LLMs and Model-Driven Engineering: Lebioda et al. (Petrovic et al., 2024) demonstrate LLM-assisted automotive software development through Ecore model instance creation followed by OCL constraint validation, while Alaoui Mdaghri et al. (Alaoui Mdaghri et al., 2025) propose leveraging LLMs for DSL modeling from natural language descriptions within an iterative validation framework. Similarly, Patil et al. (Patil et al., 2025) show that specification-driven LLM code generation combined with formal verification can produce safety-critical embedded automotive software even without iterative backprompting (Patil et al., 2025). However, unconstrained sampling frequently violates syntax rules or domain-specific constraints.
Grammar- and prefix-constrained decoding mitigates these failures by intersecting the model’s token distribution with a formal grammar, ensuring that every incremental prefix remains well formed while incurring negligible loss in fluency. This has been demonstrated by incremental and semantically constrained approaches such as PICARD and Synchromesh (Scholak et al., 2021; Poesia et al., 2021), and further optimized for subword alignment and efficiency by DOMINO (Beurer-Kellner et al., 2024; Geng et al., 2024).
Beyond syntax enforcement, knowledge-graph-augmented prompting injects authoritative domain triples—for example, AUTOSAR component hierarchies—into the LLM context, grounding generation and reducing hallucinations (Zhang et al., 2024). Finally, iterative LLM+SMT loops are emerging: an LLM proposes an artifact, an SMT solver checks constraints, counterexamples are fed back, and the model (or an auxiliary synthesizer) repairs defects until the artifact satisfies all checks (Tang et al., 2024). Early studies report dramatic reductions in constraint violations, pointing toward pipelines in which artifacts are generated under explicit, checkable constraints—a capability largely missing from classical MDE practice.
2. Related Work
2.1. Model Transformations in MDA
Model-Driven Architecture (MDA) employs model-to-model (M2M) transformations (e.g., ATL, QVT, Triple Graph Grammars (TGG)) and model-to-text (M2T) transformations (e.g., Acceleo, Xtend) to automate structured artifact derivation (Ma et al., 2025a; Kulkarni et al., 2023; Kahani et al., 2019). These transformation languages provide explicit, systematic mappings between abstraction levels, in contrast to ad hoc, manually curated integration efforts that are vulnerable to semantic drift (see Section 1). However, while they improve repeatability, they typically do not offer the kind of explicit semantic preservation guarantees that a colimit-based unified meta-model construction can provide.
2.2. Specification and Certification in Regulated Domains
Safety-critical sectors require that every development artifact conform to normative schemas and be accompanied by auditable evidence. Standards such as DO-178C and ISO 26262 mandate rigorous conformance and full traceability, and domain-specific coding and safety guidelines—including MISRA-C (Motor Industry Software Reliability Association C guidelines) and SOTIF (Safety Of The Intended Functionality)—require structured assurance cases. Tool qualification (e.g., qualification of Simulink auto-coders), bidirectional trace links, and safety arguments structured using Goal Structuring Notation (GSN) are routinely required; empirical studies document the effort needed to curate evidence that ties each safety claim to verified artifacts (El-Gnainy et al., 2024; Nair et al., 2014).
As a result, model-driven toolchains in automotive, avionics, and healthcare often embed schema validators and certification artifact generators to ensure that outputs such as AUTOSAR XML and AADL designs satisfy domain rules “by construction.” This reduces the manual burden at audit time but further entrenches specialized, domain-qualified tooling that is costly to evolve.
2.3. Formal Verification and Proof-Carrying Artifacts
Formal methods enrich modeling languages with machine-checkable semantics. Contracts in the form of assume/guarantee pairs and Object Constraint Language (OCL) invariants enable compositional reasoning; model checkers such as SPIN and NuSMV, as well as Satisfiability Modulo Theories (SMT)-based analyzers, are used to validate behavioral properties before code generation. Interactive proof assistants (e.g., Isabelle, Coq, PVS) remain essential wherever deep functional guarantees are required.
The proof-carrying code (PCC) paradigm attaches machine-verifiable correctness certificates to generated artifacts, providing a theoretical foundation for safety-critical practice, though complete PCC-style adoption in industrial pipelines is still limited. Solver-aided synthesis and repair techniques push this further: Max-SAT-based approaches such as DirectFix compute minimal patches that re-establish property satisfaction (Orvalho et al., 2024; Mechtaev et al., 2015).
Recent domain studies also demonstrate scalability (Charalambous et al., 2023). For instance, AUTOSAR XML models have been translated into timed-automata templates and exhaustively analyzed with UPPAAL for end-to-end latency verification (Zhang et al., 2023), and Architecture Analysis & Design Language (AADL) models enriched with a Safety Annex have undergone model-checking-driven fault injection to enumerate minimal cut sets (Stewart et al., 2021). The PRISM pipeline—which integrates the Unified Meta-Model (UMM), Integrated Constraint Model (ICM), and Constraint-Guided Verifiable Generation (CVG) stages—inherits this “lift-to-formal-semantics” philosophy while additionally attaching proof-carrying certificates to each generated artifact (Ma et al., 2025b).
2.4. LLM-Assisted Generation under Structural Constraints
Large Language Models (LLMs), including Codex and Claude, can already draft substantial portions of code, configuration files, and AUTOSAR-style XML directly from natural-language prompts. However, naive (unconstrained) sampling frequently violates grammar rules, schema cardinalities, naming conventions, and cross-reference requirements, rendering raw outputs unusable in safety-critical pipelines.
To address this, prior work explores grammar- and prefix-constrained decoding, in which decoding proceeds under a context-free grammar or a deterministic automaton so that every incremental prefix remains syntactically valid, as demonstrated by Synchromesh and PICARD (Scholak et al., 2021), with negligible fluency loss. An alternative strategy, grammar prompting (Wang et al., 2023b), guides the LLM to first synthesize a minimally sufficient grammar tailored to the specific generation task, and then to generate content under that task-specific grammar. This treats constraint specification as an LLM-interpretable guide rather than a hard automaton mask.
More recent decoding algorithms such as DOMINO enforce constraints at the subword level and combine speculative decoding with precomputation to achieve near-zero runtime overhead while still honoring a target grammar (Beurer-Kellner et al., 2024). Follow-up work on fast grammar-constrained decoding shows how to align grammar tokens with the model tokenizer and to build the automaton efficiently, reducing offline preprocessing cost by more than an order of magnitude while preserving efficient online masking (Park et al., 2025b).
In parallel, reinforcement-learning-style schema alignment aims to train LLMs to emit JSON that is provably valid with respect to large, real-world schemas. Nevertheless, even frontier models systematically fail on complex industrial schemas, which has motivated dedicated benchmarks such as SchemaBench and analyses such as JSONSchemaBench that characterize the trade-off between “over-constrained” engines (which reject semantically valid but grammatically atypical structures) and “under-constrained” engines (which accept ill-formed outputs and thus require repair) (Lu et al., 2025). Retrieval- and knowledge-graph-augmented prompting grounds generation in authoritative domain ontologies (for example, AUTOSAR component hierarchies), mitigating hallucinations. Finally, Ferrari and Spoletini show that LLMs can synthesize first-order temporal predicates and safety conditions directly from regulatory prose, effectively extracting normative logical constraints from natural-language standards (Ferrari and Spoletini, 2025).
Emerging industrial pipelines iterate on these trends: an LLM drafts an artifact; for structural layers, a supported subset of JSON Schema / regular expressions / Generalized Backus–Naur Form (GBNF) is compiled into deterministic automata to guide decoding via prefix-safe masking and online tracing; post-generation semantic and logic validators (e.g., SHACL, SMT) then check semantic and logical constraints, capture machine-checkable evidence, and drive automated repair. PRISM generalizes this pattern into a unified, auditable “verify-as-you-generate” loop.
A critical challenge identified in constrained decoding research is distribution distortion: while grammar masking guarantees structural validity, it can bias LLM generation toward the shortest valid completions, yielding artifacts that are syntactically correct but semantically impoverished (Park et al., 2025a). This effect arises because hard constraint masking eliminates longer yet equally valid continuations, thereby distorting the LLM’s learned probability distribution.
Grammar-Aligned Decoding (GAD) (Park et al., 2025a) addresses this through adaptive sampling with approximate expected futures (ASAp): rather than deterministically masking all invalid tokens, GAD performs limited forward simulation to estimate which candidate tokens can still lead to a grammatical completion, and samples from a distribution that balances grammaticality with fidelity to the base LLM’s preferences. Empirical results on code generation and structured data extraction show that GAD produces outputs with higher likelihood under the base model while maintaining grammatical correctness.
However, GAD’s forward simulation introduces computational overhead that scales with grammar complexity and generation length. For high-assurance domains that generate large-scale artifacts (e.g., thousands of lines) under deeply nested industrial grammars, this overhead can be prohibitive. Moreover, sectors with strict safety-critical requirements may prioritize deterministic structural guarantees over probabilistic grammaticality assurances, because even low-probability violations can have unacceptable consequences. These domain-specific constraints motivate architectures that stratify enforcement across multiple layers with different computational and assurance characteristics, as explored in Section 3.
3. Methodology: The UMM–ICM–CVG Framework
3.1. Overview
Figure 1. PRISM links domain requirements to verifiable generation. Natural-language requirements and existing meta-models are consolidated into a Unified Meta-Model (UMM) and an Integrated Constraint Model (ICM). Generation is guided by structural constraints (Layer-1, enforced during decoding) and then checked by semantic / logical validators (Layer-2) after decoding. The validation signals, together with the decode-time audit record, form evidence that can be used to automatically repair the artifact in a closed loop.:
Motivation and Design Goals.
Industrial, safety-critical domains (e.g., AUTOSAR-compliant automotive software, legal compliance workflows) require not only functionally plausible artifacts but also verifiably correct artifacts that can be audited by certification authorities (Yang et al., 2024; Wang et al., 2023a). Unconstrained Large Language Models (LLMs) excel at understanding domain intent and producing structured drafts, but they do not by themselves guarantee regulatory conformance, global consistency, or traceable justification. In contrast, Model-Driven Engineering (MDE) offers rigor through meta-models, transformation rules, and formal analysis, yet is brittle to evolving standards and costly to maintain as domains change. PRISM addresses this gap as an end-to-end high-assurance generation pipeline that couples the flexibility of LLMs with the accountability of formal methods. The pipeline is organized around three tightly connected components: (1) the Unified Meta-Model (UMM), which provides an auditable semantic backbone across heterogeneous domain sources; (2) the Integrated Constraint Model (ICM), which aggregates structural, semantic, and logical constraints; and (3) the Constraint-guided Verifiable Generation (CVG) process, which enforces constraints during generation, certifies artifacts after generation, and drives targeted repair when violations are found. Together, these components allow PRISM to automate the 80% of routine artifact construction while still producing machine-checkable evidence for expert review on the remaining 20%.
Core Components.
PRISM coordinates LLMs and formal methods in a relay pattern: the UMM and ICM formalize what “correct” means, and CVG—which includes Audit-Guided Repair (AGR)—uses those definitions to control generation, verify conformance, and close the loop with targeted repairs when inconsistencies are detected. We summarize each component below.
Unified Meta-Model (UMM). The UMM is not a passive schema repository; it is the semantic contract that both constrains generation and explains to auditors what each generated element is supposed to mean. To build the UMM, PRISM merges heterogeneous meta-models and domain ontologies (e.g., AUTOSAR XSD, legacy UML variants) into a coherent, typed graph $M_{\text{UMM}}=(V,E,\mathcal{C}_{\text{struct}})$, where nodes denote domain entities (classes, attributes) and edges record structural relations. This merge is carried out using a mathematically well-defined unification procedure (instantiated in our implementation as a category-theoretic colimit construction (Diskin and Maibaum, 2012)), which ensures that overlapping concepts from different sources are consistently identified and traceable. The point is not that the colimit construction itself is a contribution, but that it lets us claim semantic coherence and auditability: every node in the UMM can be traced back to its source models. For generation, PRISM materializes relevant subgraphs of the UMM as retrieval-augmented context—e.g., JSON-style schema fragments and textual descriptions—which are injected into LLM prompts to give the model a precise structural and conceptual scaffold.
Integrated Constraint Model (ICM). The ICM collects, normalizes, and compiles constraints that the generated artifact must satisfy. It is populated via two coordinated channels that mirror the strengths of MDE and LLMs:
  • Channel 1 (Deductive, schema-driven). Traditional model transformations extract explicit invariants from domain meta-models: field cardinalities, type hierarchies, mandatory references, and other structural rules. These invariants are organized into a multi-sorted family $(\mathcal{C}_{\text{lex}},\mathcal{C}_{\text{struct}},\mathcal{C}_{\text{sem}},\mathcal{C}_{\text{logic}})$ and compiled into executable artifacts at two layers: Layer-1 (L1) guards are realized as JSON Schema / Regex / GBNF constraints and determinized finite-state controllers for generation-time enforcement; Layer-2 (L2) validators are realized through semantic checkers (e.g., SHACL shapes for graph constraints) and logic checkers (e.g., SMT solvers for numeric/temporal properties) for post-generation validation. Formal analysis provides completeness guarantees: Theorem 8 states that, under bounded unfolding depth $d$, satisfying the compiled constraints implies satisfying the original meta-model invariants.
  • Channel 2 (Inductive, LLM-assisted). Many safety or regulatory requirements in practice are not fully captured by XSDs or UML—for example, “an OperationInvokedEvent must reference an existing operation,” or “mode transitions must form a directed acyclic graph.” Channel 2 uses an LLM extractor $E_{\theta}$ to read natural-language specifications (e.g., AUTOSAR Software Component Template PDFs) that have been augmented with inlined UMM definitions (see Figure 3). $E_{\theta}$ proposes candidate L2-only constraints (expressed as semantic or logic assertions). Before these constraints are admitted into the ICM, PRISM links each predicate back to UMM entities via entity linking (Definition 4) and checks their semantic compatibility (Definition 6), then runs consistency check to reject contradictory rules (Algorithm 2).
This dual-channel design allows PRISM to ingest both structured sources (Channel 1) and free-text specifications (Channel 2) without sacrificing rigor: LLM-proposed rules do not bypass formal scrutiny, but instead become auditable constraints that are grounded in—and indexed by—the UMM.
Constraint-guided Verifiable Generation (CVG). CVG is the execution core of PRISM. It governs how artifacts are produced, checked, and—if necessary—repaired. It proceeds in three stages:
  1. Stage 1:
    Generation with real-time constraint enforcement. Layer-1 constraints are compiled into a prefix-closed automaton (deterministic finite automaton for regular structures or bounded pushdown automaton for context-free patterns, unfolded to depth $d$; see Section 3.4). During decoding, at each step $t$, the executor computes the allowed token set $M_{t}=\{y\in\Sigma:\delta(s_{t},y)\neq\bot\}$, where $s_{t}$ is the current automaton state and $\delta$ is the transition function. Invalid continuations are masked before the LLM samples the next token. This guarantees prefix safety: every partial output always remains on a path that could complete to a structurally valid artifact (Theorem 12). In parallel, PRISM records an audit trail $\tau_{t}=\langle s_{t},M_{t},y_{t},\delta(s_{t},y_{t}),\Delta t\rangle$, which captures the control decisions that led to each generated token for later auditing.
  2. Stage 2:
    Post-generation formal verification. After the candidate artifact is produced, PRISM runs Layer-2 validators. Semantic validators check constraints such as cross-file reference integrity and role consistency; logic validators check temporal and numeric properties such as ordering constraints and resource bounds. Each validator emits a machine-checkable certificate $\pi_{\bullet}\in\{\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}}\}$ (e.g., DFA acceptance traces, semantic validation reports, logic proofs/unsat cores). Crucially, these certificates are independent of the generator: an external auditor can re-check correctness using only the artifact $a$, the evidence bundle $\Pi$, and the validator identifiers—without having to trust the LLM itself (Definition 13).
  3. Stage 3:
    Audit-Guided Repair (AGR). If validation fails, PRISM does not blindly re-generate from scratch. Instead, AGR analyzes the composite evidence $\Pi=(\pi_{\text{struct}}\otimes\pi_{\text{sem}}\otimes\pi_{\text{logic}})\oplus\tau$ to identify the root cause. Structural violations are localized using the DFA state in $\tau$ (e.g., which mandatory element was skipped); semantic violations are localized using violation paths from semantic validators (e.g., which reference is dangling); logical violations are localized using unsat cores from logic solvers (e.g., which constraints are mutually inconsistent) (Algorithm 4). AGR then invokes the LLM in a constrained repair mode: the prompt includes the violation diagnostics plus the relevant UMM/ICM context, guiding the model to propose targeted fixes rather than producing an unrelated new artifact.
System Flow and Human Oversight.
Figure 1 summarizes how these components interact. Natural-language requirements are first interpreted by an LLM-based NLP Processing Layer, while explicit domain schemas are parsed by a Knowledge-Graph Construction pipeline that builds the UMM and populates the ICM through Channel 1 and Channel 2. The Unified Constraint Layer then compiles these constraints into L1 automata and L2 validators. During Instance Model Generation, the LLM receives retrieval-augmented UMM subgraphs as context, and the Structure Decoder enforces L1 constraints through token masking while logging audit tuples $\tau_{t}$ to an Evidence Registry. After generation, the Validation & Repair Layer executes L2 validators (structural $\rightarrow$ semantic $\rightarrow$ logical), appends each certificate $\pi_{\bullet}$ to the registry, and either: (i) assembles the final verifiable artifact $\mathfrak{A}=(a,\Pi,\varphi)$ by hashing the artifact, the evidence, and a timestamp $H(a\|\Pi\|t)$; or (ii) triggers AGR to request a focused repair from the LLM or, in edge cases, a manual override by a domain expert. This “graduated automation” design lets PRISM automate routine generation while keeping human reviewers in control of exceptional cases, now equipped with machine-checkable diagnostics instead of ad hoc debugging.
Roadmap for the Remainder of the Section.
The following subsections elaborate each component. Section 3.2 details the construction of the UMM and explains how heterogeneous domain specifications are merged into a single auditable model. Section 3.3 describes how the ICM unifies deductive constraints (Channel 1) and LLM-assisted constraints (Channel 2) under the UMM, including entity linking and consistency checks. Section 3.4 presents CVG, including automaton-guided decoding (L1) and semantic/logic validation (L2), and proves prefix safety under bounded unfolding. Section 3.5 defines the composite evidence $\Pi$ and explains how PRISM packages artifacts together with machine-checkable certificates. Section 3.6 introduces the audit trail recorder, the evidence registry, and the AGR loop, and analyzes repair convergence in terms of layered violation dynamics.
3.2. Unified Meta-Model Construction: A Dual-Path Approach
Problem Formulation and Goals.
In practice, safety-critical domains present two very different starting points for model construction. Some domains (e.g., AUTOSAR, AADL, OPC UA) already publish an explicit meta-model: classes, attributes, and relationships are described in machine-readable form (XSD/XMI/Ecore) or in tightly maintained technical documentation. Other domains (e.g., regulatory and legal workflows, device compliance manuals) do not provide such an explicit schema, even though they implicitly assume a fairly well-structured universe of entities (“who is responsible,” “what object is configured,” “what reference must exist”). PRISM introduces the Unified Meta-Model (UMM) as the common semantic backbone across both cases. When an explicit meta-model exists, UMM ingests and normalizes it; when it does not, UMM inductively assembles the same kind of typed entity–relation structure from natural-language specifications with LLM assistance. In both cases, the result is a single, provenance-aware representation of the domain’s core entities and relationships that downstream components can treat as the source of truth. Its construction must satisfy three goals: (i) reconcile heterogeneous sources into one typed graph, (ii) maintain provenance so each element in the UMM can be traced back to its origin, and (iii) expose a stable structure that downstream components can rely on for generation, validation, and repair.
Graph-Based Representation.
Each source meta-model is represented as a typed directed graph.
Definition 3.0 (Meta-Model as Typed Graph).
A meta-model is represented as a typed directed graph $M=(V,E,T,\mathcal{C})$, where:
  • $V$: set of nodes representing domain entities (e.g., classes, attributes),
  • $E\subseteq V\times V$: set of edges representing relationships (e.g., associations, generalizations),
  • $T:V\cup E\rightarrow\mathcal{T}$: type mapping function assigning semantic types from a type system $\mathcal{T}$,
  • $\mathcal{C}$: set of structural constraints (cardinalities, mandatory fields, inheritance restrictions).
Given a collection of such graphs $\{M_{i}=(V_{i},E_{i},T_{i},\mathcal{C}_{i})\}_{i\in I}$, the Unified Meta-Model $M_{\text{UMM}}$ is obtained by merging them into a single typed knowledge structure. In our implementation, this merge follows a colimit-style typed graph unification in the sense of (Diskin and Maibaum, 2012): overlapping entities and relations across different $M_{i}$ are identified and merged through structure-preserving homomorphisms. This provides two desirable properties for PRISM: (1) semantic coherence—concepts that represent the same domain notion across sources are aligned rather than duplicated; and (2) auditability—every node and edge in $M_{\text{UMM}}$ retains provenance links to its source model(s), allowing downstream validators and human reviewers to trace “where this concept came from.”
Figure 2. Two ways to build the Unified Meta-Model (UMM). Path S1 handles structured sources: when a machine-readable meta-model (for example AUTOSAR, AADL, OPC UA) already exists, PRISM ingests and normalizes it. Path S2 handles unstructured sources: when only natural-language requirements are available, PRISM uses an LLM to induce domain entities and relations directly from text. Deployments typically follow one of these two paths, depending on which input is available.:
Dual-Path Construction Strategy.
PRISM supports two alternative construction paths for the UMM (Figure 2). In practice, deployments usually select one of these paths: either the domain already ships a machine-readable meta-model (Path S1), or it only has natural-language specifications with no canonical XSD/UML (Path S2). Both paths produce the same kind of output: a typed graph $M_{\text{UMM}}$ with provenance annotations. We describe the two paths below.
Path S1: Structured Meta-Model Transformation.
When an explicit domain meta-model exists, PRISM reuses an established dual-stage transformation process (Ma et al., 2025a). This process proceeds as follows:
  1. (1)
    Structural Extraction. XMI files are parsed to recover inheritance DAGs via topological sorting. XSD schemas are parsed to recover static data constraints such as field cardinalities and type restrictions. The result is a set of class/attribute definitions and their structural relationships.
  2. (2)
    Semantic Fusion. DP-Fusion (Dual-Path Fusion) merges the class hierarchy derived from XMI with the data-level constraints extracted from XSD. Conflicting field definitions are resolved by priority rules (e.g., XSD constraints override inherited defaults), yielding a single consistent view of entities, attributes, and associations.
  3. (3)
    Behavioral Annotation. Domain annotations (including stereotypes and association semantics) are carried forward as behavioral hooks: for example, reference-lookup patterns or CRUD-style method signatures that indicate how instances of a class are expected to interact.
Path S1 therefore produces a UMM instance in which: (i) entities, attributes, and relations reflect the authoritative schema; and (ii) selected behavioral expectations are explicitly attached to those entities. Within PRISM, this serves as the canonical semantic backbone whenever the domain already provides a formalized meta-model.
Path S2: LLM-Assisted Induction from Unstructured Sources.
Some domains lack any clean XSD/XMI-style schema. Instead, what counts as a “component,” a “reference,” or an “obligation” is only defined informally in prose. Path S2 lifts those informal descriptions into the same UMM representation through a three-stage pipeline:
Stage 1: Context-Enriched Document Segmentation. A corpus $D=\{d_{1},\ldots,d_{n}\}$ of natural-language specifications is segmented by a learned boundary detector $B$, which identifies sections likely to define entities, attributes, or relationships. For each segment $s_{i}$, PRISM injects domain context into the prompt (e.g., known entity names, typical field names, role labels from prior segments or from a seed ontology). This grounding step constrains the extractor’s vocabulary and reduces hallucinated concepts.
Stage 2: LLM-Based Entity and Relation Induction. An LLM extractor $E_{\theta}$, parameterized by domain knowledge $\theta$, maps each context-enriched segment to a bundle of candidate model elements:
(1) $E_{\theta}:\text{Segment}\rightarrow\mathcal{C}_{\text{raw}}\times[0,1],$
where $\mathcal{C}_{\text{raw}}$ is a set of proposed entities, attributes, and relations, and the confidence score in $[0,1]$ is used to discard low-certainty candidates. All candidates follow a fixed induction schema that records: a proposed identifier (name, ID), its intended role or parent entity, its described properties or fields in natural language, and provenance metadata (document reference, section index, confidence).
Stage 3: Consolidation into the UMM. The induced candidates are reconciled against the evolving $M_{\text{UMM}}$. We apply fuzzy alignment to associate each candidate with an existing UMM node when possible:
(2) $\alpha:\mathcal{C}_{\text{raw}}\rightharpoonup V_{\text{UMM}},$
where $\alpha(c)=\bot$ indicates that $c$ does not match any known node. If $\alpha(c)\neq\bot$, the candidate is treated as an alias or refinement of the matched node, and its provenance metadata is attached to that node. If $\alpha(c)=\bot$, PRISM tentatively adds a new node or relation into $M_{\text{UMM}}$, tagged with provenance and marked for optional human review when the stakes are high (e.g., safety-critical components). Crucially, at this stage we are only inducing and reconciling domain entities, attributes, and relations; constraint extraction and formal consistency checking are handled later by the Integrated Constraint Model (ICM) in Section 3.3. No SHACL or SMT reasoning is applied here.
Path S2 thus yields a UMM even in domains with no formal schema at all. The result is again a typed, provenance-aware graph $M_{\text{UMM}}$ that defines what “objects” exist, how they relate, and which textual sources justified their inclusion.
Scalability Note.
Although UMM construction is primarily an integration activity rather than a new theoretical contribution, its runtime characteristics matter for practical deployment. Let $n_{i}=|V_{i}|$ be the number of entities extracted from source $M_{i}$, and let $m$ be the number of detected overlaps across all sources. The reconciliation procedure maintains disjoint sets of candidate-equivalent nodes and performs unions as they are aligned. Its complexity is summarized below.
Lemma 2 (Construction Complexity).
The dual-path construction algorithm (including node alignment and union over overlaps) runs in
(3) $\mathcal{O}\!\biggl(\textstyle\sum_{i}n_{i}+m\,\alpha\!\bigl(\sum_{i}n_{i}\bigr)\biggr)$
time and
(4) $\mathcal{O}\!\bigl(\textstyle\sum_{i}n_{i}\bigr)$
space, where $\alpha$ is the inverse Ackermann function from disjoint-set analysis.
Sketch.
Candidate overlaps are proposed via bounded subgraph matching to identify entities that are likely semantically identical. These candidates are merged using union-find with path compression. Union-find yields near-constant amortized merge cost, which gives the stated near-linear bounds in practice. ∎
The near-linear scaling implied by Lemma 2 allows PRISM to ingest industrial meta-models with thousands of entities without incurring the $O(n^{2})$ blow-up typical of manual pairwise weaving.
Relation to Existing Approaches.
Classical model weaving frameworks (e.g., AMW) rely on manually authored correspondence models, while graph transformation systems (e.g., AGG) require hand-written rules for every pair of source models. Purely LLM-driven extraction has also been explored (Chen et al., 2023), but without grounding the induced concepts in a persistent, provenance-tracked ontology, such approaches risk hallucinated entities and silently inconsistent references. PRISM differs in three respects:
  • Source Adaptability. Path S1 ingests explicit XSD/XMI-style meta-models through deterministic model transformations (Ma et al., 2025a). Path S2 induces a meta-model from prose using an LLM guided by domain context. Deployers choose whichever path matches their domain reality.
  • Semantic Preservation. In both paths, newly discovered or merged entities are either aligned to an existing node in $M_{\text{UMM}}$ via $\alpha$ or added as a new node annotated with provenance and flagged for optional review. This prevents semantic drift and keeps the UMM auditable.
  • Scalability. Because reconciliation proceeds via incremental union rather than exhaustive pairwise weaving, PRISM scales to realistic industrial domains, as summarized in Lemma 2.
In summary, UMM construction is the step that turns messy, heterogeneous domain knowledge into a single, provenance-aware semantic backbone. All later phases of PRISM—constraint modeling, generation-time enforcement, post-generation validation, and audit-guided repair—assume this backbone as the shared point of truth.
3.3. Constraint Extraction and Integrated Constraint Model (ICM)
Role of the ICM.
The Unified Meta-Model (UMM) provides PRISM with a single, auditable semantic backbone. The Integrated Constraint Model (ICM) builds on that backbone by collecting, normalizing, and compiling the domain rules that generated artifacts must satisfy. These rules range from low-level structural requirements (e.g., mandatory fields, cardinalities) to high-level semantic and logical invariants (e.g., “an OperationInvokedEvent must reference an existing operation”), and they originate from both machine-readable meta-models and natural-language specifications. This subsection formalizes how such constraints are represented, how they are extracted through two complementary channels, how they are grounded in the UMM, and how they are compiled into executable Layer-1 (L1) and Layer-2 (L2) validators for generation-time enforcement and post-generation verification.
Formalization of Constraint Spaces.
To reason uniformly about constraints from heterogeneous sources, we model them in a typed semantic space.
Definition 3.0 (Constraint Space).
A constraint space $\mathcal{C}=(\mathcal{L},\mathcal{S},\models,\leq)$ consists of:
  • $\mathcal{L}$: a logical language for expressing constraints,
  • $\mathcal{S}$: a semantic domain of valid artifact structures,
  • $\models$: a satisfaction relation $\mathcal{S}\times\mathcal{L}\rightarrow\{\top,\bot\}$,
  • $\leq$: an ordering on $\mathcal{L}$ capturing relative constraint strength.
Different inputs to PRISM (e.g., AUTOSAR XSD/XMI schemas vs. regulatory PDFs) induce different constraint spaces. Before integration, these must be semantically aligned with the UMM so that downstream validators can interpret them consistently.
Definition 3.0 (Partial Alignment Mapping).
Let $\mathcal{P}$ be the set of extracted constraint patterns and let $U$ be the set of UMM sorts. A partial alignment is a (possibly partial) mapping
(5) $\alpha:\mathcal{P}\rightharpoonup U$
where $\alpha(p)=\bot$ denotes that pattern $p$ has no direct anchor in $U$ and must instead be retained at an abstract layer $\mathcal{C}_{\mathrm{abs}}$.
Intuitively, $\alpha$ tells us which extracted rules are already grounded in the UMM and which must be quarantined for review or additional interpretation.
Dual-Channel Constraint Acquisition.
ICM population proceeds through two coordinated channels, targeting persistent, reusable domain constraints. Both channels produce structured intermediate representations that are compiled into executable validators and stored with provenance. The channels are:
Channel 1: Deductive Extraction from Structured Meta-Models. Given a source meta-model $M=(V,E,T,\mathcal{C})$ as defined in Section 3.2, we apply programmatic operators that extract constraints directly from the schema:
$\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{struct}}$ $:M\to 2^{\mathcal{C}_{\mathrm{struct}}},$
$\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{sem}}$ $:M\to 2^{\mathcal{C}_{\mathrm{sem}}},$
(6) $\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{log}}$ $:M\to 2^{\mathcal{C}_{\mathrm{log}}}.$
Here, $\mathcal{C}_{\mathrm{struct}}$ captures structural invariants (cardinality, mandatory fields, inheritance restrictions), $\mathcal{C}_{\mathrm{sem}}$ captures semantic constraints (reference integrity, role consistency), and $\mathcal{C}_{\mathrm{log}}$ captures logical/temporal/numeric constraints.
L2 Validator Instantiation.
The abstract semantic and logic validators $\mathcal{V}_{\mathrm{sem}}$ and $\mathcal{V}_{\mathrm{log}}$ in $\mathcal{I}$ can be instantiated through various verification formalisms depending on domain requirements. In our implementation, we instantiate semantic validators using SHACL (Shapes Constraint Language) (Knublauch and Kontokostas, 2017) for graph-structural constraints such as reference integrity, cardinality restrictions, and type compatibility. SHACL naturally captures cross-reference semantics in XML/RDF-based artifacts like AUTOSAR ARXML configurations. For logic validators, we employ SMT (Satisfiability Modulo Theories) solvers (de Moura and Bjørner, 2008), which efficiently handle bounded numeric reasoning, temporal constraints, and resource allocation predicates common in automotive timing specifications and legal precedence rules.
This instantiation choice reflects the structural characteristics of our evaluation domains (AUTOSAR, Brussels I bis). The validator interface remains extensible: alternative formalisms such as Alloy for relational constraints, OWL reasoners for ontology-based validation, or domain-specific checkers can substitute or complement SHACL/SMT where appropriate. The key architectural requirement is that validators produce machine-checkable certificates $\pi_{\text{sem}}$ and $\pi_{\text{logic}}$ with violation localization metadata sufficient to drive Audit-Guided Repair (Section 3.6).
These extracted constraints are then compiled for enforcement at appropriate layers:
(7) $\mathrm{Compile}_{\mathrm{L1}}$ $:\mathcal{C}_{\mathrm{struct}}\to\{\text{DFA/GBNF, JSON Schema}\},$
$\mathrm{Compile}^{\mathrm{sem}}_{\mathrm{L2}}$ $:\mathcal{C}_{\mathrm{sem}}\to\mathcal{V}_{\mathrm{sem}},$
(8) $\mathrm{Compile}^{\mathrm{log}}_{\mathrm{L2}}$ $:\mathcal{C}_{\mathrm{log}}\to\mathcal{V}_{\mathrm{log}}.$
where $\mathcal{V}_{\mathrm{sem}}$ and $\mathcal{V}_{\mathrm{log}}$ denote the spaces of semantic and logic validators, respectively. In our implementation, these are instantiated as SHACL shapes and SMT-LIB2 formulas (Section 3.3). The result is that Channel 1 provides deterministic constraints with explicit provenance in the original schema, and these constraints become part of the persistent ICM.
Channel 2: Inductive Extraction from Natural-Language Specifications. Many industrial domains encode mission-critical rules in prose rather than structured schemas: AUTOSAR PDFs, regulatory guidance, safety standards, legal jurisdiction rules. Channel 2 uses an LLM-based extractor to convert free-form text into structured constraint candidates, following prior work on LLMs as structured information extractors (Dagdelen et al., 2024; Lu et al., 2022; Wei et al., 2020; Jiao et al., 2024). Formally,
$\mathrm{Extract}^{\mathrm{LLM}}_{\mathrm{sem}}$ $:\mathcal{D}\to 2^{\mathcal{C}_{\mathrm{sem}}^{\text{structured}}},$
(9) $\mathrm{Extract}^{\mathrm{LLM}}_{\mathrm{log}}$ $:\mathcal{D}\to 2^{\mathcal{C}_{\mathrm{log}}^{\text{structured}}}.$
These outputs are intermediate JSON-like specifications that undergo entity linking $\alpha$ (Definition 4), semantic compatibility validation (Definition 6), and compilation to L2 validators before admission to the ICM. This prevents unconstrained LLM extraction from directly polluting the validator set.
Figure 3. LLM-based extraction from natural-language specifications. The extractor proposes structured constraint candidates and aligns each candidate to entities in the UMM. Only candidates that align cleanly and pass semantic checks are committed to the Integrated Constraint Model (ICM); the rest are flagged for human review instead of being turned directly into validators.:
Layer Assignment: L1 vs. L2.
Channel 1 feeds $\mathcal{C}_{\mathrm{struct}}$ to L1, where constraints determinize to prefix-closed automata (DFA/GBNF, JSON Schema) for token-by-token masking during decoding. Channel 2 contributes only to L2, as free-form constraints may require global context unsuitable for incremental enforcement. Both channels populate $\mathcal{C}_{\mathrm{sem}}$ and $\mathcal{C}_{\mathrm{log}}$ in L2, compiled to semantic and logic validators applied post-generation. This separation preserves deterministic safety during generation while enabling rich semantic/logic checks afterward.
ICM as Persistent Knowledge.
We now formalize how both channels populate a persistent, auditable repository of constraints.
Figure 4. ICM as a multi-sorted algebra layered over the UMM. White nodes denote UMM entities; gray nodes encode semantic and logical constraints extracted from Channel 1 (deductive) and Channel 2 (LLM-assisted). Morphisms relate abstract constraints to class-/attribute-level constraints, enabling cross-layer reasoning and repair prioritization.:
Definition 3.0 (Integrated Constraint Model).
An ICM is a tuple $\mathcal{I}=(\mathcal{U},\mathcal{X},\mathcal{F},\Gamma)$ where:
  • $\mathcal{U}$: the underlying UMM providing sorts and signatures;
  • where $\mathcal{X}=\{\mathcal{C}_{\mathrm{struct}},\mathcal{C}_{\mathrm{sem}},\mathcal{C}_{\mathrm{log}}\}$: constraint algebras populated by Channel 1 and Channel 2;
  • $\mathcal{F}=\{f_{ij}:\mathcal{C}_{i}\rightarrow\mathcal{C}_{j}\}$: constraint morphisms supporting cross-layer reasoning;
  • $\Gamma$: a dependency lattice with partial order $\preceq$, which we later use to prioritize repair (Section 3.6).
The morphisms $\mathcal{F}$ capture how abstract invariants (e.g., “all mode transitions are acyclic”) induce obligations on concrete fields and references. The lattice $\Gamma$ then orders those obligations, which is later consumed by Audit-Guided Repair (AGR).
ICM Construction via Dual-Channel Extraction.
Algorithm 1 summarizes how PRISM assembles the ICM from both structured schemas and unstructured documents. This process is persistent: it populates the long-lived constraint repository (the ICM $\mathcal{I}$), rather than just synthesizing a one-off validator.
Algorithm 1 ICM Construction via Dual-Channel Extraction
1:Meta-model $M=(V_{M},E_{M},T_{M},\mathcal{C}_{M})$, UMM graph $(V_{\text{UMM}},E_{\text{UMM}})$, Spec docs $D=\{d_{1},\ldots,d_{n}\}$, Constraint schema $\mathcal{S}_{c}$
2:ICM graph $(V_{\text{ICM}},E_{\text{ICM}})$
3:function BuildICM($M$, $V_{\text{UMM}}$, $D$, $\mathcal{S}_{c}$)
4:  $V_{\text{ICM}}\leftarrow\emptyset$; $E_{\text{ICM}}\leftarrow\emptyset$
5:  for $\phi\in\mathcal{C}_{M}$ do $\triangleright$ Channel 1: Deductive extraction
6:    if $\phi\in\mathcal{C}_{\mathrm{sem}}\cup\mathcal{C}_{\mathrm{log}}$ then
7:     $c\leftarrow\text{createConstraintNode}(\phi)$
8:     add $c$ to $V_{\text{ICM}}$ and link to $V_{\text{UMM}}$ via $E_{\text{ICM}}$
9:    end if
10:  end for
11:  for doc $d\in D$ do $\triangleright$ Channel 2: LLM-assisted extraction
12:    $d^{\prime}\leftarrow\text{injectContext}(d,V_{\text{UMM}})$
13:    $\mathcal{C}_{\text{raw}}\leftarrow\text{extractStructured}(d^{\prime},V_{\text{UMM}},\mathcal{S}_{c})$
14:    for $c_{\text{raw}}\in\mathcal{C}_{\text{raw}}$ do
15:     $\alpha(c_{\text{raw}})\leftarrow\text{linkToUMM}(c_{\text{raw}},V_{\text{UMM}})$
16:     if $\alpha(c_{\text{raw}})=\bot$ then
17:      flag for human review; continue
18:     end if
19:     if validateSemanticCompatibility($c_{\text{raw}}$, $M$) then
20:      add $c_{\text{raw}}$ to $V_{\text{ICM}}$ with anchor $\alpha(c_{\text{raw}})$
21:     else
22:      repair_and_relink($c_{\text{raw}}$); repeat
23:     end if
24:    end for
25:  end for
26:  return $(V_{\text{ICM}},E_{\text{ICM}})$
27:end function
The central safeguard is that Channel 2 output does not automatically enter the persistent ICM $\mathcal{I}$ or any downstream validator. Each candidate constraint must (i) be anchored to a concrete UMM entity via $\alpha$, and (ii) pass semantic compatibility checks before being admitted.
LLM-Assisted Constraint Extraction with Entity Grounding.
Channel 1 recovers explicit constraints from structured schemas. Channel 2 is responsible for implicit constraints described only in prose, such as: “an OperationInvokedEvent must reference an existing operation” or “mode transitions must form a directed acyclic graph.” Algorithm 2 formalizes this pipeline.
Algorithm 2 Channel 2: LLM-Assisted Constraint Extraction for ICM
1:Document segment $s$, UMM entities $V_{\text{UMM}}$, constraint schema $\mathcal{S}_{c}$
2:Structured constraint candidates $\mathcal{C}^{\text{cand}}=\{c_{1}^{\text{cand}},\ldots,c_{n}^{\text{cand}}\}$
3:function ExtractStructuredConstraints($s$, $V_{\text{UMM}}$, $\mathcal{S}_{c}$)
4:  $\text{ctx}\leftarrow\text{serializeUMM}(V_{\text{UMM}})$
5:  $\mathcal{C}_{\text{raw}}\leftarrow\text{LLM}(\text{buildPrompt}(s,\text{ctx},\mathcal{S}_{c}),T{=}0.3)$
6:  $\mathcal{C}^{\text{cand}}\leftarrow\emptyset$
7:  for $c_{\text{raw}}\in\mathcal{C}_{\text{raw}}$ do
8:    $(v,\text{score})\leftarrow\text{fuzzyMatch}(c_{\text{raw}}.\text{target},V_{\text{UMM}})$
9:    if $v=\bot$ or $\text{score}<\theta_{\text{link}}$ then
10:     log_for_review($c_{\text{raw}}$, $v$, score); continue
11:    end if
12:    $c^{\text{cand}}\leftarrow\text{createStructuredForm}(c_{\text{raw}},v)$
13:    $c^{\text{cand}}.\text{type}\leftarrow\text{classifyConstraint}(c_{\text{raw}})$
14:    $c^{\text{cand}}.\text{anchor}\leftarrow v$
15:    $c^{\text{cand}}.\text{metadata}\leftarrow\langle\text{doc\_id},\text{para},c_{\text{raw}}.\text{conf},\text{score}\rangle$
16:    if validateSemanticCompatibility($c^{\text{cand}}$, $V_{\text{UMM}}$) then
17:     add $c^{\text{cand}}$ to $\mathcal{C}^{\text{cand}}$
18:    else
19:     log_for_review($c^{\text{cand}}$, reason=”incompatible”)
20:    end if
21:  end for
22:  return $\mathcal{C}^{\text{cand}}$
23:end function
The key mechanism is Step 2: entity grounding. Instead of accepting every LLM hypothesis, we force each proposed constraint to attach to an existing UMM node with confidence above threshold $\theta_{\text{link}}$. This (i) filters hallucinated entities and broken references, and (ii) creates a traceable link from each natural-language rule to a formal constraint and, through the UMM, to the certified domain ontology. Constraints that fail linking or compatibility are logged for expert review rather than silently admitted into the ICM.
Semantic Compatibility and Extraction Consistency.
To ensure that accepted constraints do not contradict the UMM, we require semantic compatibility.
Definition 3.0 (Semantic Compatibility).
Let $\mathcal{S}$ be the semantic domain of admissible artifacts from Definition 3, and let $\models$ be the satisfaction relation. A constraint $c$ is semantically compatible with a meta-model $M$ if and only if
(10) $\exists a\in\mathcal{S}:(a\models M)\land(a\models c).$
Intuitively, $c$ is compatible with $M$ if there exists at least one admissible artifact $a$ that can satisfy both $M$ and $c$ simultaneously, i.e., $c$ does not impose an impossible obligation relative to the domain described by $M$.
We now relate Channel 1 (schema-driven) and Channel 2 (LLM-assisted) extraction. For brevity, define:
$\mathrm{Extract}^{\mathrm{KG}}(M)\;\;=\;\;\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{struct}}(M)\;\cup\;\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{sem}}(M)\;\cup\;\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{log}}(M),$
and
$\mathrm{Extract}^{\mathrm{LLM}}(D)\;\;=\;\;\mathrm{Extract}^{\mathrm{LLM}}_{\mathrm{sem}}(D)\;\cup\;\mathrm{Extract}^{\mathrm{LLM}}_{\mathrm{log}}(D),$
where $D$ is the specification corpus.
Theorem 7 (Extraction Consistency with Partial Alignment).
For any valid specification corpus $D$ and its corresponding meta-model $M$, the extracted constraint sets satisfy
(11) $\bigl(\mathrm{Extract}^{\mathrm{KG}}(M)\cap\mathrm{Extract}^{\mathrm{LLM}}(D)\bigr)\;\cup\;\mathcal{C}_{\mathrm{abs}}\;\neq\;\varnothing,$
and every constraint $c\in\mathrm{Extract}^{\mathrm{LLM}}(D)$ obeys
(12) $\alpha(c)\neq\bot\;\Longrightarrow\;\bigl(\alpha(c)\in U\ \land\ c\text{ is semantically compatible with }M\bigr),$
where $\alpha$ is the partial alignment map (Definition 4), $U$ is the set of UMM sorts, and $\mathcal{C}_{\mathrm{abs}}$ is the quarantined abstract layer for rules that have no direct UMM anchor.
In words: either a Channel 2 rule anchors to a UMM entity with acceptable compatibility, or it is quarantined in $\mathcal{C}_{\mathrm{abs}}$ rather than being silently admitted into the ICM.
Proof sketch.
Let $\mathcal{T}:\mathcal{P}\rightarrow U\cup\{\bot\}$ be the pattern-to-model mapping. For any pattern $p\in D$ with $\mathcal{T}(p)=u\neq\bot$, $u\in U$ serves as an anchor establishing overlap with $\mathrm{Extract}_{\mathrm{ded}}(M)$. Patterns with $\mathcal{T}(p)=\bot$ are routed to $\mathcal{C}_{\mathrm{abs}}$ by Definition 4. Thus the combined set is non-empty, and compatibility follows from how $E_{\theta}$ is constructed to respect $M$. ∎
The theorem formalizes the discipline imposed by partial alignment $\alpha$: either a constraint grounds in the UMM and is provably compatible, or it is quarantined (in $\mathcal{C}_{\mathrm{abs}}$) for review instead of silently entering the validator pipeline.
Dynamic Constraint Synthesis (Runtime, Non-ICM).
In addition to the static ICM, PRISM supports request-specific constraints that apply only to a single generation run. For example, a user generating one AUTOSAR subsystem may require “all port connections in this subsystem must use compatible data types” without permanently altering the global repository. To accommodate such ad hoc requirements, we synthesize ephemeral L2 validators at runtime:
(13) $\mathrm{Synthesize}^{\mathrm{NLP}}:\mathcal{R}_{\text{dynamic}}\to\{\text{SHACL, SMT}\}.$
These validators are fed directly into L2 verification during CVG (Section 3.4), but they are not persisted in the ICM unless they pass solver-backed validation and are deemed reusable.
Algorithm 3 describes this guarded synthesis loop.
Algorithm 3 Dynamic NLP Constraint Synthesis (CVG Runtime)
1:Per-request NL requirement $r\in\mathcal{R}_{\text{dynamic}}$, UMM context $V_{\text{UMM}}$
2:Executable L2 validators $\mathcal{V}_{\text{L2}}=\{\mathcal{V}_{\text{sem}},\mathcal{V}_{\text{log}}\}$ (instantiated as SHACL and SMT)
3:function SynthesizeDynamicConstraints($r$, $V_{\text{UMM}}$)
4:  $\text{ctx}\leftarrow\text{selectRelevantEntities}(r,V_{\text{UMM}})$
5:  $\mathcal{C}_{\text{formal}}\leftarrow\text{LLM}(\text{buildSynthesisPrompt}(r,\text{ctx}),T{=}0.2)$
6:  $\mathcal{V}_{\text{SHACL}}\leftarrow\emptyset$; $\mathcal{V}_{\text{SMT}}\leftarrow\emptyset$
7:  for $c\in\mathcal{C}_{\text{formal}}$ do
8:    if $c.\text{type}=\text{``semantic''}$ then
9:     add $\text{compileToSHACL}(c,V_{\text{UMM}})$ to $\mathcal{V}_{\text{SHACL}}$
10:    else if $c.\text{type}=\text{``logic''}$ then
11:     add $\text{compileToSMT}(c)$ to $\mathcal{V}_{\text{SMT}}$
12:    end if
13:  end for
14:  $\text{conflicts}\leftarrow\text{detectConflicts}(\mathcal{V}_{\text{SHACL}}\cup\mathcal{V}_{\text{SMT}},r)$
15:  if $\text{conflicts}\neq\emptyset$ then
16:    resolve_via_relaxation(conflicts)
17:  end if
18:  return $\mathcal{V}_{\text{L2}}$
19:end function
Conceptually, Algorithms 1 and 2 populate the long-lived constraint repository (ICM) that applies across generation tasks. Algorithm 3 instead creates validators that exist only for the current request, unless they are later promoted (via solver-backed validation) into the persistent Formal Constraints Database. This policy prevents one-off requirements from polluting the global constraint store or contradicting established domain invariants.
Compilation to Executable Validators.
Finally, constraints admitted into the ICM must be turned into actual checkers used by CVG. Figure 5 illustrates PRISM’s stratified compilation pipeline.
Figure 5. Stratified compilation with LLM fallback and repository policy. Left: Channel 1 programmatic mapping compiles UMM/ICM constraints to L1 validators (JSON Schema, DFA, GBNF) and L2 validators (SHACL, SMT), preserving semantics. Right: when programmatic mapping is incomplete, an LLM-based pipeline (Ma et al., 2025b) synthesizes missing constraints with solver-backed validation; only verified outputs are persisted in the ICM storage.:
Following (Barrett et al., 2009; W3C RDF Data Shapes Working Group, 2017), we define compilation functions:
(14) $\kappa_{\mathrm{lex}}$ $:\mathcal{C}_{\mathrm{lex}}\rightarrow\mathrm{GBNF}$
(15) $\kappa_{\mathrm{struct}}$ $:\mathcal{C}_{\mathrm{struct}}\rightarrow\mathrm{DFA}_{\mathrm{prefix}}$
(16) $\kappa_{\mathrm{sem}}$ $:\mathcal{C}_{\mathrm{sem}}\rightarrow\mathcal{V}_{\mathrm{sem}}$
(17) $\kappa_{\mathrm{logic}}$ $:\mathcal{C}_{\mathrm{logic}}\rightarrow\mathcal{V}_{\mathrm{log}}$
Remark. In our implementation, context-free GBNF/CFG fragments are enforced at generation time using PDA/LR-style controllers. For bounded depth $d$, we apply finite unfolding to obtain an equivalent DFA for unified token masking and trace logging across structured decoding.
Theorem 8 (Conditional Soundness and Completeness).
Fix an unfolding bound $d$ for recursive structures. Let $\kappa_{\text{struct}}:\mathcal{C}_{\text{struct}}\!\to\!\text{DFA}$, $\kappa_{\text{sem}}:\mathcal{C}_{\text{sem}}\!\to\!\text{SHACL}$, $\kappa_{\text{logic}}:\mathcal{C}_{\text{logic}}\!\to\!\text{SMT}$, and write $\kappa$ for their union on $\mathcal{C}$ with cross-layer deduplication. Then:
  1. (i)
    Soundness. For all $c\in\mathcal{C}$ and $a\in\mathbb{A}$:  $a\models c\Rightarrow a\models\kappa(c)$.
  2. (ii)
    Conditional completeness. For all $c\in\mathcal{C}$ and $a\in\mathbb{A}$ with $\text{maxDepth}(a)\leq d$:  $a\models\kappa(c)\Rightarrow a\models c$.
Proof sketch.
(i) By construction: DFA acceptance implies satisfaction of the structural fragment; Semantic validators encode graph-structural constraints (e.g., SHACL shapes for first-order relations); logic validators capture temporal and numeric properties (e.g., SMT formulas for bounded arithmetic). Each $\kappa_{\bullet}$ is a semantics-preserving encoding.
(ii) For recursive structures, finite unfolding to depth $d$ yields an automaton whose accepted language over-approximates the target up to depth $d$. Any artifact $a$ with $\text{maxDepth}(a)\leq d$ that satisfies $\kappa(c)$ admits a witness (accepting run / model) reconstructing a proof that $a\models c$. Cross-layer deduplication ensures that constraints assigned to the most specific executable layer remain entailed in stronger layers. ∎
Together, these results guarantee that constraints admitted into the ICM can be enforced in two complementary ways: L1 constraints prevent structurally invalid outputs during generation, and L2 validators certify semantic and logical correctness after generation. The evidence produced by these validators is later combined with audit trails to drive Audit-Guided Repair (Section 3.6).
3.4. Unified Automaton Execution with Theoretical Guarantees
Role in CVG.
Constraint-Guided Verifiable Generation (CVG) enforces Layer-1 (L1) structural constraints during decoding and Layer-2 (L2) semantic/logic constraints after decoding (Section 3.3). This subsection explains how L1 constraints are compiled into a unified automaton that steers the Large Language Model (LLM) token-by-token, how this execution is audited, and why L1 and L2 must remain distinct from a computability perspective. We further show how these mechanisms provide prefix safety, bounded structural repair, and machine-checkable evidence for certification and Audit-Guided Repair (AGR) (Section 3.6).
Figure 6. Unified automaton execution for L1 enforcement during LLM decoding. Input constraints: L1 structural constraints ($\mathcal{C}_{\mathrm{struct}}$ from the ICM; Section 3.3) are compiled into executable automata. Finite-state fragments (JSON Schema / Regex / FSM) are determinized to prefix-closed DFAs. GBNF/CFG fragments are executed via PDA/LR; for unified masking and logging, bounded-depth unfolding (depth $d$) produces an equivalent DFA $\mathcal{A}_{d}$. Runtime decoding loop: At decoding step $t$, the executor computes $M_{t}=\{y\in\Sigma\mid\delta(s_{t},y)\neq\bot\}$, masks invalid tokens in the LLM logits, samples a valid token $y_{t}$, and transitions to $s_{t+1}=\delta(s_{t},y_{t})$. Audit trail: Each step records $\tau_{t}=\langle s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t\rangle$ into the Evidence Registry. Post-generation, L2 validators (semantic checkers for graph constraints and logic checkers for temporal/numeric properties) certify the completed artifact, yielding composite evidence $\Pi=(\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}})\oplus\tau$ (Section 3.5).:
Why Layering is Necessary: Decidability and Auditability.
Existing grammar-aligned decoding methods enforce a single global grammar or schema during LLM decoding and can align external grammars with subword vocabularies, using speculative decoding to keep runtime overhead negligible while preserving grammar validity (Scholak et al., 2021; Park et al., 2025a; Deljouyi, 2024; Beurer-Kellner et al., 2024; Park et al., 2025b; Lu et al., 2025). However, large-scale evaluations (e.g., SchemaBench / JSONSchemaBench) report two persistent failure modes: (i) over-constraining, where grammars block structurally unusual but semantically valid outputs, and (ii) under-constraining, where syntactically valid outputs still violate cross-reference integrity, safety rules, or timing constraints (Geng et al., 2025). The root cause is computability: not all constraints are prefix-decidable.
Layer-1 constraints (L1). Structural and lexical invariants—“element A must contain child B,” “attribute X must be an integer,” “balanced tags,” etc.—are local and prefix-decidable. Their satisfiability at step $t$ depends only on the partially generated artifact plus bounded lookahead. Such constraints can be compiled into finite automata (DFA/PDA) that, at each decoding step, expose the set of tokens that keep the output on a valid path. They are therefore enforceable during generation.
Layer-2 constraints (L2). Semantic and logic invariants—“every OperationInvokedEvent must reference a defined Operation,” “mode transitions must form an acyclic graph,” “all port connections must be type-compatible”—are global. Their satisfiability depends on full-artifact knowledge (cross-file symbol tables, temporal relations, numerical bounds). Deciding these properties while decoding is infeasible because it would require reasoning about content not yet generated. They must therefore be checked after generation using semantic and logic validators (Section 3.3).
This is not an engineering convenience but a theoretical split. L1 maximizes what can be proven safe online, and L2 certifies global semantics offline. The separation underpins PRISM’s claim of constraint-guided, auditable generation in safety-critical domains.
Definition 3.0 (Constraint Enforcement Hierarchy).
Let $\mathcal{H}$ be a family of enforcement strategies, ordered by $\preceq$. For $h_{i},h_{j}\in\mathcal{H}$ we write $h_{i}\preceq h_{j}$ iff: (i) $\mathcal{L}(h_{i})\supseteq\mathcal{L}(h_{j})$, i.e., $h_{i}$ permits a superset of sequences accepted by $h_{j}$, and (ii) $\mathcal{P}(h_{i})\leq\mathcal{P}(h_{j})$, where $\mathcal{P}(\cdot)$ orders correctness guarantees. In PRISM:
  • Layer-1 (Generation-Time Enforcement). Prefix-decidable structural constraints $\mathcal{C}_{\mathrm{struct}}$ are compiled into DFA/PDA automata for token-level masking. L1 enforces structural and lexical correctness through reachability-preserving transitions.
  • Layer-2 (Validation-Time Certification). Global semantic/logic constraints $\mathcal{C}_{\mathrm{sem}}$ and $\mathcal{C}_{\mathrm{log}}$ are compiled into L2 semantic and logic validators and run post-generation. L2 emits machine-checkable certificates for semantic and logical correctness.
Under this hierarchy, L1 has stronger online enforceability guarantees but weaker expressiveness; L2 has broader expressiveness but applies post hoc. Both are required for safety-critical assurance.
Unified Automaton Execution: Formal Semantics.
All L1 constraints are compiled (Section 3.3) into an automaton that provides generation-time guidance and an auditable trace. Let $s_{t}$ denote the automaton configuration (state or stack encoding) at decoding step $t$. We define the allowed token set by
(18) $M_{t}=\mathrm{Allow}(s_{t})=\{y\in\Sigma\mid\delta(s_{t},y)\neq\bot\},$
where $\Sigma$ is the token vocabulary and $\delta$ is the automaton transition function.
Compilation proceeds as follows:
  • Finite-state fragments (JSON Schema / Regex / FSM). These are determinized into a prefix-closed deterministic finite automaton (DFA) $\mathcal{A}=(Q,\Sigma,\delta,q_{0},F)$. For this pure DFA case, we write $q_{t}\in Q$ for the automaton state at decoding step $t$ (so here $s_{t}=q_{t}$), and every reachable non-final state $q_{t}\notin F$ still satisfies $\exists w:\delta^{*}(q_{t},w)\in F$ (accepting reachability).
  • Context-free fragments (GBNF / CFG). These are enforced via a PDA/LR-style automaton that maintains a parsing stack. For unified token masking and audit logging, we apply bounded-depth unfolding to depth $d$ (the same depth bound $d$ used in Theorem 8) to obtain an equivalent DFA $\mathcal{A}_{d}$ whose states encode stack configurations up to depth $d$. Depth $d$ trades off structural expressiveness, automaton size, and runtime cost.
The generation loop is then:
  1. 1.
    The executor queries $s_{t}$ to compute $M_{t}$.
  2. 2.
    The LLM produces logits $z_{t}\in\mathbb{R}^{|\Sigma|}$.
  3. 3.
    Tokens $y\notin M_{t}$ are masked by setting $z^{\prime}_{t}[y]=$-$\infty$. A valid token $y_{t}$ is sampled from $\mathrm{softmax}(z^{\prime}_{t})$.
  4. 4.
    The automaton transitions to $s_{t+1}=\delta(s_{t},y_{t})$. We record an audit tuple $\tau_{t}=\langle s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t\rangle$, where $s_{t+1}=\delta(s_{t},y_{t})$. Optionally a debug variant $\widetilde{\tau}_{t}$ serializes $M_{t}$ itself when needed.
The trace $\tau=\{\tau_{1},\ldots,\tau_{T}\}$ is streamed to the Evidence Registry. From $\tau$ we extract a DFA acceptance path $\rho$ and generate $\pi_{\text{struct}}=\langle\rho,\text{cert}_{L1}\rangle$. Each run is bound to a specific compiled automaton via automaton_id =H(Schema/GBNF source)=H($\text${Schema/GBNF source}), so that external auditors can re-validate L1 conformance using only the artifact, $\pi_{\text{struct}}$, and the automaton definition—without trusting the LLM itself. After generation, L2 validators (semantic validators for $\mathcal{C}_{\mathrm{sem}}$, logic validators for $\mathcal{C}_{\mathrm{log}}$) operate independently on the completed artifact, producing $\pi_{\text{sem}}$ and $\pi_{\text{logic}}$ that, together with $\tau$, form the composite evidence $\Pi$ (Section 3.5).
Expressiveness Baseline and Controlled Restriction.
Before analyzing quality trade-offs, we state an upper bound on expressiveness for unconstrained decoding.
Theorem 10 (Maximum Expressiveness of Unconstrained Strategy).
Let $h_{\mathrm{free}}$ denote unconstrained LLM decoding. Then $\mathcal{L}(h_{\mathrm{free}})=\mathcal{L}_{\text{LLM}}$, the full language model distribution over token sequences.
Proof sketch.
$h_{\mathrm{free}}$ samples directly from $P_{\theta}(y_{t}\mid y_{<t})$ with no masking. Any constrained strategy prunes tokens not in $M_{t}$, restricting the reachable sequences to a (possibly strict) subset. Thus $\mathcal{L}(h_{\mathrm{free}})\supseteq\mathcal{L}(h_{\text{constrained}})$ for any constrained strategy. ∎
Unconstrained decoding maximizes linguistic fluency and variety but does not guarantee structural, semantic, or logic correctness. Our L1/L2 split deliberately trades some expressiveness for verifiability and auditability (cf. the repair convergence analysis in Section 3.6).
Generation Quality vs. Structural Guarantees.
Enforcing DFA-based masking at L1 alters the model’s token distribution. This can bias the model toward minimally compliant completions, an effect observed in grammar-aligned decoding literature (Park et al., 2025a). We characterize this effect via parameterized grammars.
Definition 3.0 (Parameterized GBNF).
A parameterized GBNF is a tuple $\mathcal{G}_{\theta}=(N,\Sigma,R_{\theta},S)$ where $N$ is the set of non-terminals, $\Sigma$ is the token vocabulary, $S\in N$ is the start symbol, and $R_{\theta}$ is a set of production rules parameterized by $\theta\in\Theta$. Each production $A\to_{\theta}\alpha$ may include parameters $\theta_{k}$ that enable/disable alternatives or adjust rule preferences, where $\alpha\in(\Sigma\cup N\cup\{\theta_{k}\})^{*}$.
Execution semantics. GBNF/CFG fragments are enforced by a PDA/LR automaton maintaining a parse stack. When unified token masking and auditable traces are required, we apply bounded unfolding to depth $d$ to obtain an equivalent DFA $\mathcal{A}_{d}$ whose states encode stack configurations up to depth $d$. This preserves reachability guarantees (and thus prefix safety) within the bound $d$, while making the enforcement compatible with the DFA-style masking loop above.
Even with parameterized grammars, strict masking can distort generation quality. To mitigate this, PRISM employs three runtime strategies tailored to safety-critical model-driven engineering:
  1. (i)
    Minimum structural coverage. Before allowing EOS, the executor enforces configurable completeness thresholds: minimum token counts per structural component, presence of all semantically critical “optional” elements, and coverage of domain-specific fields. This prevents premature termination that would otherwise satisfy the DFA but yield under-specified artifacts.
  2. (ii)
    Two-stage generation with targeted refinement. Phase 1 produces a structurally valid draft under tight DFA/PDA constraints (guaranteeing correctness). Phase 2 applies targeted refinement using AGR (Section 3.6), relaxing only semantic/logic aspects to enrich detail and style while preserving the structural invariants already certified by L1.
  3. (iii)
    Calibrated unfolding depth. We tune the unfolding bound $d$ from Theorem 8 to preserve essential branching in $\mathcal{A}_{d}$. Shallower unfolding ($d<d_{\text{max}}$) reduces automaton size and runtime cost while still permitting diverse valid continuations, mitigating over-pruning and mode collapse.
In practice, these heuristics preserve usefulness and readability of generated artifacts while maintaining formal guarantees.
Verification-Guided DFA Masking and Prefix Safety.
L1 enforcement guarantees that every prefix remains correctable to an L1-valid artifact. We compile all structural constraints in $\mathcal{C}_{\mathrm{struct}}$ into a single prefix-closed DFA $\mathcal{A}=(Q,\Sigma,\delta,q_{0},F)$ (or its bounded unfolding $\mathcal{A}_{d}$), where every non-final state $q\notin F$ has a path $q\leadsto F$ to acceptance. During generation, after emitting prefix $y_{1..t}$ reaching $q_{t}$, the executor tests each candidate token $y\in\Sigma$ by computing $q^{\prime}=\delta(q_{t},y)$ and admits $y$ into $M_{t}$ iff $q^{\prime}\leadsto F$. If EOS occurs in non-final $q_{t}\notin F$, the executor computes the shortest accepting suffix $\mathcal{T}^{\star}(q_{t})$ via breadth-first search and appends it automatically, yielding structural repair bound $\lambda_{\text{struct}}\leq 1$.
Theorem 12 (Prefix Safety and Bounded Structural Repair).
Under the unified automaton execution framework, for any prefix $y_{1..t}$ produced by the verification-guided decoder over a prefix-closed DFA $\mathcal{A}$ (including bounded-unfolded $\mathcal{A}_{d}$), the current state $q_{t}$ satisfies: (i) $q_{t}\leadsto F$(there exists a continuation to acceptance); (ii) there exists a unique shortest completion $\mathcal{T}^{\star}(q_{t})$ of length $L(q_{t})=\min\{|w|:\delta^{*}(q_{t},w)\in F\}$; (iii) applying at most one structural edit based on $\mathcal{T}^{\star}(q_{t})$ yields an artifact that satisfies all L1 structural constraints.
Proof sketch.
(i) and (ii) follow from the reachability-based masking policy and the breadth-first computation of $\mathcal{T}^{\star}(q_{t})$. For (iii), if the final state $q_{T}\notin F$ but $q_{T}\leadsto F$, appending $\mathcal{T}^{\star}(q_{T})$ performs a single closure step under the same structural parent in the parse tree. By construction this is the minimal edit that reaches $F$, bounding the structural repair parameter by $\lambda_{\text{struct}}\leq 1$. Empirically, we observe $\lambda_{\text{struct}}\approx 0.03$ (3% of cases require such repair) on AUTOSAR benchmarks (Section 4.3). ∎
Complexity Analysis.
Let $Q$ be the DFA state set after dead-state elimination, $k=|Q|$ the number of reachable states, and $|\Sigma|$ the token vocabulary size. Each decoding step examines at most $k\cdot|\Sigma|$ candidate transitions to derive $M_{t}$, yielding time complexity $\mathcal{O}(k\cdot|\Sigma|)$ per step and space complexity $\mathcal{O}(k)$ for state storage. The one-step structural closure (computing $\mathcal{T}^{\star}$ by breadth-first search) costs $\mathcal{O}(k)$. In practice, with $k\approx 10^{3}$ for AUTOSAR-scale schemas and $|\Sigma|\approx 10^{4}$ for modern LLM vocabularies, this overhead remains tractable (<1<1ms/step on commodity hardware).
3.5. Constraint-Guided Verifiable Generation
Role in PRISM.
Constraint-Guided Verifiable Generation (CVG) is the final stage of the PRISM pipeline. Given an artifact $a$ generated under Layer-1 (L1) structural guidance (Section 3.4) and validated under Layer-2 (L2) semantic/logic certification (Section 3.3), CVG assembles machine-checkable evidence that proves conformance to domain constraints and binds that evidence to the artifact. The result is a verifiable artifact that can be audited by an external authority without rerunning the (possibly proprietary) generation process. Unlike purely post-hoc verification approaches (Misu et al., 2024; ParthasarathyGaurav et al., 2024), PRISM records generation-time traces and post-generation validation results and fuses them into a single composite certificate.
Theoretical Foundation: Map of This Section.
For clarity, we organize the formal development as follows:
  1. (i)
    Verifiable artifacts and composite evidence (Definition 13): how we package an artifact $a$ with its evidence $\Pi$ and a verifier $\varphi$.
  2. (ii)
    Evidence composition (Definitions 1416): how structural, semantic, and logic traces are combined using $\otimes$ (sequential composition) and $\oplus$ (audit enrichment).
  3. (iii)
    Sound compilation (Theorem 8, Section 3.3): why L1/L2 validators preserve the semantics of source constraints $\mathcal{C}_{\mathrm{struct}}$, $\mathcal{C}_{\mathrm{sem}}$, $\mathcal{C}_{\mathrm{log}}$ under bounded depth $d$.
  4. (iv)
    End-to-end correctness: Section 3.4 establishes prefix safety and bounded structural repair (Theorem 12); later, Theorem 18 (Section 3.5) establishes incremental trace soundness for whole-artifact guarantees.
Together, these elements ensure that each generated artifact is paired with independently checkable, regulator-ready evidence $\Pi$, not just with an informal “the model said so” claim.
Definition 3.0 (Verifiable Artifact with Composite Evidence).
A verifiable artifact is a tuple $\mathfrak{A}=(a,\Pi,\varphi)$ where:
  • $a\in\mathcal{L}$ is the generated artifact (e.g., an AUTOSAR ARXML instance);
  • $\Pi$ is a composite evidence object aggregating layer-specific traces and runtime audits;
  • $\varphi:\Pi\to\{\top,\bot\}$ is a verifier that evaluates $\Pi$ and returns whether the evidence proves that $a$ satisfies the required constraints.
Evidence Composition Operators.
We next formalize how $\Pi$ is assembled from per-layer traces.
Let $\mathcal{T}$ denote the space of validation traces with temporal ordering $\preceq_{t}$. A validation trace $\pi=(T,V,\sigma)$ comprises timestamps $T$, validator outcomes $V$, and a satisfaction mapping $\sigma$ from constraints to validation status.
Sequential Composition ($\otimes$). For traces $\pi_{1}=(T_{1},V_{1},\sigma_{1})$ and $\pi_{2}=(T_{2},V_{2},\sigma_{2})$:
(19) $\pi_{1}\otimes\pi_{2}$ $=(T_{1}\cup T_{2},\;V_{1}\cup V_{2},\;\sigma_{1}\cup\sigma_{2})$
$\quad\text{subject to }\max(T_{1})\preceq_{t}\min(T_{2}),$
i.e., later-stage validation (e.g., L2 semantic checks) must occur after earlier-stage validation (L1 structural checks).
Audit Enrichment ($\oplus$). Given a validation evidence object $\Pi_{0}$ and a runtime audit trail $\tau=\{\tau_{1},\ldots,\tau_{T}\}$ (Definition 17):
$\Pi_{0}\oplus\tau\;=\;\Pi_{0}\cup\{\text{metadata}(\tau)\},$
where $\text{metadata}(\tau)$ captures process-level observability (e.g., automaton state transitions, masking statistics, timing), without modifying Boolean pass/fail states.
Verifier composition. The verifier $\varphi$ is required to satisfy:
$\varphi(\pi_{1}\otimes\pi_{2})=\varphi(\pi_{1})\wedge\varphi(\pi_{2}),\qquad\varphi(\Pi_{0}\oplus\tau)=\varphi(\Pi_{0}).$
Thus, $\oplus$ enriches evidence with audit metadata without changing correctness, while $\otimes$ enforces conjunctive validity across layers.
Definition 3.0 (Composite Evidence and Composition Operators).
Let $\pi_{\text{struct}}$, $\pi_{\text{sem}}$, $\pi_{\text{logic}}$ be the per-layer validation traces for structural, semantic, and logic constraints, and let $\tau$ be the runtime audit trail.
For notational alignment with the system diagrams, define $\pi_{1}:=\pi_{\text{struct}}$ and $\pi_{2}:=\pi_{\text{sem}}\otimes\pi_{\text{logic}}$.
We then write the composite evidence as
$\Pi=(\pi_{1}\otimes\pi_{2})\oplus\tau=\big(\pi_{\text{struct}}\otimes\pi_{\text{sem}}\otimes\pi_{\text{logic}}\big)\oplus\tau.$
Operationally:
  • $\otimes$ (sequential composition) appends validator outputs in causal order (L1 $\rightarrow$ L2), preserving timestamps and dependency structure.
  • $\oplus$ (audit enrichment) injects runtime audit metadata $\tau$ into the composed traces without altering acceptance.
Finally, $\varphi(\Pi)=\top$ if and only if $\varphi(\pi_{\text{struct}})=\varphi(\pi_{\text{sem}})=\varphi(\pi_{\text{logic}})=\top$. When $\varphi(\Pi)=\top$, we seal the result using a temporal hash $H(a\,\|\,\Pi\,\|\,t)$ to bind artifact $a$, evidence $\Pi$, and timestamp $t$ into an immutable record.
Lemma 15 (Audit Enrichment Conservativity).
Let $\Pi_{0}=\pi_{\text{struct}}\otimes\pi_{\text{sem}}\otimes\pi_{\text{logic}}$ and $\Pi=\Pi_{0}\oplus\tau$. If $\varphi(\Pi_{0})=b$ for $b\in\{\top,\bot\}$, then $\varphi(\Pi)=b$.
Proof sketch.
By Definition 14, $\oplus$ appends audit metadata $\tau$ but does not alter the Boolean outcomes of $\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}}$. Since $\varphi$ is defined as the conjunction of those outcomes, $\varphi(\Pi)=\varphi(\Pi_{0})$. ∎
Lemma 15 guarantees that attaching process-level provenance (forensics, performance metrics) cannot “flip” a pass into a fail or vice versa. Evidence is extensible but judgment is stable.
Compositional Verification Architecture.
Figure 7. PRISM’s validation and evidence composition. The generated artifact and the decode-time audit trail are checked by structural, semantic, and logic validators derived from the ICM. Each stage emits a machine-checkable trace. These traces are bundled into a single evidence package that is used both to accept or reject the artifact and to drive targeted repair when checks fail.:
Figure 7 shows that verification is modular. Each validator operates on the completed artifact $a$ but targets one constraint family: L1 structural correctness ($\mathcal{C}_{\mathrm{struct}}$), L2 semantic conformance ($\mathcal{C}_{\mathrm{sem}}$), and L2 logic/temporal consistency ($\mathcal{C}_{\mathrm{log}}$). This modularity yields three critical properties: (i) independent auditability: third parties can re-check a single layer using only $a$ and the corresponding $\pi_{\bullet}$; (ii) incremental repair: a semantic or logic violation at L2 does not erase the structural guarantees at L1, enabling targeted AGR; (iii) composable evidence: $\Pi$ is literally the conjunction (via $\otimes$) of per-layer certificates.
Layer-Specific Traces.
Definition 3.0 (Compositional Verification Trace).
Given constraint families $\mathcal{C}=\mathcal{C}_{\mathrm{struct}}\cup\mathcal{C}_{\mathrm{sem}}\cup\mathcal{C}_{\mathrm{log}}$, the composite verification trace for artifact $a$ is:
$\mathcal{T}_{a}\;=\;\bigotimes_{i\in\{\text{struct},\text{sem},\text{logic}\}}\mathcal{T}_{i}\;\oplus\;\tau,$
where:
  • $\mathcal{T}_{\text{struct}}=\langle Q,\delta,q_{0},F,\rho\rangle$: DFA execution trace for L1 structural constraints. $Q$ is the DFA state set, $\delta$ the transition function, $q_{0}$ the initial state, $F$ the accepting states, and $\rho:\{1,\ldots,T\}\to Q$ the run induced by generation. Acceptance requires $\rho(T)\in F$.
  • $\mathcal{T}_{\text{sem}}=\langle\mathcal{S},\mathcal{V},\sigma\rangle$: Semantic validation report for L2 semantic constraints. $\mathcal{S}=\{s_{1},\ldots,s_{n}\}$ are semantic constraint specifications (e.g., SHACL shapes), $\mathcal{V}\subseteq\mathcal{S}$ are violated shapes, and $\sigma:\mathcal{S}\to\{\text{conforms},\text{violates}\}$. Acceptance requires $\mathcal{V}=\emptyset$.
  • $\mathcal{T}_{\text{logic}}=\langle\Phi,\mathcal{M},\mu\rangle$: Logic validation certificate for L2 logic/temporal/numeric constraints. $\Phi$ are logic constraint formulas (e.g., SMT-LIB2), $\mathcal{M}$ is a satisfying model (if $\Phi$ is satisfiable) or an unsat-core $\mathcal{U}\subseteq\Phi$, and $\mu$ is the solver proof object. Acceptance requires $\Phi$ to be satisfiable.
  • $\tau$: runtime audit trail (Definition 17) capturing generation-time decisions and constraint enforcement, sealed by a temporal hash $H(a\,\|\,\Pi\,\|\,t)$ for tamper evidence.
Runtime Audit Trail: Capturing Generation Decisions.
The runtime audit trail $\tau$ is the process-level provenance of how the LLM was guided by L1 constraints (Section 3.4). It complements the post-hoc validators by documenting how the artifact was produced.
Definition 3.0 (Audit Tuple Structure).
At decoding step $t$, CVG records:
$\tau_{t}\;=\;\langle s_{t},\;|M_{t}|,\;y_{t},\;s_{t+1},\;\Delta t\rangle$
where:
  • $s_{t}\in Q$ is the automaton configuration (DFA state or PDA stack) before emitting the token;
  • $|M_{t}|\in\mathbb{N}$ is the size of the allowed token set $M_{t}=\mathrm{Allow}(s_{t})=\{y\in\Sigma\mid\delta(s_{t},y)\neq\bot\}$, which quantifies how restrictive the mask was;
  • $y_{t}\in\Sigma$ is the token actually selected after masking;
  • $s_{t+1}=\delta(s_{t},y_{t})$ is the successor configuration;
  • $\Delta t\in\mathbb{R}^{+}$ is per-step latency for performance profiling.
The full audit trail is the sequence $\tau=\{\tau_{1},\ldots,\tau_{T}\}$.
Memory-Aware Audit Recording.
To keep runtime overhead tractable, the default mode records only
$(s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t)$ per step. Storing $|M_{t}|$ instead of the full $M_{t}$ keeps per-step cost $\mathcal{O}(1)$ rather than $\mathcal{O}(|\Sigma|)$. When full forensic mode is enabled, and when $|M_{t}|<1000$, the framework may also log the entire $M_{t}$ for that step. The transition $\delta(s_{t},y_{t})$ is stored implicitly via $(s_{t},s_{t+1})$, reducing storage from $\mathcal{O}(|Q|\cdot|\Sigma|)$ to $\mathcal{O}(T)$.
Automaton Identity Binding.
Each audit trail $\tau$ is bound to a specific compiled automaton via
$\textit{automaton\_id}=H(\text{Schema/GBNF source}),$
where $H$ is a cryptographic hash. This lets an auditor reconstruct the DFA/PDA from the same source constraints (ICM-derived $\mathcal{C}_{\mathrm{struct}}$) and check that $\tau$ is a valid accepting run without trusting the LLM or rerunning generation. In production, automaton_id can simply be a version-control commit hash for the constraint schema.
Evidence Registry: Append-Only Storage.
All evidence—including $\tau$, $\pi_{\text{struct}}$, $\pi_{\text{sem}}$, $\pi_{\text{logic}}$, and the composed $\Pi$—is stored in an append-only Evidence Registry. To balance auditability with throughput, the registry uses a two-tier strategy: (i) an in-memory bounded LRU cache to keep active trails for near-term refinement and AGR; and (ii) asynchronous persistence to append-only durable storage. Entries are content-addressed by cryptographic hash, enabling deduplication when multiple artifacts share constraint subsets. A $\Pi$-manifest records $\Pi=(\pi_{\text{struct}}\otimes\pi_{\text{sem}}\otimes\pi_{\text{logic}})\oplus\tau$ together with version metadata (v1, v2, final), supporting iterative repair cycles.
Incrementality and Layer Scope.
L1 validation is prefix-decidable (Section 3.4): at each decoding step, the DFA run $\rho$ can be extended and checked. By contrast, L2 semantic and logic validators require the complete artifact to build global symbol tables, evaluate cross-file references, or solve temporal/logical constraints. This asymmetry yields the following guarantee.
Theorem 18 (Incremental Trace Soundness).
Let $a^{(t)}=y_{1}y_{2}\cdots y_{t}$ be a partial artifact at decoding step $t$. If the incremental structural trace $\mathcal{T}^{(t)}_{\text{struct}}$ satisfies $\varphi(\mathcal{T}^{(t)}_{\text{struct}})$$\top$ (i.e., the DFA run $\rho^{(t)}$ remains reachable to an accepting state), then for any structurally valid completion $a=a^{(t)}\cdot w$, there exists a trace extension $\Delta\Pi$ such that $\varphi(\mathcal{T}^{(t)}_{\text{struct}}\oplus\Delta\Pi)=\top$.
Proof sketch.
By Theorem 12, if the current DFA state $q_{t}$ is still reachable to some accepting state $q_{f}\in F$, there exists a shortest accepting suffix $\mathcal{T}^{\star}(q_{t})$. Appending this suffix yields a structurally valid completion. For conjunctions of multiple structural constraints, the sequential composition operator $\otimes$ preserves conjunctive satisfaction. Thus any prefix that has not violated L1 constraints can be extended to a fully L1-valid artifact. ∎
Theorem 18 formalizes prefix safety for evidence: if we have not violated L1 yet, we can finish in compliance. L2 (semantic/logic) evidence, however, is appended only once the artifact is complete and global context is known.
Engineering Considerations for Deployment.
The CVG runtime and Evidence Registry must serve both high-throughput generation and high-assurance auditing. To that end:
  • Configurable audit granularity. Deployments can choose: full logging (record every $\tau_{t}$ and, when $|M_{t}|$ is small, full $M_{t}$), summary mode (record only aggregate statistics and final DFA acceptance), or key-event logging (log only structural boundaries, cross-file references, and validator feedback points). This tunes overhead versus forensic richness.
  • Backend abstraction. Different grammar-constrained decoders (e.g., LR/PDA backends, DFA masking engines) expose different internal states. The audit layer normalizes these into a canonical tuple $(s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t)$ so that evidence is portable across backends.
  • Deferred context binding. In high-throughput settings with shared schemas, automaton_id binding can be deferred until evidence composition, avoiding expensive schema hashing on the hot path and attaching it only when $\Pi$ is finalized.
Table 1. Constraint hierarchy and evidence mapping. Each constraint family is enforced at its natural layer and emits a corresponding audit/validation artifact.
Type Layer-1 (Structural) Layer-2 (Semantic / Logic) Evidence Artifact
Structural JSON Schema / GBNF $\to$ DFA masking Cross-file reference integrity (deferred) $|M_{t}|$, rejection traces, $\pi_{\text{struct}}$
Semantic Enumerations / local vocab SHACL (hierarchy, reference consistency) SHACL reports in $\pi_{\text{sem}}$
Logic Local numeric guards SMT (temporal, numeric, safety logic) SMT models / unsat-cores in $\pi_{\text{logic}}$
Table 1 summarizes how constraint families map to enforcement layers and to evidence artifacts. Some constraints (e.g., enumerations, basic numeric guards) appear in both layers: L1 enforces the prefix-decidable fragment via DFA masking (preventing illegal tokens), while L2 re-checks global consistency via SHACL/SMT. This cross-layer assignment matches the compilation strategy in Section 3.3 and preserves both preventive guarantees during generation and certifying guarantees after generation.
3.6. Audit-Guided Intelligent Repair
Role in PRISM.
By the time an artifact reaches Audit-Guided Repair (AGR), Layer-1 structural constraints have already been enforced. Prefix-closed DFA/PDA masking plus one-step completion (Theorem 12) guarantee that the draft artifact satisfies $\mathcal{C}_{\mathrm{struct}}$. Residual violations are therefore almost entirely Layer-2 failures: semantic graph obligations ($\mathcal{C}_{\mathrm{sem}}$, via SHACL) and logical/temporal/numeric obligations ($\mathcal{C}_{\mathrm{log}}$, via SMT). Traditional responses are either full regeneration (expensive, nondeterministic) or manual patching without guidance (error-prone). AGR replaces both with evidence-driven, localized repair.
Formally, AGR consumes the composite evidence
[…]
and turns validator diagnostics into two coordinated repair routes:
  1. (i)
    Automated semantic/logic repair. The framework invokes an LLM in a tightly constrained mode to synthesize minimal edits that resolve the specific SHACL/SMT violation, using ICM-grounded context.
  2. (ii)
    Model-driven human repair. The artifact is also materialized into a meta-model-aware configuration editor (“Human Review Layer / MDE Editor” in Figure 1) built from the UMM (Path S1, Section 3.2). A domain engineer applies guided fixes in a structured UI that directly edits typed model classes, not raw text.
Both routes are driven by the same evidence $\Pi$, and both feed successful fixes back into PRISM as reusable constraints (“constraint promotion”), so future generations avoid repeating the violation. The two routes differ only in who performs the edit (LLM vs. human) and how the edit is applied (direct text patch vs. meta-model–aware configuration update). This “graduated automation” design is explicitly supported in the PRISM architecture: the Validation & Repair Layer and the Human Review Layer exchange artifacts, evidence, and compiled validators through the Evidence Registry and Configuration Editor.:contentReference[oaicite:1]index=1
Evidence-Driven Violation Localization.
Given a failed artifact $(a,\Pi,\varphi)$ with $\varphi(\Pi)=\bot$, AGR decomposes $\Pi$ to localize violations with Layer-2 precision:
  • Semantic violations ($\pi_{\text{sem}}$). The SHACL validator returns, for each failing shape: (i) a violation path (an address into $a$ such as an AUTOSAR element or field), (ii) the expected condition (e.g., “must reference an existing Operation”), and (iii) a shape identifier. During ICM compilation (Section 3.3), each shape is annotated with ICM provenance: which ICM entry generated this rule, and which UMM entity it is anchored to. AGR uses this to (a) locate the exact broken reference in $a$, and (b) retrieve the authoritative rule and valid target types.
  • Logic violations ($\pi_{\text{logic}}$). The SMT validator produces an unsat-core $\mathcal{U}\subseteq\Phi$ that pinpoints the minimal inconsistent subset of numeric/temporal predicates. Each predicate in $\mathcal{U}$ is back-linked to concrete fields in $a$ and to its source obligation in the ICM (e.g., timing bounds, resource-usage limits). AGR can therefore highlight which values conflict and under which domain rule.
  • Structural trace ($\pi_{\text{struct}}$). The DFA run $\rho$ and runtime audit $\tau$ remain attached for auditability, but structural incompleteness has already been resolved before AGR via shortest completion $\mathcal{T}^{\star}$ (Theorem 12). AGR typically does not need further structural edits.
In other words, AGR does not have to “search for the bug”: SHACL/SMT produce machine-readable failure tickets that name the failing node, the violated obligation, and its ICM/UMM provenance. This is crucial for both automated and human-assisted repair.
Two Repair Routes: Automated Patch vs. MDE-Guided Edit.
The localized violations are then repaired through one of two routes.
Route A: Automated semantic/logic patching. AGR builds a repair prompt that includes: (i) the failing region of $a$, (ii) the violating constraint with natural-language explanation from ICM provenance, (iii) the expected target type(s) from the UMM, and (iv) any candidate values or references retrieved from the knowledge graph. The LLM (run at low temperature) proposes a minimal edit—for example, replacing an invalid cross-file reference with a valid one of the correct type, or reconciling two conflicting numeric bounds to satisfy the SMT model. The patched artifact $a^{\prime}$ is then locally re-validated on just that region to confirm the violation is cleared without introducing new ones.
Route B: Human-in-the-loop MDE editing. PRISM can also project the artifact $a$ into a configuration editor generated from Path S1 (Section 3.2). Path S1 converts domain XMI/XSD into a Unified Meta-Model (UMM) with typed entities, attributes, associations, and behavioral annotations (CRUD signatures, reference-lookup hooks). This UMM is compiled into strongly typed model classes plus a UI/editor that enforces field types, cardinalities, and reference pickers. In regulated domains (e.g., AUTOSAR), engineers already work with such meta-model-driven editors to author or adjust ARXML configurations.
AGR attaches validator diagnostics to that editor: for each SHACL/SMT violation, the Configuration Editor highlights the offending field(s), surfaces the failing rule (via ICM provenance), and offers valid candidates or safe ranges. A domain expert can then apply a compliant fix using familiar MDE-style interactions (selecting a valid referenced Operation from a dropdown, adjusting a timing bound within admissible limits), instead of free-text editing. After the human edit, the updated artifact is re-validated and re-materialized back into PRISM’s pipeline.:contentReference[oaicite:2]index=2
This route is not an afterthought: it is the fallback path in PRISM’s “graduated automation” model. Routine violations are auto-repaired, but high-stakes or ambiguous cases go through a configuration workflow that is already accepted by certification processes in safety-critical engineering: configure $\rightarrow$ validate $\rightarrow$ repair in an MDE tool.
Constraint Promotion (Feedback to Generation).
Whether a violation is fixed automatically or via the Configuration Editor, AGR performs constraint promotion: it records the violated SHACL shape or SMT clause (together with its ICM provenance and UMM anchor) and feeds it forward into future generation contexts. Concretely, the promoted constraint can: (i) be injected into retrieval-augmented prompts provided to the LLM during Instance Model Generation, (ii) tighten decoder-side guards for specific entity types, or (iii) be persisted in the ICM as a reusable domain rule if it was previously missing. As a result, the system does not merely “patch and forget”; it accumulates domain obligations and prevents recurring violations on subsequent artifacts.
Dependency-Aware Scheduling and Convergence.
AGR orders fixes using the constraint lattice $\Gamma$ (Definition 5), which encodes precedence between obligations (e.g., well-formed reference bindings take priority over numeric tuning). Algorithm 4 summarises the workflow, now with two explicit repair routes and promotion:
Algorithm 4 Layer-2-Guided Repair with Human/LLM Routes and Constraint Promotion
1:Artifact $a$, Composite evidence $\Pi$, Constraint set $\mathcal{C}$
2:Repaired artifact $a^{*}$ or MANUAL_REVIEW; updated generation context
[…]
8:for $v\in\mathcal{V}_{\text{sorted}}$ do
9:  $\text{prov}\leftarrow\text{ICMProvenance}(v)$ $\triangleright$ Back-pointer to ICM entry and UMM anchor
10:  $\text{ctx}\leftarrow\text{RetrieveFromKG}(v.\text{targetEntity},\text{prov})$
11:  if isAutoRepairable(v) then
12:    $\text{prompt}\leftarrow\text{BuildRepairPrompt}(v,\text{ctx},\Pi)$
13:    $a\leftarrow\text{LLM}(\text{prompt},T{=}0.2)$ $\triangleright$ Route A: constrained LLM patch
14:  else
15:    $a\leftarrow\text{ApplyEditorFix}(a,v,\text{prov},\text{ctx})$ $\triangleright$ Route B: human edit in UMM-derived Configuration Editor
16:  end if
17:  $\text{PromoteConstraintToGeneration}(\text{prov})$ $\triangleright$ Persist rule / tighten future prompts and guards
18:  $\mathcal{R}\leftarrow\text{LocalRevalidate}(a,v,\mathcal{C})$
19:  if $\mathcal{R}\neq\emptyset$ then
20:    return MANUAL_REVIEW
21:  end if
22:end for
23:return $a$ as $a^{*}$
The repair convergence guarantee still holds under two observations: (i) Layer-1 structural conformance is already guaranteed and is not re-broken by AGR; and (ii) each remaining violation is either auto-fixable with probability >0.5>0.5 using ICM-grounded prompts, or is surfaced to a domain engineer through a configuration workflow that encodes the UMM and ICM constraints and is already standard practice in model-driven engineering toolchains (configure $\rightarrow$ validate $\rightarrow$ repair). This prevents blind retry loops and yields the empirical $\leq 1$-iteration repair convergence reported in Section 4.3.
4. Evaluation
4.1. Overview and Research Questions
The evaluation validates the PRISM framework through three complementary research questions that progress from single-artifact generation with comprehensive constraint enforcement to multi-artifact system generation under scalability stress, culminating in cross-domain transferability assessment where explicit meta-models are unavailable. This progression addresses three fundamental challenges in LLM-driven artifact generation for regulated domains: establishing formal correctness guarantees for individual artifacts, maintaining consistency across interdependent artifacts at system scale, and demonstrating architectural generalizability beyond domains with rich formal specifications.
We structure the evaluation around the following research questions:
RQ1 (Single-File Generation)::
Can layered constraint enforcement with evidence-carrying generation achieve both structural correctness and high semantic compliance on single-file AUTOSAR components while converging to valid outputs within one repair iteration?
RQ2 (Multi-File System Generation)::
Can the PRISM framework scale to multi-file AUTOSAR systems while maintaining structural correctness, and what are the dominant failure modes when cross-file dependencies introduce global semantic constraints?
RQ3 (Cross-Domain Transferability)::
Can the UMM-ICM-CVG architecture extend to domains without explicit meta-models through inductive constraint extraction while maintaining layered constraint enforcement and evidence-carrying generation?
The primary evaluation domain is AUTOSAR (AUTomotive Open System ARchitecture), a mature model-driven engineering standard for automotive software that provides well-defined meta-models, comprehensive constraint specifications, and industrial validation tooling. AUTOSAR serves as an ideal testbed for RQ1 and RQ2 because it exhibits the complexity characteristics of safety-critical artifact generation—hierarchical component structures, cross-file reference integrity requirements, and rich semantic constraints—while offering authoritative schemas for validation. For RQ3, we evaluate the framework on jurisdiction determination under the Brussels I bis Regulation in Private International Law, a domain lacking explicit meta-models but requiring formal constraint reasoning over precedence hierarchies and prerequisite conditions. This cross-domain probe assesses whether the framework’s modular architecture—separating meta-model construction pathways (S1 deductive transformation versus S2 inductive extraction), constraint compilation (ICM), and constraint enforcement (CVG)—supports deployment beyond engineering domains with established formal artifacts.
Evaluation methodology employs quantitative metrics for structural correctness (XSD validation), semantic consistency (SHACL validation for graph constraints, SMT validation for logic constraints), cross-file reference integrity, repair convergence efficiency, and computational cost. We complement automated metrics with expert review for system-level AUTOSAR configurations, assessing requirements traceability, architectural quality, engineering usability, and toolchain integration outcomes. All experiments use reproducible configurations with fixed random seeds, documented model versions, and archived generation artifacts to support external validation.
4.2. AUTOSAR Domain Preparation
To ensure reproducibility and external validation of the AUTOSAR experiments, we establish a three-layer domain representation that serves as the foundation for all subsequent generation, validation, and repair procedures. This representation comprises the Unified Meta-Model (UMM), the Instance-Constraint Mapping (ICM), and an executable Knowledge Graph (KG), each constructed through systematic transformation and extraction processes.
4.2.1. Unified Meta-Model Construction
The UMM provides a canonical, machine-readable view of AUTOSAR structure. We derive the UMM exclusively through deterministic model transformation from official AUTOSAR metamodel artifacts in XSD (XML Schema Definition) and XMI (XML Metadata Interchange) formats, without relying on unconstrained large language model inference or speculative schema induction. This corresponds to the S1 pathway described in Section 3, where authoritative platform metamodels exist and are lifted into normalized representations.
The transformation process flattens and reconciles the AUTOSAR metamodel into an explicit graph of core types—including software components, ports, interfaces, runnables, signals, and timing attributes—together with their structural relations such as containment, reference targets, allowed cardinalities, and typed links. The UMM thus acts as a canonical vocabulary that names each concept, fixes its attributes, and records which other concepts it may reference. This vocabulary consistency is essential for downstream constraint compilation and validation.
4.2.2. Instance-Constraint Mapping Construction
Building upon the UMM, we construct an ICM that enriches the meta-model with structured constraints extracted from AUTOSAR PDF specifications and related normative documents. The ICM extraction process follows the dual-channel approach described in Section 3.3, combining deductive extraction from formal schemas with inductive extraction from natural-language requirements.
The extracted constraints encompass four primary categories: mandatory element presence requirements, allowed value ranges for attributes, reference integrity obligations that specify compatible target types, and timing or behavioral expectations expressed in machine-checkable form. Each constraint is linked back to UMM types and relations rather than concrete instances, ensuring reusability. For example, the ICM does not specify that a particular component instance must have field $f$; instead, it declares that any element of metamodel type $T$ must expose role $r$ with property $p$. This type-level linkage enables the same constraint to be applied wherever an element of type $T$ is instantiated or validated.
The resulting ICM transforms natural-language obligations from specification documents into structured, typed constraints aligned with the UMM vocabulary, bridging the gap between informal domain knowledge and formal verification requirements. Applying this process to the AUTOSAR Software Component Template specification (AUTOSAR_TPS_SoftwareComponentTemplate) yielded 1,161 normative constraints governing component architecture, port semantics, interface contracts, and behavioral invariants. The entity linking mechanism successfully anchored 1,045 constraints to verified UMM entities, achieving 90% precision. The remaining constraints flagged for human review primarily involved terminological variations and format inconsistencies in the source specification, which undergo manual validation before inclusion in the final ICM. This precision level demonstrates that the dual-channel extraction approach operates at sufficient reliability for production deployment while maintaining the integrity guarantees required for downstream certification workflows.
4.2.3. Knowledge Graph Materialization
The final preparation step materializes a navigable, queryable KG that fuses UMM types and relations with ICM typed constraints. Nodes in the KG correspond to metamodel entities such as software component types, port types, and interface types, while edges capture both structural relations—containment hierarchies, reference targets, cardinalities—and attached constraints, including requirements like ”must provide signal ss” or ”must reference interface of category cc”.
This KG serves dual roles in the experimental environment. First, as a retrieval substrate for prompt construction, the KG enables targeted extraction of relevant metamodel subgraphs rather than arbitrary prose. When generating a specific component, the system retrieves KG subgraphs exposing expected ports, interfaces, and required attributes in a form already aligned with the AUTOSAR metamodel, serving as structural blueprints for generation. Second, as a dynamic schema assembly engine, the KG supports on-demand composition of structural schemas and constraint snippets—such as required fields, allowed references, and value obligations—for downstream generation and validation. Because ICM constraints are linked at the metamodel level, the system can construct appropriate schemas for any candidate component configuration without manual per-instance rule rewriting.
In summary, the AUTOSAR domain preparation yields a three-layer knowledge substrate: the UMM obtained through deterministic transformation from official artifacts, the ICM formed through dual-channel constraint extraction and entity linking, and the KG that operationalizes both structure and constraints for retrieval and dynamic schema construction. This layered representation, containing 1,161 structured constraints over more than 1,000 metamodel entities and 3,049 edges, underpins all AUTOSAR experiments reported in this section.
4.3. RQ1: Single-File AUTOSAR Component Generation
RQ1 evaluates the complete PRISM pipeline on single-component AUTOSAR generation, examining whether UMM-ICM-guided retrieval, layered constraint enforcement, and Audit-Guided Repair deliver structural correctness, semantic compliance, and efficient convergence.
4.3.1. Experimental Setup
Research Question.
RQ1 follows §4.1: single-file AUTOSAR generation with L1+L2 enforcement and one-iteration repair convergence (metrics defined below).
Baseline Comparisons.
We evaluate four pipeline configurations that represent progressively structured approaches from the recent LLM literature, each implemented using the same base model to isolate the impact of constraint enforcement strategies:
  1. (1)
    vLLM (pure prompting baseline): Direct generation without retrieval augmentation or Layer-1 enforcement, mirroring common practice in LLM-based information extraction systems where pretrained models are directly prompted to emit structured records (Dagdelen et al., 2024; Lu et al., 2022).
  2. (2)
    vLLM+RAG: Augments prompts with facts retrieved from the AUTOSAR UMM and KG, analogous to instruction-following text-to-database agents that ground extraction in existing schemas (Jiao et al., 2024).
  3. (3)
    vLLM+RAG+JSON Schema: Enforces Layer-1 constraints via intermediate JSON instances deterministically projected to ARXML, approximating schema-aligned decoding (Lu et al., 2025).
  4. (4)
    vLLM+RAG+GBNF: Applies grammar-constrained decoding using GBNF (GBNF Notation Format) derived from AUTOSAR specifications, masking invalid continuations at each generation step to allow only grammar-admissible prefixes, following the constrained decoding paradigm (Scholak et al., 2021; Beurer-Kellner et al., 2024; Park et al., 2025b).
All pipelines use the same backend model (DeepSeek-R1-Distill-Qwen-32B served through vLLM) with fixed random seeds (42, 1001, 20250701), temperature 0.7, and top-p sampling at 0.9 to ensure reproducibility. We evaluate 60 representative AUTOSAR components under three prompt regimes that vary the amount of natural-language specification context provided: Min (minimal requirements), Std (standard documentation), and Full (comprehensive specifications including timing and behavioral constraints).
Evaluation Metrics.
We measure three dimensions of correctness and efficiency:
  • Structural correctness: XSD validation pass rate indicating syntactic compliance with AUTOSAR schemas.
[…]
  • Repair efficiency: Average repair iterations required for AGR to converge to valid artifacts.
  • Evidence coverage: Audit trail completeness measuring the percentage of generation steps with recorded structural proofs $\pi_{\text{struct}}$.
  • Computational cost: End-to-end latency in seconds and token counts for input and output.
For semantic validation, we focus on a representative subset of the 1,161 extracted constraints covering mandatory element presence, reference integrity, cardinality restrictions, and type compatibility rules. This subset selection addresses practical constraints: AUTOSAR data used in this study are subject to confidentiality restrictions, and commercial configuration tools limit automated harness construction for the full constraint suite. Integration of the complete constraint set remains ongoing work. Within this tested subset, SHACL and SMT (Satisfiability Modulo Theories) validators cover nearly identical violation patterns; we therefore report SHACL results only.
4.3.2. Results
Computational Cost.
Table 2 presents end-to-end latency and token consumption across pipelines and prompt regimes. The pure vLLM baseline exhibits moderate latency and minimal input tokens due to lack of retrieval augmentation, but generates longer outputs containing structural errors. Adding RAG increases input context substantially (from 154–227 tokens to 992–6,649 tokens depending on regime) while improving output quality. Constrained decoding pipelines (GBNF and JSON Schema) reduce output tokens by enforcing compact, schema-compliant generation, with JSON Schema achieving the lowest latency due to deterministic intermediate representation.
Table 2. Computational cost for RQ1: end-to-end latency and token counts by pipeline and prompt regime. All measurements use DeepSeek-R1-Distill-Qwen-32B on the same hardware.
Pipeline Configuration Latency (seconds) Input Tokens Output Tokens
vLLM (Min) 21.87 154 812
vLLM (Std) 29.41 188 1,279
vLLM (Full) 46.31 227 2,013
vLLM+RAG (Min) 17.51 992 743
vLLM+RAG (Std) 33.19 2,534 1,397
vLLM+RAG (Full) 85.60 4,493 3,610
vLLM+RAG+GBNF (Min) 8.01 828 155
vLLM+RAG+GBNF (Std) 59.94 2,896 1,446
vLLM+RAG+GBNF (Full) 123.74 6,649 3,118
vLLM+RAG+JSON Schema (Min) 3.83 848 119
vLLM+RAG+JSON Schema (Std) 13.38 2,605 523
vLLM+RAG+JSON Schema (Full) 31.56 2,915 1,291
Evidence Coverage.
For the vLLM+RAG+JSON Schema pipeline, the structured-output path records complete audit coverage with per-token events and Layer-1 candidate sets, producing structural proof $\pi_{\text{struct}}$ for all 60 test cases across all three prompt regimes. Table 3 summarizes audit statistics. The allowed token set size averages 156 tokens per generation step with minimum of 1 (fully constrained choices) and maximum of 650 (unconstrained text fields), demonstrating effective constraint propagation through the DFA. Audit trace density approximates three events per accepted token, providing fine-grained provenance for verification. The vLLM+RAG+GBNF pipeline was not instrumented with the same audit recorder in our current implementation, representing an engineering limitation rather than fundamental incompatibility with GBNF-based constraint enforcement.
Table 3. Layer-1 auditing statistics for vLLM+RAG+JSON Schema pipeline, showing audit coverage and constraint effectiveness across prompt regimes.
Prompt Regime Allowed Token Set (median / min) Generation Steps Audit Coverage Structural Proof $\pi_{\text{struct}}$ Availability
Min 156 / 1 (max $\leq$ 650) 119 100% 100%
Std 156 / 1 (max $\leq$ 650) 524 100% 100%
Full 156 / 1 (max $\leq$ 650) 1,292 100% 100%
Correctness and Repair Convergence.
Table 4 presents structural and semantic correctness rates alongside repair iteration counts. The pure vLLM baseline fails all XSD validation checks, confirming that unconstrained generation produces structurally malformed ARXML. Adding RAG alone achieves 100% XSD compliance through improved context understanding but yields only 30% SHACL compliance with baseline retrieval, indicating that structural correctness does not guarantee semantic constraint satisfaction.
[…]
Repair convergence validates the AGR mechanism: all pipelines with Layer-1 enforcement converge within one iteration on average, compared to two iterations for the unconstrained baseline. For comparison, an LLM-API configuration using the same SHACL validation suite achieves 50% XSD compliance and 10% SHACL compliance, demonstrating the limitations of API-based generation without systematic constraint enforcement.
[…]
Method XSD Pass Rate $_{\text{base}} () SHACL$$_{\text{enh}}$ (%) Average Repair Iterations
vLLM 0 0 2
vLLM+RAG 100 30 1
vLLM+RAG+GBNF 100 30 1
vLLM+RAG+JSON Schema 100 30 100 1
LLM-API (reference) 50 10 1
4.3.3. Analysis and Discussion
Validation of Layered Constraint Architecture.
The progression from 0% to 100% XSD compliance (pure prompting to constrained decoding) and from 30% to 100% SHACL compliance (baseline RAG to constraint-aware RAG) validates the core architectural hypothesis of PRISM: structural correctness can be guaranteed at generation time through Layer-1 enforcement, while semantic correctness benefits from explicit constraint materialization during the generation phase rather than solely relying on post-hoc validation.
This finding addresses a fundamental challenge in LLM-driven artifact generation: implicit constraint satisfaction from context alone proves insufficient for domains with rich semantic invariants. The baseline RAG configuration achieves 30% SHACL compliance when the LLM must infer constraints from generic metamodel descriptions, demonstrating inherent limitations of unstructured context. In contrast, constraint-aware retrieval that explicitly tags and injects ICM constraints into prompts enables the LLM to satisfy semantic requirements proactively, reducing AGR to handling residual violations from untested constraint categories.
Evidence-Carrying Generation and Audit Guarantees.
The 100% audit coverage and structural proof availability for constrained pipelines demonstrate that evidence-carrying generation is practically achievable without prohibitive overhead. The audit trail density of approximately three events per token provides sufficient granularity for root-cause diagnosis during repair while maintaining reasonable storage requirements. The recorded DFA traces enable deterministic reconstruction of generation decisions, supporting both reproducibility requirements in regulated domains and diagnostic capabilities for failed validation.
Repair Efficiency and Convergence Properties.
The consistent one-iteration convergence for pipelines with Layer-1 enforcement empirically validates the bounded repair iterations property discussed in Section 3.6. The key insight underlying this efficiency is dependency-aware prioritization: AGR exploits the constraint hierarchy to repair structural violations before semantic violations, preventing cascading failures that would require multiple repair rounds.
Constraint Subset and Generalization.
The perfect SHACL compliance achieved on the tested constraint subset should be interpreted within its validation scope. The subset covers core constraint categories representative of typical AUTOSAR component generation scenarios—mandatory element presence, reference integrity, cardinality restrictions, and type compatibility. However, extension to the full 1,161-constraint suite requires addressing practical challenges including confidentiality restrictions on AUTOSAR data and tooling limitations for automated validation harness construction. The ongoing integration effort aims to expand coverage while maintaining the demonstrated correctness guarantees.
Computational Trade-offs.
The latency and token cost results reveal computational trade-offs inherent in constrained generation. Constraint-aware RAG substantially increases input tokens (up to 6,649 tokens in the Full regime) but reduces output tokens through compact, compliant generation. The JSON Schema pipeline achieves lowest latency by exploiting deterministic intermediate representations, while GBNF exhibits higher latency due to per-token automaton state transitions.
Synthesis.
RQ1 validates that the complete PRISM pipeline—encompassing UMM-ICM-guided retrieval, layered constraint enforcement, evidence capture, and AGR—delivers verifiable AUTOSAR component generation with formal correctness guarantees and efficient repair convergence. The results demonstrate that layered constraints, evidence composition, and targeted constraint injection collectively enable trustworthy LLM-assisted software engineering in regulated domains. Critically, this validation encompasses the end-to-end pipeline rather than isolated Layer-1 decoding, confirming that proof-carrying generation with multi-layer verification provides a viable pathway for deploying LLMs in safety-critical artifact generation.
4.4. RQ2: Multi-File System Generation at Scale
The second research question extends the evaluation from single-component generation (RQ1) to system-level, multi-file AUTOSAR configurations, investigating the feasibility boundaries and architectural challenges that emerge when LLM-driven generation must maintain consistency across multiple interdependent artifacts.
4.4.1. Experimental Setup
Figure 8. Two-phase AUTOSAR generation, validation, and repair pipeline. Phase 1 synthesizes a high-level blueprint from NLP requirements and AUTOSAR standards, populating the UMM/ICM database. Phase 2 performs constrained LLM decoding to generate component-specific ARXML with JSON Schema enforcement. Post-generation validation applies XSD, SHACL, and SMT checks, with Audit-Guided Repair generating targeted fixes for any violations to produce a verified Decision Package.:
Research Question and Motivation.
RQ2 tests scalability from single components to multi-file AUTOSAR systems, stressing cross-file references and global invariants that do not arise in RQ1.
Dual-Phase Pipeline Architecture.
To address these challenges, we adopt a blueprint-guided dual-phase generation strategy that decomposes system-level generation into two stages, each with distinct objectives and constraint enforcement mechanisms:
  1. (1)
    Phase 1 (Blueprint Synthesis): The system first generates a JSON-structured blueprint specifying the component inventory, dependency graph, file layout, and cross-component interface contracts. This blueprint serves as a global coordination artifact that establishes naming conventions, reference targets, and architectural topology before detailed component generation begins. The blueprint generation employs JSON Schema constraints to ensure structural validity and completeness of the system plan.
  2. (2)
    Phase 2 (Component Assembly): Guided by the blueprint, the system generates individual ARXML files for each component, with interfaces emitted as separate files referencing a shared basic-types definition. Each generation step uses the blueprint as context to resolve cross-component dependencies and maintain reference consistency. Component-level generation employs the same JSON Schema-constrained decoding used in Phase 1, ensuring structural compliance at the artifact level.
This dual-phase decomposition addresses scalability through separation of concerns: Phase 1 handles global architectural decisions that require system-wide reasoning, while Phase 2 focuses on component-level detail generation constrained by the established blueprint. The blueprint acts as a contract between phases, reducing the context window requirements for Phase 2 generation while providing explicit targets for cross-file references.
Model Selection and Constraint Enforcement.
Unlike RQ1, which evaluated local vLLM deployments with comprehensive audit trail recording, RQ2 employs a frontier API-based model (GPT-5) for both phases. This choice reflects three considerations:
  • Architectural reasoning capabilities: System-level blueprint synthesis requires stronger planning and reasoning abilities to maintain global invariants across complex dependency graphs. Frontier models demonstrate superior performance on multi-step reasoning tasks that involve decomposing high-level requirements into structured component inventories and dependency relationships.
  • Practical deployment scenarios: Many organizations deploy LLM-based tooling through API access rather than self-hosted inference, particularly for occasional system generation tasks where the overhead of local infrastructure deployment is unjustified. RQ2 evaluates a representative deployment scenario complementary to the local vLLM evaluation in RQ1.
  • Constraint enforcement preservation: Although API-based generation precludes the fine-grained audit trail recording demonstrated in RQ1, both phases employ JSON Schema-constrained decoding through the API provider’s structured output capabilities. Phase 1 blueprint generation uses a JSON Schema specifying component inventory structure, dependency graph format, and file layout constraints. Phase 2 component generation uses JSON Schemas derived from AUTOSAR metamodel fragments, ensuring that individual ARXML files satisfy structural requirements before deterministic XML serialization. Thus, Layer-1 constraint enforcement remains active, though without the token-level audit provenance available in the vLLM pipeline.
To balance creative architectural exploration with structural conformance, we adopt stage-specific temperature settings: Phase 1 blueprint synthesis uses temperature 0.7 to encourage diverse architectural solutions, while Phase 2 detailed generation uses temperature 0.3 for more deterministic output, and interface file emission uses temperature 0.2 for maximum consistency. Each system is generated once without seed variation, reflecting single-shot engineering workflows common in industrial practice.
Dataset and Evaluation Metrics.
We curate 20 AUTOSAR systems across three complexity tiers designed to stress different aspects of cross-file generation. The dataset spans three distinct architectural configurations that progressively challenge the framework’s ability to maintain cross-file consistency and semantic correctness.
The Simple tier comprises 7 cases with 27 components total, featuring chain topologies suitable for sensor-processor-actuator pipelines. Representative scenarios include data collection systems, control loops, and monitoring applications. Each system contains an average of 3 to 4 components, generating 52 ARXML files across the tier. The Middle tier consists of 7 cases with 45 components total, employing multi-branch tree topologies that represent sensor fusion architectures and multi-domain controllers integrating powertrain, body, and ADAS subsystems. Systems in this tier average 6 to 7 components each, producing 108 ARXML files. The Complex tier includes 6 cases with 54 components total, utilizing mesh topologies with bidirectional dependencies that model sophisticated scenarios such as autonomous driving stacks, by-wire chassis control, and vehicle-cloud coordination. These systems average 8 to 10 components each and generate 124 ARXML files.
We evaluate generated systems across six primary dimensions. File completeness measures whether all blueprint-declared files are produced without truncation, assessed post-generation for all 284 files across the 20 systems as a binary complete-or-incomplete determination. Structural correctness quantifies syntactic compliance with AUTOSAR XSD schemas through xmllint validation, reporting the XSD validation pass rate for all generated ARXML files. Cross-file reference integrity calculates the resolution rate as the proportion of resolvable references over total <REF> elements, measured by parsing all reference targets and verifying their existence in the generated artifact set. Toolchain usability evaluates whether systems successfully import into commercial AUTOSAR configuration tools, specifically Vector DaVinci Developer, yielding a binary pass-or-fail outcome per system. Generation cost captures both token counts and wall-clock time, measured separately for Phase 1 blueprint synthesis and Phase 2 component assembly across all 20 systems.
Expert review follows a structured protocol evaluating four dimensions: requirements alignment assesses whether generated systems satisfy functional specifications and maintain traceability to original requirements; architectural quality evaluates dependency consistency and type-level correctness; engineering quality measures naming conventions, structural modularity, and estimated repair effort; toolchain integration validates whether systems import successfully into commercial AUTOSAR configuration environments. Reviews are conducted by a domain expert with five years of AUTOSAR development experience and aggregated at the group level to identify systematic patterns rather than system-specific idiosyncrasies.
4.4.2. Results
Structural Correctness and Completeness.
All 20 systems achieve 100% file completeness and 100% XSD validation pass rates across all three complexity tiers, encompassing 284 generated ARXML files (Table 5). Cross-file reference integrity exhibits tier-dependent degradation: Simple systems maintain 100% reference resolution (605 of 605 references resolved), Middle systems decline to 99.4% (1,744 of 1,754 references), and Complex systems reach 92.8% (1,936 of 2,086 references). The overall reference resolution rate across all systems is 96.4% (4,285 of 4,445 references).
Table 5. Structural correctness and cross-file reference integrity by complexity tier. All systems achieve perfect file completeness and XSD compliance, while reference resolution degrades with increasing system complexity.
Complexity Systems Total Files Completeness XSD Pass Rate Total References Resolved References Resolution Rate
Simple 7 52 52/52 (100%) 52/52 (100%) 605 605 100.0%
Middle 7 108 108/108 (100%) 108/108 (100%) 1,754 1,744 99.4%
Complex 6 124 124/124 (100%) 124/124 (100%) 2,086 1,936 92.8%
Total 20 284 284/284 (100%) 284/284 (100%) 4,445 4,285 96.4%
Table 6. Expert review findings aggregated by complexity tier. Scores use 5-point scales for quantitative dimensions; C3 uses categorical assessment (minor/moderate/extensive); D1 reports binary import status. Findings illustrate representative failure patterns observed within each tier rather than universal characteristics.
Dimension Scores Representative Findings
Simple Tier (7 systems, 27 components)
Requirements Alignment A1: 100%; A2: 100%; A3: 5/5 End-to-end functional intent preserved across chain topology. All interfaces emitted as standalone files. Dataflow matches blueprint specifications without omissions.
Architectural Quality B1: 5/5; B2: 5/5 Linear pipeline topology with correctly paired Provided/Required ports. Sender-Receiver and Client-Server interface kinds consistent with specified types.
Engineering Quality C1: 5/5; C2: 5/5; C3: minor edits Consistent naming conventions and modular file separation. No architectural changes required.
Toolchain Integration D1: success All systems import cleanly into commercial AUTOSAR configuration tooling.
Middle Tier (7 systems, 45 components)
Requirements Alignment A1: 100%; A2: 100%; A3: 3/5 Representative issue in one system: Diagnostic coordination module both provides and requires the same Client-Server interface, violating client-server separation. UDS service provided-ports absent on coordination/router components where requirements specified them.
Architectural Quality B1: 2/5; B2: 5/5 Representative issue: Client-Server role confusion in diagnostic subsystem. Local type matching preserved, but global coordination pattern violated.
Engineering Quality C1: 5/5; C2: 5/5; C3: moderate edits Representative remedy: Remove spurious provided services on diagnostic modules; add missing UDS exposure ports.
Toolchain Integration D1: success All systems import despite semantic inconsistencies.
Complex Tier (6 systems, 54 components)
Requirements Alignment A1: 100%; A2: 100%; A3: 2/5 Representative issues: SecOC chain lacks complete tag generation/verification steps. Health module architecture contradicts observer pattern—pulls business data via required-ports rather than consuming pushed health events.
Architectural Quality B1: 1/5; B2: 5/5 Representative issue: Health module pulls business data via required-ports to infer health, contradicting intended observer pattern where producers push health states via provided-ports.
Engineering Quality C1: 5/5; C2: 5/5; C3: partial rewrite Representative remedy: Refactor health monitoring to pure consumer of health events; reestablish SecOC end-to-end tag flow.
Toolchain Integration D1: success All systems import successfully. Structural validity does not guarantee semantic correctness.
Generation Cost and Scalability.
Blueprint synthesis (Phase 1) incurs moderate cost averaging 17,548 to 26,617 input tokens and 127 to 275 seconds across complexity tiers. Component assembly (Phase 2) dominates computational cost, scaling super-linearly with system complexity: Simple systems average 63,959 output tokens and 195 seconds, Middle systems rise to 184,605 tokens and 530 seconds, and Complex systems reach 280,038 tokens and 717 seconds. The approximate 1.5× token increase and 1.35× time increase from Middle to Complex tiers reflect the combinatorial explosion of cross-file reference formation and context integration requirements as dependency graph density increases.
Expert Review Findings.
Table 6 presents aggregated expert assessment across the four evaluation dimensions. For Simple systems, all dimensions score maximally: requirements alignment achieves 100% functional coverage and interface completeness with full traceability (A3: 5/5), architectural quality maintains perfect dependency consistency and type matching (B1: 5/5, B2: 5/5), engineering quality exhibits consistent naming and modular structure requiring only minor edits (C1: 5/5, C2: 5/5, C3: minor), and all systems import successfully into commercial tooling (D1: success).
Middle systems maintain perfect functional coverage and interface completeness but exhibit reduced traceability (A3: 3/5). A representative failure pattern observed in one Middle-tier system involves diagnostic coordination components: a diagnostic module incorrectly both provides and requires the same Client-Server interface, violating the client-server separation principle fundamental to AUTOSAR diagnostics architecture. Additionally, Unified Diagnostic Services (UDS) exposure is incomplete, with diagnostic provided-ports absent on coordination and router components where the requirements specified them. Architectural quality assessment reflects these issues with dependency consistency degrading to B1: 2/5 while local type matching remains correct (B2: 5/5). Engineering quality scores remain high (C1: 5/5, C2: 5/5), with moderate repair effort (C3: moderate edits) required to remove spurious provided services and add missing UDS ports. All Middle systems import successfully despite these semantic inconsistencies.
Complex systems encounter more severe semantic deviations. A representative Complex-tier system exhibits two major issues: the SecOC (Secure Onboard Communication) authentication chain lacks complete tag generation and verification steps, breaking the end-to-end security guarantee; and health monitoring components contradict the intended observer pattern by pulling business data via required-ports to infer health status, rather than consuming health state events pushed via provided-ports as specified. Requirements traceability declines to A3: 2/5, and dependency consistency reaches B1: 1/5, indicating fundamental architectural misalignment. The health monitoring pattern violation exemplifies architectural role drift: the generated component topology is locally type-correct (B2: 5/5) but violates the global observer discipline specified in system requirements. Repair effort escalates to C3: partial rewrite, requiring refactoring of health monitoring subsystems and re-threading of SecOC tag flows. Despite these issues, all Complex systems import into tooling, demonstrating that structural validity alone does not guarantee semantic correctness for system-level configurations.
4.4.3. Analysis and Discussion
Structural Scaling Success and the Blueprint Role.
The structural correctness results presented in Table 5 extend RQ1’s finding that Layer-1 constraint enforcement guarantees structural correctness from single-file to system-level scenarios. Blueprint-guided dual-phase generation with JSON Schema constraint enforcement successfully addresses the syntactic challenges of multi-file artifact generation across all complexity tiers. The blueprint serves three critical functions: establishing globally consistent naming conventions that prevent identifier collisions across files, defining file layout and dependency structure that guides Phase 2 component generation, and providing explicit reference targets that reduce ambiguity during cross-file reference formation.
However, the blueprint’s effectiveness is limited to structural and local semantic properties. While it successfully stabilizes file organization and local type matching (B2 consistently scores 5/5 across all tiers), it fails to enforce global architectural invariants that span multiple components. Expert review findings demonstrate this limitation clearly: dependency consistency (B1) degrades from 5/5 in Simple systems to 2/5 in Middle systems and 1/5 in Complex systems, even as local type matching remains perfect. This divergence indicates that the blueprint captures explicit dependencies declared in component interfaces but fails to encode implicit global constraints such as client-server role separation in diagnostic subsystems or observer-style event propagation in health monitoring architectures.
The Cross-File Reference Boundary.
The progressive degradation pattern observed in Table 5 identifies cross-file reference semantics as the dominant scalability boundary for LLM-driven multi-file generation. This failure mode differs qualitatively from the structural errors observed in RQ1’s unconstrained baseline: rather than producing malformed XML that fails schema validation, the multi-file pipeline generates syntactically valid artifacts that contain semantically inconsistent cross-file references.
Both failure modes—reference hallucination and architectural role drift—stem from limited context windows and independent Phase 2 generation, which preclude global reasoning. Models resolve references using only local context (the current component being generated plus blueprint metadata), lacking the full system view needed to disambiguate similarly-named candidates or detect global pattern violations such as unidirectional data flow or exclusive role assignment.
This finding has important implications for deploying LLM-based generation in multi-file scenarios. While Layer-1 structural constraints prevent local syntactic errors, they cannot alone guarantee global semantic consistency. Addressing the reference resolution boundary requires complementary mechanisms beyond constrained decoding, including explicit reference target canonicalization, stricter interface scoping to reduce choice proliferation, and post-generation validation of system-level invariants before finalization.
Complexity-Dependent Failure Progression.
The tier-specific expert review findings reveal that failure modes intensify with system complexity in predictable patterns. Middle systems encounter isolated role confusion within specific subsystems, affecting localized dependency chains while preserving overall system structure. These failures typically require moderate edits that preserve the generated architectural skeleton. Complex systems exhibit cascading semantic violations that span multiple subsystems, necessitating partial architectural rewrites to restore global invariant satisfaction.
This progression suggests a complexity threshold beyond which purely generative approaches require substantial human refinement. Simple systems with linear topologies and low branching factors remain within the capability envelope of LLM-based generation augmented with blueprint guidance and JSON Schema constraints. Middle systems approach the boundary, achieving high structural quality with localized semantic issues amenable to targeted repair. Complex systems exceed current capabilities for fully autonomous generation, producing structurally sound but semantically inconsistent artifacts that serve as starting points for expert-guided refinement rather than deployment-ready outputs.
Engineering Implications and Mitigation Strategies.
The observed failure modes inform practical deployment strategies for LLM-driven multi-file generation. Three mitigation approaches emerge from the experimental analysis.
First, enhanced blueprint specifications can encode global invariants explicitly. Rather than listing only component interfaces and dependencies, blueprints should specify role assignments determining which components act as clients versus servers in each interaction, communication patterns such as unidirectional data flow and observer relationships, and cross-cutting concerns including security chains and health monitoring topology. This requires extending Phase 1 generation to produce richer architectural specifications that guide Phase 2 generation more tightly.
Second, incremental validation with early feedback can detect semantic violations during Phase 2 generation rather than post-hoc. After generating each component, validating its consistency with previously generated components and the blueprint enables early identification of reference errors and role conflicts. This validation-driven generation approach trades computational cost through multiple validation passes for improved output quality and reduced need for post-generation repair.
Third, hybrid generation with targeted human review acknowledges the complexity boundary identified in this evaluation. For Simple and Middle systems, automated generation with lightweight review focuses on verifying reference resolution and checking for known failure patterns. For Complex systems, generation produces architectural scaffolding that human experts refine to satisfy global invariants, leveraging LLM capabilities for routine structural generation while reserving architectural reasoning for human expertise. This graduated automation model balances efficiency gains from automation with quality requirements for safety-critical deployments.
Model Capabilities and API-Based Generation.
The structural correctness achieved through API-based generation with JSON Schema constraint enforcement provides guarantees comparable to vLLM-based constrained decoding demonstrated in RQ1, despite lacking token-level audit trails. However, the cross-file semantic failures observed in Middle and Complex tiers indicate that stronger base models, while improving blueprint quality and component coherence, do not alone solve the global consistency problem. Even frontier models operating under structural constraints struggle with system-level reasoning when context limitations prevent full-system visibility during component generation.
This finding implies that addressing the multi-file scalability boundary requires architectural innovations—enhanced blueprint specifications, incremental validation, hybrid human-machine workflows—rather than solely relying on improved base model capabilities. While stronger models may reduce the frequency or severity of semantic violations, the fundamental challenge of maintaining global invariants under local generation constraints persists across model families.
4.5. RQ3: Cross-Domain Transferability via UMM S2 Path
We evaluate the framework in a domain without an explicit meta-model by exercising the UMM S2 path (inductive meta-model construction from natural-language sources). This cross-domain probe uses Private International Law (PIL) jurisdiction determination under the Brussels I bis Regulation as a representative case study, and asks whether the UMM–ICM–CVG architecture can be transferred into a high-stakes legal reasoning setting that does not provide a pre-existing, formally specified metamodel. This experiment is intended as an architectural validation of applicability, not as an attempt to build a production-ready legal AI system or to claim comprehensive doctrinal coverage.
4.5.1. Experimental Setup
We examine a legal domain without a formal schema (Brussels I bis). Prior work notes domain hallucination risks (Magesh et al., 2025); we therefore use the UMM S2 path to induce entities/relations/constraints from text and test whether layered enforcement reduces doctrinal errors (Magesh et al., 2025).
UMM S2 Pipeline Architecture.
The evaluation employs the S2 pathway described in Section 3, comprising four stages:
  1. (1)
    LLM-induced meta-model instantiation: Given a fact pattern describing a cross-border dispute, the system induces a UMM mini-instance capturing entities (parties, contracts, properties), roles (plaintiff, defendant, contracting parties), and connecting factors (domicile, contract performance location, property situs) together with their applicability preconditions.
  2. (2)
    Provision linking (ICM construction): The induced factors are mapped to candidate Brussels I bis provisions. For example, immovable property location links to Article 24(1) (exclusive jurisdiction), contract performance location links to Article 7(1)(b) (special jurisdiction), and a choice-of-court clause links to Article 25 (prorogation), optionally paired with Article 31 in parallel-proceedings / first-seised scenarios.
  3. (3)
    Promote-aware constrained decoding (Layer-1): JSON Schema-constrained generation enforces a fixed priority hierarchy when selecting a jurisdictional basis: exclusive jurisdiction (Article 24) outranks agreement-based jurisdiction (Articles 25 and 31), which outranks lis pendens stays (Article 31 alone), which outranks special jurisdiction (Article 7), which in turn outranks general jurisdiction (Article 4) and appearance-based jurisdiction (Article 26). The system commits to one promoted basis before emitting natural-language justification.
  4. (4)
    Post-hoc verification and repair (Layers 2–3): Deterministic auditing checks legal constraints: Article 24 overrides lower bases; appearance (Article 26) cannot displace exclusive jurisdiction; Article 25 invocation requires proper formality and is paired with Article 31 analysis in parallel proceedings; and missing formal prerequisites (e.g., insufficiently supported choice-of-court agreement) force abstention instead of overconfident assignment.
The terminal output is a Decision Package containing (i) an Evidence Registry, which records applicable provisions, connecting factors, and any missing prerequisites, and (ii) a Structured Decision JSON, which encodes the jurisdictional conclusion, the selected forum type, and cited provisions in machine-auditable form.
Figure 9. UMM S2 pipeline architecture for cross-domain Private International Law evaluation. The pipeline instantiates a UMM mini-instance from fact patterns, links induced factors to Brussels I bis provisions through ICM construction, applies promote-aware constrained decoding with JSON Schema enforcement, and performs multi-layer verification with deterministic repair. Output comprises Evidence Registry and Structured Decision JSON forming a complete Decision Package.:
Comparative Configurations.
We compare three configurations using the same API-served model to isolate architectural effects:
  • Baseline (LLM-only): Single-pass prompting without retrieval, schema constraints, or audit. The model receives only the fact pattern and produces free-form jurisdiction analysis.
  • RAG: Retrieval-augmented prompting in the style of commercial legal assistants studied in (Magesh et al., 2025). The model receives the fact pattern plus top-$k$ Brussels I bis provisions retrieved for that scenario, but generation remains unconstrained: it may assert a jurisdictional basis even if statutory preconditions are not actually satisfied.
  • PRISM: Full UMM $\rightarrow$ ICM $\rightarrow$ CVG S2 pipeline. The system first induces the structured mini-instance (UMM), links factors to provisions (ICM), then performs promote-aware JSON Schema-constrained decoding followed by deterministic verification and repair. This configuration enforces precedence, exclusivity, pairing requirements, and formality gates before finalizing the decision.
Dataset and Evaluation Metrics.
The evaluation uses 30 expert-authored Brussels I bis jurisdiction cases spanning the full hierarchy of bases of jurisdiction: exclusive (Article 24), agreement-based (Articles 25 and 31), lis pendens / first-seised stay (Article 31), special jurisdiction (Article 7), general jurisdiction (Article 4), and appearance (Article 26). Each case was drafted and doctrinally validated by a subject-matter expert in European civil procedure. As noted above, this dataset is intended to probe architectural transfer, not to claim full coverage of real-world PIL litigation.
[…]
4.5.2. Results
Table 7 summarizes results across the three configurations. PRISM improves correctness, citation faithfulness, structural compliance, and priority alignment relative to both Baseline and RAG. RAG improves over the Baseline in several dimensions but still fails to guarantee constraint compliance.
Table 7. Cross-domain evaluation on 30 Brussels I bis jurisdiction cases comparing Baseline (LLM-only), RAG retrieval augmentation, and the full PRISM pipeline.
Metric Baseline RAG PRISM
Legal Correctness 0.200 0.300 0.467
Citation Precision 0.250 0.428 0.778
Abstention Quality 0.500 0.800 0.800
Promotion Accuracy 0.367 0.567 0.933
Schema Compliance 1.000
Rule Satisfaction 1.000
Legal Correctness and Citation Quality.
[…]
Constraint Compliance, Abstention, and Priority Alignment.
PRISM achieves perfect Schema-Pass (1.000) and Rule-OK (1.000), meaning every final output is syntactically valid and satisfies precedence, exclusivity, and pairing constraints. Promotion Accuracy (alignment between the claimed forum type and the strongest legally available basis) increases from 0.367 (Baseline) to 0.567 (RAG) and 0.933 (PRISM), indicating that PRISM’s promote-aware decoding and deterministic repair suppress “everything applies” answers and force adherence to the Brussels I bis priority ladder.
Abstention Quality—the ability to defer judgment when required information is missing—is 0.500 for Baseline, 0.800 for RAG, and 0.800 for PRISM. RAG benefits here because retrieved provisions explicitly enumerate preconditions; the model can recognize that those preconditions are not satisfied and refuse to commit. PRISM preserves that epistemic caution while additionally enforcing that abstention is required when, for example, Article 25 formality prerequisites are underspecified.
Computational Cost.
PRISM incurs higher cost due to multi-stage generation, schema-constrained decoding, and post-hoc verification. This overhead is the direct cost of guaranteeing structured legality.
4.5.3. Analysis and Discussion
Architectural Transferability.
These results show that the UMM–ICM–CVG architecture transfers to a non-engineering, text-governed domain (PIL jurisdiction) via the S2 inductive pathway. Without any pre-existing metamodel, the system can induce a structured mini-instance, align facts to candidate provisions, enforce hierarchical precedence, and output auditable decisions. PRISM delivers higher Legal-Correct@1 than both Baseline and RAG (0.467 vs. 0.200 / 0.300), far higher Promotion Accuracy, and perfect Schema-Pass / Rule-OK. This indicates that layered constraint enforcement meaningfully outperforms unconstrained prompting and improves upon retrieval augmentation alone (Magesh et al., 2025).
Discussion.
RAG improves correctness (0.300 vs. 0.200 Baseline) and abstention (0.800 vs. 0.500), echoing (Magesh et al., 2025), but does not enforce doctrinal validity. PRISM’s UMM $\rightarrow$ ICM $\rightarrow$ CVG pipeline adds precedence rules, exclusivity constraints, and formality gates, yielding 0.933 Promotion Accuracy, 0.778 Citation Precision, and perfect structural legality (Schema-Pass = Rule-OK = 1.000). Each case logs UMM instance, factor-provision mappings, decoding trace, and repair actions, providing end-to-end provenance for practitioner review.
This 30-case Brussels I bis evaluation is a feasibility probe testing architectural transfer, not claiming exhaustive legal coverage or deployment readiness. Scaling requires broader doctrinal coverage, multi-jurisdictional support, and human oversight in professional workflows.
5. Conclusion and Future Work
This paper presented PRISM, a high-assurance generation framework integrating LLMs with model-driven engineering and formal methods to produce verifiable, auditable artifacts in safety- and regulation-critical domains. PRISM combines a Unified Meta-Model (UMM) for typed domain semantics, an Integrated Constraint Model (ICM) for constraint aggregation, and Constraint-guided Verifiable Generation (CVG) with Audit-Guided Repair (AGR) to produce verifiable artifacts $\mathfrak{A}=(a,\Pi,\varphi)$. Constraints are stratified across two layers: L1 enforces structural rules during decoding (Theorem 12), while L2 validates semantic/logic constraints post-generation with machine-checkable certificates.
Positioning.
[…]
Limitations.
  • Incremental semantic/logic validation. Our current L2 semantic and logic validators (instantiated as SHACL and SMT) operate on completed artifacts. While L1 is prefix-decidable and enforces $\mathcal{C}_{\mathrm{struct}}$ online, incremental L2 checking of $\mathcal{C}_{\mathrm{sem}}$ and $\mathcal{C}_{\mathrm{log}}$ during long generation (e.g., multi-file systems) remains future work.
  • System-level semantics at scale. In RQ2, architectural drift and dangling cross-file references appear in more complex systems. PRISM can surface these violations and drive AGR, but does not yet guarantee global semantic invariants for arbitrarily dense dependency graphs.
  • Cost and latency. DFA/PDA masking with per-step audit logging introduces decoding overhead, and SHACL/SMT validation plus AGR adds post-generation latency. In regulated workflows this is acceptable, but embedded/real-time deployment (e.g., ECU configuration on-device) may require lighter-weight variants.
  • Human review remains essential. PRISM narrows the surface that humans must review and provides $\Pi$ and $\tau$ to justify remaining issues, but it does not claim unsupervised certification. Certified deployment still requires domain experts and regulatory sign-off.
Future Work.
We outline several directions for advancing PRISM and the broader UMM–ICM–CVG paradigm:
  1. 1)
    Incremental semantic and logic validation. Beyond the post-artifact L2 validation described in Limitations, we aim to integrate streaming SHACL engines (Ye et al., 2023) and incremental SMT solvers accepting partial formulas, enabling Layer-2 feedback during generation. A key challenge is exposing partial, request-scoped constraints without exponential solver call blow-up.
  2. 2)
    Fine-grained constraint strategies. Our current L1 execution uses a unified automaton that encodes JSON Schema / Regex / FSM fragments and bounded-unfolded GBNF. Future variants will mix enforcement modes at block granularity: for critical subtrees we can require fully deterministic automata and one-step closure; for descriptive regions we can allow looser local grammars. This opens the door to adaptive enforcement policies that trade completeness of $\pi_{\text{struct}}$ against linguistic naturalness, guided by live feedback from AGR.
  3. 3)
    Audit-guided fine-tuning and distribution-aware decoding. Although PRISM is model-agnostic and achieved the reported results without domain-specific fine-tuning, the framework exposes systematic opportunities to leverage audit trails for targeted model improvement. The runtime audit logs $\tau$ and AGR repair traces capture precise failure modes and correction patterns that reflect domain-specific constraint violations. These audit-derived signals can guide efficient fine-tuning in vertical domains by identifying high-value constraint patterns where model alignment would reduce repair overhead. Pre-training or fine-tuning on (UMM, ICM, $\Pi$)-annotated corpora could bias $p_{\text{LLM}}$ toward emitting structures that already satisfy $\mathcal{C}_{\mathrm{struct}}$ and respect high-value semantic invariants, reducing AGR workload and improving first-pass success rates, particularly for the $\sim$20% of architecturally dense cases that currently require iterative repair.
    A critical consideration is that grammar-constrained decoding (GCD) may induce distribution distortion by aggressively masking tokens to enforce structural constraints. Recent grammar-aligned decoding (GAD) (Park et al., 2025a) methods aim to respect grammar constraints while preserving the base model’s token distribution. Future work should explore hybrid approaches that combine audit-guided fine-tuning with distribution-aware sampling techniques. Selectively applying stricter GCD enforcement only to safety-critical structural regions while employing GAD-style preservation in descriptive regions could balance formal correctness guarantees with linguistic naturalness. The audit trail $\tau$ provides direct empirical evidence of where enforcement overhead is justified versus where relaxed sampling suffices, enabling data-driven calibration of this enforcement/naturalness trade-off.
    Practitioners must balance infrastructure cost (e.g., vLLM with full per-token auditing) against fine-tuning investment. A systematic decision framework is needed: for a given domain and deployment context, should one invest in targeted fine-tuning of $p_{\text{LLM}}$, or rely on stronger automaton enforcement plus richer KG/RAG retrieval, or adopt a hybrid? This trade-off depends on artifact volume, update frequency of $\mathcal{C}_{\mathrm{sem}}$$\mathcal{C}_{\mathrm{log}}$, and the acceptable amount of human-in-the-loop review for residual hard cases. The layered constraint architecture (UMM–ICM–CVG) is orthogonal to and synergistic with fine-tuning: fine-tuning attempts to align $p_{\text{LLM}}$ with $\mathcal{C}$ before decoding, while CVG guarantees prefix safety, bounded closure, and machine-checkable $\Pi$ during and after decoding. Quantifying this synergy and automating the enforcement/tuning mix is an important next step.
  4. 4)
    Domain applicability and scenario recommendations. The UMM–ICM–CVG architecture exhibits varying degrees of suitability across different LLM application scenarios. For transformations from natural language or multimodal inputs to structured artifacts—the focus of this work—PRISM provides formal correctness guarantees and audit trails that address the trustworthiness requirements of regulated domains. The framework is particularly well-suited to domains characterized by explicit meta-models (or meta-models that can be inductively extracted from specifications), comprehensive constraint specifications spanning structural, semantic, and logical dimensions, and validation requirements that demand machine-checkable evidence of conformance.
    Conversely, direct structured-to-structured transformations that demand perfect consistency may not benefit from LLM-based generation, as deterministic rule-based transformations typically suffice and avoid probabilistic uncertainty. However, structured-to-natural-language generation represents a promising application domain where UMM-based abstraction can strengthen LLM comprehension of source structures and reduce semantic drift. By materializing persistent UMM representations of structured data sources, the framework enables LLM-based analysis and explanation tasks to retrieve accurate semantic metadata and constraint context via RAG mechanisms. This UMM-guided comprehension pathway prevents the model from hallucinating incorrect structural interpretations and ensures that natural language outputs remain grounded in verified domain semantics. Such scenarios include generating documentation from AUTOSAR configurations, producing compliance reports from legal decision structures, or synthesizing explanatory narratives from complex engineering artifacts where the UMM serves as a semantic anchor preventing drift between the structured source and natural language output.
Take-away.
PRISM operationalizes evidence-carrying artifact generation for regulated domains through layered constraint enforcement (L1 structural automata, L2 SHACL/SMT validation) and first-class provenance. This UMM–ICM–CVG pattern offers a practical path toward trustworthy AI-assisted engineering and legal reasoning workflows requiring formal verification and auditability.
Acknowledgements.
We thank Xinyue Yang (School of Law, Wuhan University) for helpful discussions on Brussels I bis, legal analysis regarding exclusive jurisdiction, choice-of-court formalities, and for reviewing preliminary examples and gold labels. All remaining errors are our own. This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant 32427801 (National Major Scientific Research Instrument Development Project). We thank the anonymous reviewers for their valuable feedback.