### v2

# ATLAS: Artifact Generation Through Layered Constraints and LLM $\times$ MDE Synergy

Abstract.
ATLAS unifies Large Language Models with Model-Driven Engineering to generate regulator-ready artifacts and machine-checkable evidence for safety- and compliance-critical domains.
ATLAS integrates three pillars: a Unified Meta-Model (UMM) reconciles heterogeneous schemas and regulatory text into a single semantic space; an Integrated Constraint Model (ICM) extends our prior Dual-Stage (S2D2) extraction logic to compile layered requirements into deterministic generation-time automata (Layer 1) and post-generation validators (Layer 2); and Constraint-Guided Verifiable Generation (CVG) applies these through two-layer enforcement—Layer 1 structural constraints drive prefix-safe decoding while Layer 2 semantic/logical validation produces machine-checkable certificates. When violations occur, ATLAS performs audit-guided repair and records generation traces for compliance review.
We evaluate ATLAS in automotive software engineering (AUTOSAR) and cross-border legal jurisdiction (Brussels I bis). ATLAS produces structurally valid, auditable artifacts that integrate with existing tooling and substantially reduce manual remediation effort, validating a graduated automation paradigm that automates routine construction while empowering experts to resolve complex semantic ambiguities through machine-checkable evidence.
model-driven engineering; meta-model integration; constraint-guided generation; large language models; formal verification; traceable artifacts
1. Introduction
1.1. From PIM–PSM Transformations to Unified Meta-Models
Model-driven architecture (MDA) advocated raising design effort from code to high-level platform-independent models (PIMs), then systematically refining them into platform-specific models (PSMs) and code (omg2003mda; 10.5555/3103551; mardanikoraniModelDrivenEngineeringTechniques2023). Recent advances in meta-model integration aim to provide unified views across heterogeneous domain descriptions by merging multiple domain-specific languages (DSLs) into a single unified meta-model (UMM) (amaralUsingMetamodelComposition2023). However, substantial human curation is still required to align concepts and avoid semantic drift during integration, making the “unification power of models” an active research challenge rather than a solved problem.
1.2. Conformance and Assurance in Safety-Critical Domains
Safety-critical domains demand rigorous conformance to normative standards and bidirectional traceability across the entire development lifecycle (delavaraModelbasedSpecificationSafety2016). In automotive and avionics practice, models such as AUTOSAR (AUTomotive Open System ARchitecture) XML (ARXML) or the Architecture Analysis & Design Language (AADL) must satisfy prescribed structural and semantic rules; certification bodies additionally require that every requirement, design element, code unit, and test case be traceable in both directions (delavaraModelbasedSpecificationSafety2016).
Despite mature certification processes, a practical trade-off remains between assurance and agility: most industrial toolchains rely on rigorous tool qualification rather than per-artifact verification. While effective, this process-heavy approach assumes the generator is static and trusted, creating a toolchain that is expensive to qualify and difficult to evolve when standards or requirements change.
1.3. Pain Points of Conventional MDE Pipelines
Industrial MDE pipelines suffer from three long-standing limitations: (i) Evolution overhead, where brittle chained transformations require costly rework upon schema updates; (ii) Rigidity of assurance mechanisms, where relying on pre-qualified tools restricts the ability to adopt new features or custom heuristics without breaking the chain of trust, often forcing developers back into manual review for unsupported constraints; and (iii) Missing solver-based self-repair, where automated repair spanning structural and logical layers remains confined to research prototypes (mundlerTypeConstrainedCodeGeneration2025; houxinyiLargeLanguageModels2024). These gaps hinder continuous compliance and force recurring review cycles.
1.4. LLMs and Knowledge-Guided Generation: A New Opportunity
Large Language Models (LLMs) such as Codex and GPT, together with open-source peers including StarCoder and DeepSeek-Coder, can already generate compilable multi-file projects directly from natural-language prompts, albeit with varying degrees of structural and semantic correctness (liuNoNeedLift2024; Idrisov2024; Wong2023). Recent work has begun exploring the synergy between LLMs and Model-Driven Engineering: Lebioda et al. (Petrovic_2024) demonstrate LLM-assisted automotive software development through Ecore model instance creation followed by OCL constraint validation, while Alaoui Mdaghri et al. (Alaoui2025) propose leveraging LLMs for DSL modeling from natural language descriptions within an iterative validation framework. Similarly, Patil et al. (Patil2025) show that specification-driven LLM code generation combined with formal verification can produce safety-critical embedded automotive software even without iterative backprompting (Patil2025). However, unconstrained sampling frequently violates syntax rules or domain-specific constraints.
Grammar- and prefix-constrained decoding mitigates these failures by intersecting the model’s token distribution with a formal grammar, ensuring that every incremental prefix remains well formed while incurring negligible loss in fluency. This has been demonstrated by incremental and semantically constrained approaches such as PICARD and Synchromesh (scholakPICARDParsingIncrementally2021; poesiaSynchromeshReliableCode2021), and further optimized for subword alignment and efficiency by DOMINO (zotero-item-1597; gengGrammarConstrainedDecodingStructured2024).
Beyond syntax enforcement, knowledge-graph-augmented prompting injects authoritative domain triples—for example, AUTOSAR component hierarchies—into the LLM context, grounding generation and reducing hallucinations (zhangKnowGPTKnowledgeGraph2024). Finally, iterative feedback-driven refinement is emerging: an LLM proposes an artifact, external validators provide diagnostics, and the model repairs defects based on this feedback (tangCodeRepairLLMs2024). This aligns with recent findings on “verbal reinforcement learning” and self-debugging (shinn2023reflexion; chen2024selfdebug), pointing toward pipelines in which artifacts are generated under explicit, checkable constraints. Early studies report dramatic reductions in constraint violations, pointing toward pipelines in which artifacts are generated under explicit, checkable constraints—a capability largely missing from classical MDE practice.
2. Related Work
2.1. Model Transformations in MDA
Model-Driven Architecture (MDA) automates structured artifact derivation via model-to-model and model-to-text transformations (DualStageFramework; kulkarniSymbioticApproachLeveraging2023; kahaniSurveyClassificationModel2019). While these systematic mappings improve repeatability over ad hoc integration (see Section 1), they typically suffer from inherent rigidity and platform dependence, making it difficult to inject custom logic or to cost-effectively construct meta-models for domains lacking explicit schemas. Our previous work, S2D2 (DualStageFramework), addressed some of these rigidities by introducing Dual-Path Fusion for static structure generation. ATLAS evolves this by moving from static code generation to dynamic constraint enforcement for LLMs.
2.2. Specification and Certification in Regulated Domains
Safety-critical sectors mandate rigorous conformance to standards (e.g., DO-178C, ISO 26262) and structured assurance cases (e.g., MISRA-C, SOTIF). Empirical studies highlight the significant effort required to curate auditable evidence linking safety claims to artifacts (el-gnainyAIEnhancedAUTOSARConfiguration2024a; nairExtendedSystematicLiterature2014). Consequently, industrial toolchains often embed rigid schema validators to ensure “correctness-by-construction,” reducing audit burdens but entrenching specialized, costly-to-evolve tooling.
2.3. Formal Verification and Proof-Carrying Artifacts
Formal methods enrich modeling with machine-checkable semantics via contracts (OCL), model checkers (SPIN, SMT), and interactive proof assistants. The proof-carrying code (PCC) paradigm extends this by attaching correctness certificates, a concept pushed further by solver-aided repair techniques like Max-SAT (orvalhoCounterexampleGuidedProgram2024; mechtaevDirectFixLookingSimple2015). Recent studies demonstrate scalability in domain-specific verification (charalambousNewEraSoftware2023), such as analyzing AUTOSAR latency with UPPAAL (zhangAutomaticModellingVerification2023) or enumerating cut sets in AADL (stewartAADLBasedSafetyAnalysis2021). The ATLAS pipeline inherits this “lift-to-formal-semantics” philosophy by integrating UMM, ICM, and CVG stages to attach proof-carrying certificates to generated artifacts (maSpecGenAutomatedGeneration2025).
2.4. LLM-Assisted Generation under Structural Constraints
Large Language Models (LLMs), including Codex and Claude, can already draft substantial portions of code, configuration files, and AUTOSAR-style XML directly from natural-language prompts. However, naive (unconstrained) sampling frequently violates grammar rules, schema cardinalities, naming conventions, and cross-reference requirements, rendering raw outputs unusable in safety-critical pipelines.
To address this, prior work explores grammar- and prefix-constrained decoding, in which decoding proceeds under a context-free grammar or a deterministic automaton so that every incremental prefix remains syntactically valid, as demonstrated by Synchromesh and PICARD (scholakPICARDParsingIncrementally2021), with negligible fluency loss. An alternative strategy, grammar prompting (wangGrammarPromptingDomainspecific2023), guides the LLM to first synthesize a minimally sufficient grammar tailored to the specific generation task, and then to generate content under that task-specific grammar. This treats constraint specification as an LLM-interpretable guide rather than a hard automaton mask.
More recent decoding algorithms such as DOMINO enforce constraints at the subword level and combine speculative decoding with precomputation to achieve near-zero runtime overhead while still honoring a target grammar (zotero-item-1597). Follow-up work on fast grammar-constrained decoding shows how to align grammar tokens with the model tokenizer and to build the automaton efficiently, reducing offline preprocessing cost by more than an order of magnitude while preserving efficient online masking (Park2025).
In parallel, reinforcement-learning-style schema alignment aims to train LLMs to emit JSON that is provably valid with respect to large, real-world schemas. Nevertheless, even frontier models systematically fail on complex industrial schemas, which has motivated dedicated benchmarks such as SchemaBench and analyses such as JSONSchemaBench that characterize the trade-off between “over-constrained” engines (which reject semantically valid but grammatically atypical structures) and “under-constrained” engines (which accept ill-formed outputs and thus require repair) (Lu2025b). Retrieval- and knowledge-graph-augmented prompting grounds generation in authoritative domain ontologies (for example, AUTOSAR component hierarchies), mitigating hallucinations. Finally, Ferrari and Spoletini show that LLMs can synthesize first-order temporal predicates and safety conditions directly from regulatory prose, effectively extracting normative logical constraints from natural-language standards (ferrariFormalRequirementsEngineering2025).
Emerging industrial pipelines iterate on these trends: an LLM drafts an artifact; for structural layers, a supported subset of JSON Schema / regular expressions / Generalized Backus–Naur Form (GBNF) is compiled into deterministic automata to guide decoding via prefix-safe masking and online tracing; post-generation semantic and logic validators (e.g., SHACL, SMT) then check semantic and logical constraints, capture machine-checkable evidence, and drive automated repair. ATLAS generalizes this pattern into a unified, auditable “verify-as-you-generate” loop.
A critical challenge identified in constrained decoding research is distribution distortion: while grammar masking guarantees structural validity, it can bias LLM generation toward the shortest valid completions, yielding artifacts that are syntactically correct but semantically impoverished (parkGrammaralignedDecoding2025). This effect arises because hard constraint masking eliminates longer yet equally valid continuations, thereby distorting the LLM’s learned probability distribution.
Grammar-Aligned Decoding (GAD) (parkGrammaralignedDecoding2025) addresses this through adaptive sampling with approximate expected futures (ASAp): rather than deterministically masking all invalid tokens, GAD performs limited forward simulation to estimate which candidate tokens can still lead to a grammatical completion, and samples from a distribution that balances grammaticality with fidelity to the base LLM’s preferences. Empirical results on code generation and structured data extraction show that GAD produces outputs with higher likelihood under the base model while maintaining grammatical correctness.
However, GAD’s forward simulation introduces computational overhead that scales with grammar complexity and generation length. For high-assurance domains that generate large-scale artifacts (e.g., thousands of lines) under deeply nested industrial grammars, this overhead can be prohibitive. Moreover, sectors with strict safety-critical requirements may prioritize deterministic structural guarantees over probabilistic grammaticality assurances, because even low-probability violations can have unacceptable consequences. These domain-specific constraints motivate architectures that stratify enforcement across multiple layers with different computational and assurance characteristics, as explored in Section 3.
3. Methodology: The UMM–ICM–CVG Framework
3.1. Overview
Figure 1. ATLAS links domain requirements to verifiable generation. Natural-language requirements and existing meta-models are consolidated into a Unified Meta-Model (UMM) and an Integrated Constraint Model (ICM). Generation is guided by structural constraints (Layer-1, enforced during decoding) and then checked by semantic / logical validators (Layer-2) after decoding. The validation signals, together with the decode-time audit record, form evidence that can be used to automatically repair the artifact in a closed loop.
Block diagram of the ATLAS pipeline. Inputs (NL requirements and meta-model specs) are merged into a unified meta-model and constraint model. L1 automaton-guided decoding enforces structural constraints during generation, L2 semantic and logic validators check the final artifact, and the audit plus validation results feed an audit-guided repair loop.
Motivation and Design Goals.
Industrial, safety-critical domains (e.g., AUTOSAR-compliant automotive software, legal compliance workflows) require not only functionally plausible artifacts but also verifiably correct artifacts that can be audited by certification authorities (yangEvaluation2024; wangMachineDeepLearning2023). Unconstrained Large Language Models (LLMs) excel at understanding domain intent and producing structured drafts, but they do not by themselves guarantee regulatory conformance, global consistency, or traceable justification. In contrast, Model-Driven Engineering (MDE) offers rigor through meta-models, transformation rules, and formal analysis, yet is brittle to evolving standards and costly to maintain as domains change. ATLAS addresses this gap as an end-to-end high-assurance generation pipeline that couples the flexibility of LLMs with the accountability of formal methods. The pipeline is organized around three tightly connected components: (1) the Unified Meta-Model (UMM), which provides an auditable semantic backbone across heterogeneous domain sources; (2) the Integrated Constraint Model (ICM), which aggregates structural, semantic, and logical constraints; and (3) the Constraint-guided Verifiable Generation (CVG) process, which enforces constraints during generation, certifies artifacts after generation, and drives targeted repair when violations are found. Together, these components allow ATLAS to automate the 80% of routine artifact construction while still producing machine-checkable evidence for expert review on the remaining 20%.
Core Components.
ATLAS coordinates LLMs and formal methods in a relay pattern: the UMM and ICM formalize what “correct” means, and CVG—which includes Audit-Guided Repair (AGR)—uses those definitions to control generation, verify conformance, and close the loop with targeted repairs when inconsistencies are detected. We summarize each component below.
Unified Meta-Model (UMM). The UMM is not a passive schema repository; it is the semantic contract that both constrains generation and explains to auditors what each generated element is supposed to mean. To build the UMM, ATLAS merges heterogeneous meta-models and domain ontologies (e.g., AUTOSAR XSD, legacy UML variants) into a coherent, typed graph $M_{\text{UMM}}=(V,E,T,\mathcal{C}_{\text{struct}})$, where nodes denote domain entities (classes, attributes) and edges record structural relations. This merge is carried out using a mathematically well-defined unification procedure (instantiated in our implementation as a category-theoretic colimit construction (diskin2012category)), which ensures that overlapping concepts from different sources are consistently identified and traceable. The point is not that the colimit construction itself is a contribution, but that it lets us claim semantic coherence and auditability: every node in the UMM can be traced back to its source models. For generation, ATLAS materializes relevant subgraphs of the UMM as retrieval-augmented context—e.g., JSON-style schema fragments and textual descriptions—which are injected into LLM prompts to give the model a precise structural and conceptual scaffold.
Integrated Constraint Model (ICM). The ICM collects, normalizes, and compiles constraints that the generated artifact must satisfy. It is populated via two coordinated channels that mirror the strengths of MDE and LLMs:
  • Channel 1 (Deductive, schema-driven). Traditional model transformations extract explicit invariants from domain meta-models: field cardinalities, type hierarchies, mandatory references, and other structural rules. These invariants are organized into a multi-sorted family $(\mathcal{C}_{\text{lex}},\mathcal{C}_{\text{struct}},\mathcal{C}_{\text{sem}},\mathcal{C}_{\text{logic}})$ and compiled into executable artifacts at two layers: Layer-1 (L1) guards are realized as JSON Schema / Regex / GBNF constraints and determinized finite-state controllers for generation-time enforcement; Layer-2 (L2) validators are realized through semantic checkers (e.g., SHACL shapes for graph constraints) and logic checkers (e.g., SMT solvers for numeric/temporal properties) for post-generation validation. Formal analysis provides completeness guarantees: Theorem 7 states that, under bounded unfolding depth $d$, satisfying the compiled constraints implies satisfying the original meta-model invariants.
  • Channel 2 (Inductive, LLM-assisted). Many safety or regulatory requirements in practice are not fully captured by XSDs or UML—for example, “an OperationInvokedEvent must reference an existing operation,” or “mode transitions must form a directed acyclic graph.” Channel 2 uses an LLM extractor $E_{\theta}$ to read natural-language specifications (e.g., AUTOSAR Software Component Template PDFs) that have been augmented with inlined UMM definitions (see Figure 3). $E_{\theta}$ proposes candidate L2-only constraints (expressed as semantic or logic assertions). Before these constraints are admitted into the ICM, ATLAS links each predicate back to UMM entities via entity linking (Definition 3) and checks their semantic compatibility (Definition 5), then runs consistency check to reject contradictory rules (Algorithm 2).
This dual-channel design allows ATLAS to ingest both structured sources (Channel 1) and free-text specifications (Channel 2) without sacrificing rigor: LLM-proposed rules do not bypass formal scrutiny, but instead become auditable constraints that are grounded in—and indexed by—the UMM.
Constraint-guided Verifiable Generation (CVG). CVG is the execution core of ATLAS. It governs how artifacts are produced, checked, and—if necessary—repaired. It proceeds in three stages:
  1. Stage 1:
    Generation with real-time constraint enforcement. Layer-1 constraints are compiled into a prefix-closed automaton (deterministic finite automaton for regular structures or bounded pushdown automaton for context-free patterns, unfolded to depth $d$; see Section 3.4). During decoding, at each step $t$, the executor computes the allowed token set $M_{t}=\{y\in\Sigma:\delta(s_{t},y)\neq\bot\}$, where $s_{t}$ is the current automaton state and $\delta$ is the transition function. Invalid continuations are masked before the LLM samples the next token. This guarantees prefix safety: every partial output always remains on a path that could complete to a structurally valid artifact (Theorem 10). In parallel, ATLAS records an audit trail $\tau_{t}=\langle s_{t},M_{t},y_{t},\delta(s_{t},y_{t}),\Delta t\rangle$, which captures the control decisions that led to each generated token for later auditing.
  2. Stage 2:
    Post-generation formal verification. After the candidate artifact is produced, ATLAS runs Layer-2 validators. Semantic validators check constraints such as cross-file reference integrity and role consistency; logic validators check temporal and numeric properties such as ordering constraints and resource bounds. Each validator emits a machine-checkable certificate $\pi_{\bullet}\in\{\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}}\}$ (e.g., DFA acceptance traces, semantic validation reports, logic proofs/unsat cores). Crucially, these certificates are independent of the generator: an external auditor can re-check correctness using only the artifact $a$, the evidence bundle $\Pi$, and the validator identifiers—without having to trust the LLM itself (Definition 11).
  3. Stage 3:
    Audit-Guided Repair (AGR). If validation fails, ATLAS does not blindly re-generate from scratch. Instead, AGR employs a validation-guided regeneration strategy. The framework analyzes the composite evidence $\Pi$ to localize violations: structural violations via DFA states in $\tau$, semantic violations via SHACL violation paths, and logical violations via SMT unsat cores (Algorithm 4). Unlike blind retries, AGR operates at file granularity but with constraint-aware guidance. It parses the diagnostics to identify specifically which constraints were violated and constructs a targeted repair prompt that explicitly injects these constraints as hard requirements. This narrows the search space and forces the LLM to attend to the specific failure modes (e.g., “missing reference” or “cardinality error”) that were missed in the first pass.
    The localized violations are then repaired through one of two routes:
    Route A: Automated validation-guided regeneration. For standard violations, AGR builds a constraint-aware repair prompt that explicitly injects the violated rules. The LLM performs a diagnostic-driven regeneration of the affected artifact, resolving the specific SHACL/SMT violation.
    Route B: Human-in-the-loop MDE editing. For complex semantic ambiguities, ATLAS projects the artifact into a context-aware configuration editor (automatically derived from the UMM, leveraging the structured transformation Path S1 detailed in Section 3.2). This transforms the human role from manual authoring to evidence-based decision making, allowing experts to resolve undecidable conflicts using machine-generated diagnostics.
System Flow and Human Oversight.
Figure 1 summarizes how these components interact. Natural-language requirements are first interpreted by an LLM-based NLP Processing Layer, while explicit domain schemas are parsed by a Knowledge-Graph Construction pipeline that builds the UMM and populates the ICM through Channel 1 and Channel 2. The Unified Constraint Layer then compiles these constraints into L1 automata and L2 validators. During Instance Model Generation, the LLM receives retrieval-augmented UMM subgraphs as context, and the Structure Decoder enforces L1 constraints through token masking while logging audit tuples $\tau_{t}$ to an Evidence Registry. The objective is to synthesize an instance model $a$ at the M1 layer that strictly conforms to the M2-level definitions in the UMM. After generation, the Validation & Repair Layer executes L2 validators (structural $\rightarrow$ semantic $\rightarrow$ logical), appends each certificate $\pi_{\bullet}$ to the registry, and either: (i) assembles the final verifiable artifact $\mathfrak{A}=(a,\Pi,\varphi)$ by hashing the artifact, the evidence, and a timestamp $H(a\|\Pi\|t)$; or (ii) triggers AGR to request a focused repair from the LLM or, in edge cases, a manual override by a domain expert. This “graduated automation” design lets ATLAS automate routine generation while keeping human reviewers in control of exceptional cases, now equipped with machine-checkable diagnostics instead of ad hoc debugging.
3.2. Unified Meta-Model Construction: A Dual-Path Approach
Problem Formulation and Goals.
In practice, safety-critical domains present two very different starting points for model construction. Some domains (e.g., AUTOSAR, AADL, OPC UA) already publish an explicit meta-model: classes, attributes, and relationships are described in machine-readable form (XSD/XMI/Ecore) or in tightly maintained technical documentation. Other domains (e.g., regulatory and legal workflows, device compliance manuals) do not provide such an explicit schema, even though they implicitly assume a fairly well-structured universe of entities (“who is responsible,” “what object is configured,” “what reference must exist”). ATLAS introduces the Unified Meta-Model (UMM) as the common semantic backbone across both cases. When an explicit meta-model exists, UMM ingests and normalizes it; when it does not, UMM inductively assembles the same kind of typed entity–relation structure from natural-language specifications with LLM assistance. In both cases, the result is a single, provenance-aware representation of the domain’s core entities and relationships that downstream components can treat as the source of truth. Its construction must satisfy three goals: (i) reconcile heterogeneous sources into one typed graph, (ii) maintain provenance so each element in the UMM can be traced back to its origin, and (iii) expose a stable structure that downstream components can rely on for generation, validation, and repair.
Graph-Based Representation.
Each source meta-model is represented as a typed directed graph.
Definition 3.0 (Meta-Model as Typed Graph).
A meta-model is represented as a typed directed graph $M=(V,E,T,\mathcal{C})$, where:
  • $V$: set of nodes representing domain entities (e.g., classes, attributes),
  • $E\subseteq V\times V$: set of edges representing relationships (e.g., associations, generalizations),
  • $T:V\cup E\rightarrow\mathcal{T}$: type mapping function assigning semantic types from a type system $\mathcal{T}$,
  • $\mathcal{C}$: set of structural constraints (cardinalities, mandatory fields, inheritance restrictions).
Given a collection of such graphs $\{M_{i}=(V_{i},E_{i},T_{i},\mathcal{C}_{i})\}_{i\in I}$, the Unified Meta-Model $M_{\text{UMM}}$ is obtained by merging them into a single typed knowledge structure. In our implementation, this merge follows a colimit-style typed graph unification in the sense of (diskin2012category): overlapping entities and relations across different $M_{i}$ are identified and merged through structure-preserving homomorphisms. This provides two desirable properties for ATLAS: (1) semantic coherence—concepts that represent the same domain notion across sources are aligned rather than duplicated; and (2) auditability—every node and edge in $M_{\text{UMM}}$ retains provenance links to its source model(s), allowing downstream validators and human reviewers to trace “where this concept came from.”
Figure 2. Two ways to build the Unified Meta-Model (UMM). Path S1 handles structured sources: when a machine-readable meta-model (for example AUTOSAR, AADL, OPC UA) already exists, ATLAS ingests and normalizes it. Path S2 handles unstructured sources: when only natural-language requirements are available, ATLAS uses an LLM to induce domain entities and relations directly from text. Deployments typically follow one of these two paths, depending on which input is available.
Two-path process for constructing the Unified Meta-Model (UMM). Path S1 applies when a machine-readable meta-model such as AUTOSAR XSD/XMI, AADL, or OPC UA Ecore already exists: ATLAS ingests and normalizes that model using model-to-model transformations. Path S2 applies when no authoritative schema exists and requirements only appear in natural-language documents: ATLAS uses an LLM-based extractor, guided by domain context, to induce entities and relations. Typical deployments choose one of these paths depending on which source type is available.
Dual-Path Construction Strategy.
ATLAS supports two alternative construction paths for the UMM (Figure 2). In practice, deployments usually select one of these paths: either the domain already ships a machine-readable meta-model (Path S1), or it only has natural-language specifications with no canonical XSD/UML (Path S2). Both paths produce the same kind of output: a typed graph $M_{\text{UMM}}$ with provenance annotations. We describe the two paths below.
Path S1: Structured Meta-Model Transformation.
When an explicit domain meta-model exists, ATLAS reuses an established dual-stage transformation process (DualStageFramework). This process proceeds as follows:
  1. (1)
    Structural Extraction. XMI files are parsed to recover inheritance DAGs via topological sorting. XSD schemas are parsed to recover static data constraints such as field cardinalities and type restrictions. The result is a set of class/attribute definitions and their structural relationships.
  2. (2)
    Semantic Fusion. DP-Fusion (Dual-Path Fusion) merges the class hierarchy derived from XMI with the data-level constraints extracted from XSD. Conflicting field definitions are resolved by priority rules (e.g., XSD constraints override inherited defaults), yielding a single consistent view of entities, attributes, and associations.
  3. (3)
    Behavioral Annotation. Domain annotations (including stereotypes and association semantics) are carried forward as behavioral hooks: for example, reference-lookup patterns or CRUD-style method signatures that indicate how instances of a class are expected to interact.
Path S1 therefore produces a UMM instance in which: (i) entities, attributes, and relations reflect the authoritative schema; and (ii) selected behavioral expectations are explicitly attached to those entities. Within ATLAS, this serves as the canonical semantic backbone whenever the domain already provides a formalized meta-model.
Path S2: LLM-Assisted Induction from Unstructured Sources.
Some domains lack any clean XSD/XMI-style schema. Instead, what counts as a “component,” a “reference,” or an “obligation” is only defined informally in prose. Path S2 lifts those informal descriptions into the same UMM representation through a three-stage pipeline:
Stage 1: Context-Enriched Document Segmentation. A corpus $D=\{d_{1},\ldots,d_{n}\}$ of natural-language specifications is segmented by a learned boundary detector $B$, which identifies sections likely to define entities, attributes, or relationships. For each segment $s_{i}$, ATLAS injects domain context into the prompt (e.g., known entity names, typical field names, role labels from prior segments or from a seed ontology). This grounding step constrains the extractor’s vocabulary and reduces hallucinated concepts.
Stage 2: LLM-Based Entity and Relation Induction. An LLM extractor $E_{\theta}$, parameterized by domain knowledge $\theta$, maps each context-enriched segment to a bundle of candidate model elements:
(1) $E_{\theta}:\text{Segment}\rightarrow\mathcal{C}_{\text{raw}}\times[0,1],$
where $\mathcal{C}_{\text{raw}}$ is a set of proposed entities, attributes, and relations, and the confidence score in $[0,1]$ is used to discard low-certainty candidates. All candidates follow a fixed induction schema that records: a proposed identifier (name, ID), its intended role or parent entity, its described properties or fields in natural language, and provenance metadata (document reference, section index, confidence).
Stage 3: Consolidation into the UMM. The induced candidates are reconciled against the evolving $M_{\text{UMM}}$. We apply fuzzy alignment to associate each candidate with an existing UMM node when possible:
(2) $\alpha:\mathcal{C}_{\text{raw}}\rightharpoonup V_{\text{UMM}},$
where $\alpha(c)=\bot$ indicates that $c$ does not match any known node. If $\alpha(c)\neq\bot$, the candidate is treated as an alias or refinement of the matched node, and its provenance metadata is attached to that node. If $\alpha(c)=\bot$, ATLAS tentatively adds a new node or relation into $M_{\text{UMM}}$, tagged with provenance and marked for optional human review when the stakes are high (e.g., safety-critical components). Crucially, at this stage we are only inducing and reconciling domain entities, attributes, and relations; the extraction of constraints are explicitly deferred to the Integrated Constraint Model (ICM) construction in Section 3.3.
Path S2 thus yields a UMM even in domains with no formal schema at all. The result is again a typed, provenance-aware graph $M_{\text{UMM}}$ that defines what “objects” exist, how they relate, and which textual sources justified their inclusion.
Scalability Note.
Although UMM construction is primarily an integration activity rather than a new theoretical contribution, its runtime characteristics matter for practical deployment. By utilizing union-find with path compression to merge candidate-equivalent nodes, the reconciliation procedure runs in near-linear time relative to the total number of entities across all sources (effectively $\mathcal{O}(\sum|V_{i}|)$). This efficiency allows ATLAS to ingest industrial-scale meta-models with thousands of entities without incurring the performance bottlenecks typical of manual pairwise weaving.
Relation to Existing Approaches.
Classical model weaving frameworks (e.g., AMW) rely on manually authored correspondence models, while graph transformation systems (e.g., AGG) require hand-written rules for every pair of source models. Purely LLM-driven extraction has also been explored (chenAutomatedDomainModeling2023), but without grounding the induced concepts in a persistent, provenance-tracked ontology, such approaches risk hallucinated entities and silently inconsistent references. ATLAS differs in three respects:
  • Source Adaptability. Path S1 ingests explicit XSD/XMI-style meta-models through deterministic model transformations (DualStageFramework). Path S2 induces a meta-model from prose using an LLM guided by domain context. Deployers choose whichever path matches their domain reality.
  • Transformation Nature (Fusion vs. Translation). Classical MDE transformations typically strive for bijectivity (or isomorphism) to ensure round-trip engineering capabilities (DualStageFramework; Stevens2010), requiring strict one-to-one mappings between source and target. However, in Path S2, the input (natural language) is inherently redundant, fragmented, and ambiguous, while the target (UMM) must be concise and unified. Consequently, adherence to bijectivity is theoretically ill-suited. ATLAS treats this process not as a translation ($A\leftrightarrow B$) but as a fusion via surjective homomorphisms (formalized by colimits). This shift allows multiple partial textual descriptions to “collapse” into a single semantic entity, rigorously resolving the redundancy and fuzzy alignments that traditional bijective transformations cannot handle.
  • Semantic Preservation. In both paths, newly discovered or merged entities are either aligned to an existing node in $M_{\text{UMM}}$ via $\alpha$ or added as a new node annotated with provenance and flagged for optional review. This prevents semantic drift and keeps the UMM auditable.
  • Scalability. Because reconciliation proceeds via incremental union rather than exhaustive pairwise weaving, ATLAS scales to realistic industrial domains, as detailed in the Scalability Note (Section 3.2).
In summary, UMM construction is the step that turns messy, heterogeneous domain knowledge into a single, provenance-aware semantic backbone. All later phases of ATLAS—constraint modeling, generation-time enforcement, post-generation validation, and audit-guided repair—assume this backbone as the shared point of truth.
3.3. Constraint Extraction and Integrated Constraint Model (ICM)
Role of the ICM.
The Unified Meta-Model (UMM) provides ATLAS with a single, auditable semantic backbone. The Integrated Constraint Model (ICM) builds on that backbone by collecting, normalizing, and compiling the domain rules that generated artifacts must satisfy. These rules range from low-level structural requirements (e.g., mandatory fields, cardinalities) to high-level semantic and logical invariants (e.g., “an OperationInvokedEvent must reference an existing operation”), and they originate from both machine-readable meta-models and natural-language specifications. This subsection formalizes how such constraints are represented, how they are extracted through two complementary channels, how they are grounded in the UMM, and how they are compiled into executable Layer-1 (L1) and Layer-2 (L2) validators for generation-time enforcement and post-generation verification.
Formalization of Constraint Spaces.
To reason uniformly about constraints from heterogeneous sources, we model them in a typed semantic space.
Definition 3.0 (Constraint Space).
A constraint space $\mathcal{C}=(\mathcal{L},\mathcal{S},\models,\leq)$ consists of:
  • $\mathcal{L}$: a logical language for expressing constraints,
  • $\mathcal{S}$: a semantic domain of valid artifact structures,
  • $\models$: a satisfaction relation $\mathcal{S}\times\mathcal{L}\rightarrow\{\top,\bot\}$,
  • $\leq$: an ordering on $\mathcal{L}$ capturing relative constraint strength.
Different inputs to ATLAS (e.g., AUTOSAR XSD/XMI schemas vs. regulatory PDFs) induce different constraint spaces. Before integration, these must be semantically aligned with the UMM so that downstream validators can interpret them consistently.
Definition 3.0 (Partial Alignment Mapping).
Let $\mathcal{P}$ be the set of extracted constraint patterns and let $U$ be the set of UMM sorts. A partial alignment is a (possibly partial) mapping
(3) $\alpha:\mathcal{P}\rightharpoonup U$
where $\alpha(p)=\bot$ denotes that pattern $p$ has no direct anchor in $U$ and must instead be retained at an abstract layer $\mathcal{C}_{\mathrm{abs}}$.
Intuitively, $\alpha$ tells us which extracted rules are already grounded in the UMM and which must be quarantined for review or additional interpretation.
Dual-Channel Constraint Acquisition.
ICM population proceeds through two coordinated channels, targeting persistent, reusable domain constraints. Both channels produce structured intermediate representations that are compiled into executable validators and stored with provenance. The channels are:
Channel 1: Deductive Extraction from Structured Meta-Models. Given a source meta-model $M=(V,E,T,\mathcal{C})$ as defined in Section 3.2, we apply programmatic operators that extract constraints directly from the schema:
$\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{struct}}$ $:M\to 2^{\mathcal{C}_{\mathrm{struct}}},$
$\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{sem}}$ $:M\to 2^{\mathcal{C}_{\mathrm{sem}}},$
(4) $\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{log}}$ $:M\to 2^{\mathcal{C}_{\mathrm{log}}}.$
Here, $\mathcal{C}_{\mathrm{struct}}$ captures structural invariants (cardinality, mandatory fields, inheritance restrictions), $\mathcal{C}_{\mathrm{sem}}$ captures semantic constraints (reference integrity, role consistency), and $\mathcal{C}_{\mathrm{log}}$ captures logical/temporal/numeric constraints.
L2 Validator Instantiation and Multi-Modal Enforcement.
The abstract semantic and logic validators $\mathcal{V}_{\mathrm{sem}}$ and $\mathcal{V}_{\mathrm{log}}$ in $\mathcal{I}$ support a multi-modal instantiation strategy to balance theoretical rigor with engineering scalability. In this work, we focus on the formal modality, instantiating semantic validators using SHACL (Shapes Constraint Language) (knublauchSHACLShapesConstraint2017) and logic validators using SMT (Satisfiability Modulo Theories) solvers (demouraZ3EfficientSMT2008). SHACL naturally captures cross-reference semantics in XML/RDF-based artifacts, while SMT efficiently handles bounded numeric reasoning and temporal constraints. This choice prioritizes machine-checkable auditability (“Proof-Carrying”) suitable for safety certification.
However, the architecture also supports a programmatic modality for industrial-scale deployment. By leveraging the strongly-typed model APIs generated via Path S1 (Section 3.2), $\mathcal{C}_{\mathrm{sem}}$ can be compiled into executable validation scripts (e.g., Java/Python assertions). To ensure these scripts are syntactically and semantically valid, we employ type-aware constrained decoding: the LLM’s generation of validation code is strictly masked by the symbol tables of the S2D2-generated APIs. This prevents the model from hallucinating non-existent method calls, effectively creating a “compiler-in-the-loop” generation process. While formal validators provide mathematical completeness for critical safety properties, these verifiable programmatic validators offer superior performance for checking thousands of routine business rules. Regardless of the modality, the key architectural requirement is that validators produce machine-readable certificates $\pi_{\text{sem}}$ and $\pi_{\text{logic}}$ with precise violation localization to drive Audit-Guided Repair.
These extracted constraints are then compiled for enforcement at appropriate layers:
(5) $\mathrm{Compile}_{\mathrm{L1}}$ $:\mathcal{C}_{\mathrm{struct}}\to\{\text{DFA/GBNF, JSON Schema}\},$
$\mathrm{Compile}^{\mathrm{sem}}_{\mathrm{L2}}$ $:\mathcal{C}_{\mathrm{sem}}\to\mathcal{V}_{\mathrm{sem}},$
(6) $\mathrm{Compile}^{\mathrm{log}}_{\mathrm{L2}}$ $:\mathcal{C}_{\mathrm{log}}\to\mathcal{V}_{\mathrm{log}}.$
where $\mathcal{V}_{\mathrm{sem}}$ and $\mathcal{V}_{\mathrm{log}}$ denote the spaces of semantic and logic validators, respectively. In our implementation, these are instantiated as SHACL shapes and SMT-LIB2 formulas (Section 3.3). The result is that Channel 1 provides deterministic constraints with explicit provenance in the original schema, and these constraints become part of the persistent ICM.
Channel 2: Inductive Extraction from Natural-Language Specifications. Many industrial domains encode mission-critical rules in prose rather than structured schemas: AUTOSAR PDFs, regulatory guidance, safety standards, legal jurisdiction rules. Channel 2 uses an LLM-based extractor to convert free-form text into structured constraint candidates, following prior work on LLMs as structured information extractors (Dagdelen2024; Lu2022; Wei2020; Jiao2024). Formally,
$\mathrm{Extract}^{\mathrm{LLM}}_{\mathrm{sem}}$ $:\mathcal{D}\to 2^{\mathcal{C}_{\mathrm{sem}}^{\text{structured}}},$
(7) $\mathrm{Extract}^{\mathrm{LLM}}_{\mathrm{log}}$ $:\mathcal{D}\to 2^{\mathcal{C}_{\mathrm{log}}^{\text{structured}}}.$
These outputs are intermediate JSON-like specifications that undergo entity linking $\alpha$ (Definition 3), semantic compatibility validation (Definition 5), and compilation to L2 validators before admission to the ICM. This prevents unconstrained LLM extraction from directly polluting the validator set.
Figure 3. LLM-based extraction from natural-language specifications. The extractor proposes structured constraint candidates and aligns each candidate to entities in the UMM. Only candidates that align cleanly and pass semantic checks are committed to the Integrated Constraint Model (ICM); the rest are flagged for human review instead of being turned directly into validators.
Illustration of Channel 2 extraction with semantic validation. Specification text is processed by an LLM-based extractor, which proposes structured JSON-like constraint candidates and links each candidate to entities in the Unified Meta-Model (UMM). Only candidates that can be anchored via alpha and that pass semantic compatibility checks are inserted into the Integrated Constraint Model (ICM). Candidates that fail these checks are not compiled directly into L2 validators but are instead flagged for review.
Layer Assignment: L1 vs. L2.
Channel 1 feeds $\mathcal{C}_{\mathrm{struct}}$ to L1, where constraints determinize to prefix-closed automata (DFA/GBNF, JSON Schema) for token-by-token masking during decoding. Channel 2 contributes only to L2, as free-form constraints may require global context unsuitable for incremental enforcement. Both channels populate $\mathcal{C}_{\mathrm{sem}}$ and $\mathcal{C}_{\mathrm{log}}$ in L2, compiled to semantic and logic validators applied post-generation. This separation preserves deterministic safety during generation while enabling rich semantic/logic checks afterward.
ICM as Persistent Knowledge.
We now formalize how both channels populate a persistent, auditable repository of constraints.
Figure 4. ICM as a multi-sorted algebra layered over the UMM. White nodes denote UMM entities; gray nodes encode semantic and logical constraints extracted from Channel 1 (deductive) and Channel 2 (LLM-assisted). Morphisms relate abstract constraints to class-/attribute-level constraints, enabling cross-layer reasoning and repair prioritization.
Diagram of the Integrated Constraint Model (ICM) represented as a multi-sorted algebra on top of the Unified Meta-Model (UMM). White nodes correspond to UMM entities. Gray nodes capture semantic and logical constraints obtained from Channel 1 (deductive extraction) and Channel 2 (LLM-assisted extraction). Morphisms connect abstract constraints to class- and attribute-level constraints, supporting cross-layer reasoning and guiding which repairs should be prioritized.
Definition 3.0 (Integrated Constraint Model).
An ICM is a tuple $\mathcal{I}=(\mathcal{U},\mathcal{X},\mathcal{F},\Gamma)$ where:
  • $\mathcal{U}$: the underlying UMM providing sorts and signatures;
  • where $\mathcal{X}=\{\mathcal{C}_{\mathrm{struct}},\mathcal{C}_{\mathrm{sem}},\mathcal{C}_{\mathrm{log}}\}$: constraint algebras populated by Channel 1 and Channel 2;
  • $\mathcal{F}=\{f_{ij}:\mathcal{C}_{i}\rightarrow\mathcal{C}_{j}\}$: constraint morphisms supporting cross-layer reasoning;
  • $\Gamma$: a dependency lattice with partial order $\preceq$, which we later use to prioritize repair (Section 3.6).
The morphisms $\mathcal{F}$ capture how abstract invariants (e.g., “all mode transitions are acyclic”) induce obligations on concrete fields and references. The lattice $\Gamma$ then orders those obligations, which is later consumed by Audit-Guided Repair (AGR).
ICM Construction via Dual-Channel Extraction.
Algorithm 1 summarizes how ATLAS assembles the ICM from both structured schemas and unstructured documents. This process is persistent: it populates the long-lived constraint repository (the ICM $\mathcal{I}$), rather than just synthesizing a one-off validator.
Algorithm 1 ICM Construction via Dual-Channel Extraction
1:Meta-model $M=(V_{M},E_{M},T_{M},\mathcal{C}_{M})$, UMM graph $(V_{\text{UMM}},E_{\text{UMM}})$, Spec docs $D=\{d_{1},\ldots,d_{n}\}$, Constraint schema $\mathcal{S}_{c}$
2:ICM graph $(V_{\text{ICM}},E_{\text{ICM}})$
3:function BuildICM($M$, $V_{\text{UMM}}$, $D$, $\mathcal{S}_{c}$)
4:  $V_{\text{ICM}}\leftarrow\emptyset$; $E_{\text{ICM}}\leftarrow\emptyset$
5:  for $\phi\in\mathcal{C}_{M}$ do $\triangleright$ Channel 1: Deductive extraction
6:    if $\phi\in\mathcal{C}_{\mathrm{sem}}\cup\mathcal{C}_{\mathrm{log}}$ then
7:     $c\leftarrow\text{createConstraintNode}(\phi)$
8:     add $c$ to $V_{\text{ICM}}$ and link to $V_{\text{UMM}}$ via $E_{\text{ICM}}$
9:    end if
10:  end for
11:  for doc $d\in D$ do $\triangleright$ Channel 2: LLM-assisted extraction
12:    $d^{\prime}\leftarrow\text{injectContext}(d,V_{\text{UMM}})$
13:    $\mathcal{C}_{\text{raw}}\leftarrow\text{extractStructured}(d^{\prime},V_{\text{UMM}},\mathcal{S}_{c})$
14:    for $c_{\text{raw}}\in\mathcal{C}_{\text{raw}}$ do
15:     $\alpha(c_{\text{raw}})\leftarrow\text{linkToUMM}(c_{\text{raw}},V_{\text{UMM}})$
16:     if $\alpha(c_{\text{raw}})=\bot$ then
17:      flag for human review; continue
18:     end if
19:     if validateSemanticCompatibility($c_{\text{raw}}$, $M$) then
20:      add $c_{\text{raw}}$ to $V_{\text{ICM}}$ with anchor $\alpha(c_{\text{raw}})$
21:     else
22:      repair_and_relink($c_{\text{raw}}$); repeat
23:     end if
24:    end for
25:  end for
26:  return $(V_{\text{ICM}},E_{\text{ICM}})$
27:end function
The central safeguard is that Channel 2 output does not automatically enter the persistent ICM $\mathcal{I}$ or any downstream validator. Each candidate constraint must (i) be anchored to a concrete UMM entity via $\alpha$, and (ii) pass semantic compatibility checks before being admitted.
LLM-Assisted Constraint Extraction with Entity Grounding.
Channel 1 recovers explicit constraints from structured schemas. Channel 2 is responsible for implicit constraints described only in prose, such as: “an OperationInvokedEvent must reference an existing operation” or “mode transitions must form a directed acyclic graph.” Algorithm 2 formalizes this pipeline.
Algorithm 2 Channel 2: LLM-Assisted Constraint Extraction for ICM
1:Document segment $s$, UMM entities $V_{\text{UMM}}$, constraint schema $\mathcal{S}_{c}$
2:Structured constraint candidates $\mathcal{C}^{\text{cand}}=\{c_{1}^{\text{cand}},\ldots,c_{n}^{\text{cand}}\}$
3:function ExtractStructuredConstraints($s$, $V_{\text{UMM}}$, $\mathcal{S}_{c}$)
4:  $\text{ctx}\leftarrow\text{serializeUMM}(V_{\text{UMM}})$
5:  $\mathcal{C}_{\text{raw}}\leftarrow\text{LLM}(\text{buildPrompt}(s,\text{ctx},\mathcal{S}_{c}),T{=}0.3)$
6:  $\mathcal{C}^{\text{cand}}\leftarrow\emptyset$
7:  for $c_{\text{raw}}\in\mathcal{C}_{\text{raw}}$ do
8:    $(v,\text{score})\leftarrow\text{fuzzyMatch}(c_{\text{raw}}.\text{target},V_{\text{UMM}})$
9:    if $v=\bot$ or $\text{score}<\theta_{\text{link}}$ then
10:     log_for_review($c_{\text{raw}}$, $v$, score); continue
11:    end if
12:    $c^{\text{cand}}\leftarrow\text{createStructuredForm}(c_{\text{raw}},v)$
13:    $c^{\text{cand}}.\text{type}\leftarrow\text{classifyConstraint}(c_{\text{raw}})$
14:    $c^{\text{cand}}.\text{anchor}\leftarrow v$
15:    $c^{\text{cand}}.\text{metadata}\leftarrow\langle\text{doc\_id},\text{para},c_{\text{raw}}.\text{conf},\text{score}\rangle$
16:    if validateSemanticCompatibility($c^{\text{cand}}$, $V_{\text{UMM}}$) then
17:     add $c^{\text{cand}}$ to $\mathcal{C}^{\text{cand}}$
18:    else
19:     log_for_review($c^{\text{cand}}$, reason=”incompatible”)
20:    end if
21:  end for
22:  return $\mathcal{C}^{\text{cand}}$
23:end function
The key mechanism is Step 2: entity grounding. Instead of accepting every LLM hypothesis, we force each proposed constraint to attach to an existing UMM node with confidence above threshold $\theta_{\text{link}}$. This (i) filters hallucinated entities and broken references, and (ii) creates a traceable link from each natural-language rule to a formal constraint and, through the UMM, to the certified domain ontology. Constraints that fail linking or compatibility are logged for expert review rather than silently admitted into the ICM.
Semantic Compatibility and Extraction Consistency.
To ensure that accepted constraints do not contradict the UMM, we require semantic compatibility.
Definition 3.0 (Semantic Compatibility).
Let $\mathcal{S}$ be the semantic domain of admissible artifacts from Definition 2, and let $\models$ be the satisfaction relation. A constraint $c$ is semantically compatible with a meta-model $M$ if and only if
(8) $\exists a\in\mathcal{S}:(a\models M)\land(a\models c).$
Intuitively, $c$ is compatible with $M$ if there exists at least one admissible artifact $a$ that can satisfy both $M$ and $c$ simultaneously, i.e., $c$ does not impose an impossible obligation relative to the domain described by $M$.
We now relate Channel 1 (schema-driven) and Channel 2 (LLM-assisted) extraction. For brevity, define:
$\mathrm{Extract}^{\mathrm{KG}}(M)\;\;=\;\;\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{struct}}(M)\;\cup\;\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{sem}}(M)\;\cup\;\mathrm{Extract}^{\mathrm{KG}}_{\mathrm{log}}(M),$
and
$\mathrm{Extract}^{\mathrm{LLM}}(D)\;\;=\;\;\mathrm{Extract}^{\mathrm{LLM}}_{\mathrm{sem}}(D)\;\cup\;\mathrm{Extract}^{\mathrm{LLM}}_{\mathrm{log}}(D),$
where $D$ is the specification corpus.
Theorem 6 (Extraction Consistency with Partial Alignment).
For any valid specification corpus $D$ and its corresponding meta-model $M$, the extracted constraint sets satisfy
(9) $\bigl(\mathrm{Extract}^{\mathrm{KG}}(M)\cap\mathrm{Extract}^{\mathrm{LLM}}(D)\bigr)\;\cup\;\mathcal{C}_{\mathrm{abs}}\;\neq\;\varnothing,$
and every constraint $c\in\mathrm{Extract}^{\mathrm{LLM}}(D)$ obeys
(10) $\alpha(c)\neq\bot\;\Longrightarrow\;\bigl(\alpha(c)\in U\ \land\ c\text{ is semantically compatible with }M\bigr),$
where $\alpha$ is the partial alignment map (Definition 3), $U$ is the set of UMM sorts, and $\mathcal{C}_{\mathrm{abs}}$ is the quarantined abstract layer for rules that have no direct UMM anchor.
In words: either a Channel 2 rule anchors to a UMM entity with acceptable compatibility, or it is quarantined in $\mathcal{C}_{\mathrm{abs}}$ rather than being silently admitted into the ICM.
Proof sketch.
Let $\mathcal{T}:\mathcal{P}\rightarrow U\cup\{\bot\}$ be the pattern-to-model mapping. For any pattern $p\in D$ with $\mathcal{T}(p)=u\neq\bot$, $u\in U$ serves as an anchor establishing overlap with $\mathrm{Extract}_{\mathrm{ded}}(M)$. Patterns with $\mathcal{T}(p)=\bot$ are routed to $\mathcal{C}_{\mathrm{abs}}$ by Definition 3. Thus the combined set is non-empty, and compatibility follows from how $E_{\theta}$ is constructed to respect $M$. ∎
The theorem formalizes the discipline imposed by partial alignment $\alpha$: either a constraint grounds in the UMM and is provably compatible, or it is quarantined (in $\mathcal{C}_{\mathrm{abs}}$) for review instead of silently entering the validator pipeline.
Dynamic Constraint Synthesis (Runtime, Non-ICM).
In addition to the static ICM, ATLAS supports request-specific constraints that apply only to a single generation run. For example, a user generating one AUTOSAR subsystem may require “all port connections in this subsystem must use compatible data types” without permanently altering the global repository. To accommodate such ad hoc requirements, we synthesize ephemeral L2 validators at runtime:
(11) $\mathrm{Synthesize}^{\mathrm{NLP}}:\mathcal{R}_{\text{dynamic}}\to\mathcal{V}_{\mathrm{L2}}.$
These validators (instantiated as SHACL or SMT in our implementation) are fed directly into L2 verification during CVG (Section 3.4), but they are not persisted in the ICM unless they pass solver-backed validation and are deemed reusable.
Algorithm 3 describes this guarded synthesis loop.
Algorithm 3 Dynamic NLP Constraint Synthesis (CVG Runtime)
1:Per-request NL requirement $r\in\mathcal{R}_{\text{dynamic}}$, UMM context $V_{\text{UMM}}$
2:Executable L2 validators $\mathcal{V}_{\text{L2}}=\{\mathcal{V}_{\text{sem}},\mathcal{V}_{\text{log}}\}$ (instantiated as SHACL and SMT)
3:function SynthesizeDynamicConstraints($r$, $V_{\text{UMM}}$)
4:  $\text{ctx}\leftarrow\text{selectRelevantEntities}(r,V_{\text{UMM}})$
5:  $\mathcal{C}_{\text{formal}}\leftarrow\text{LLM}(\text{buildSynthesisPrompt}(r,\text{ctx}),T{=}0.2)$
6:  $\mathcal{V}_{\text{SHACL}}\leftarrow\emptyset$; $\mathcal{V}_{\text{SMT}}\leftarrow\emptyset$
7:  for $c\in\mathcal{C}_{\text{formal}}$ do
8:    if $c.\text{type}=\text{``semantic''}$ then
9:     add $\text{compileToSHACL}(c,V_{\text{UMM}})$ to $\mathcal{V}_{\text{SHACL}}$
10:    else if $c.\text{type}=\text{``logic''}$ then
11:     add $\text{compileToSMT}(c)$ to $\mathcal{V}_{\text{SMT}}$
12:    end if
13:  end for
14:  $\text{conflicts}\leftarrow\text{detectConflicts}(\mathcal{V}_{\text{SHACL}}\cup\mathcal{V}_{\text{SMT}},r)$
15:  if $\text{conflicts}\neq\emptyset$ then
16:    resolve_via_relaxation(conflicts)
17:  end if
18:  return $\mathcal{V}_{\text{L2}}$
19:end function
Conceptually, Algorithms 1 and 2 populate the long-lived constraint repository (ICM) that applies across generation tasks. Algorithm 3 instead creates validators that exist only for the current request, unless they are later promoted (via solver-backed validation) into the persistent Formal Constraints Database. This policy prevents one-off requirements from polluting the global constraint store or contradicting established domain invariants.
Compilation to Executable Validators.
Finally, constraints admitted into the ICM must be turned into actual checkers used by CVG. Figure 5 illustrates ATLAS’s stratified compilation pipeline.
Figure 5. Stratified compilation with LLM fallback and repository policy. Left: Channel 1 programmatic mapping compiles UMM/ICM constraints to L1 validators (JSON Schema, DFA, GBNF) and L2 validators (SHACL, SMT), preserving semantics. Right: when programmatic mapping is incomplete, an LLM-based pipeline (maSpecGenAutomatedGeneration2025) synthesizes missing constraints with solver-backed validation; only verified outputs are persisted in the ICM storage.
Two-panel diagram of stratified compilation with LLM fallback and repository policy. Left panel: Channel 1 performs programmatic mapping from UMM/ICM constraints into L1, which enforces constraints at generation time using JSON Schema, FSM, GBNF, Regex, or DFA, and into L2, which enforces constraints post-generation via SHACL and SMT, while preserving intended semantics. Right panel: when the programmatic mapping is incomplete, a SpecGen-style pipeline invokes an LLM with solver-backed validation to synthesize missing SHACL/SMT constraints. Only constraints that pass validation are written into the Formal Constraints Database (ICM storage).
Following (barrett2009satisfiability; w3c2017shacl), we define compilation functions that map abstract constraints to executable enforcement mechanisms:
(12) $\kappa_{\mathrm{lex}}$ $:\mathcal{C}_{\mathrm{lex}}\rightarrow\mathcal{G}_{\mathrm{CFG}}$
(13) $\kappa_{\mathrm{struct}}$ $:\mathcal{C}_{\mathrm{struct}}\rightarrow\mathcal{A}_{\mathrm{prefix}}$
(14) $\kappa_{\mathrm{sem}}$ $:\mathcal{C}_{\mathrm{sem}}\rightarrow\mathcal{V}_{\mathrm{sem}}$
(15) $\kappa_{\mathrm{logic}}$ $:\mathcal{C}_{\mathrm{logic}}\rightarrow\mathcal{V}_{\mathrm{log}}$
MDE Interpretation: Compilation as Model Transformation.
Building upon the extraction methodology established in our S2D2 framework (DualStageFramework), which utilizes QVT-R to transform heterogeneous meta-models (XMI/XSD) into the UMM, ATLAS extends this chain by treating the compilation functions $\kappa$ as a subsequent set of vertical model transformations. Specifically, as visualized in Figure 6, $\kappa_{\mathrm{struct}}$ functions as a deterministic transformation rule that maps the static schema topology (from the ICM) into a dynamic state-transition system ($\mathcal{A}_{\mathrm{prefix}}$). Unlike the probabilistic nature of LLM token sampling, this transformation is mathematically rigorous and semantic-preserving (as supported by grammar-constrained decoding theory (scholakPICARDParsingIncrementally2021; poesiaSynchromeshReliableCode2021)). This distinction allows ATLAS to claim that Layer-1 enforcement is not merely a heuristic “guardrail,” but a correct-by-construction synthesis of the structural grammar, leaving only semantic ambiguities to the stochastic capabilities of the model.
Figure 6. Mapping the ATLAS constraint compilation pipeline to MDA concepts. The ICM functions as a Platform-Independent Model (PIM) of domain rules, which is deterministically transformed via $\kappa$ functions (Transformation Rules) into Platform-Specific Models (PSM) represented by L1 Prefix Automata and L2 Validators.
Remark. In our implementation, the context-free grammar $\mathcal{G}_{\mathrm{CFG}}$ is realized as GBNF/CFG fragments enforced via PDA/LR-style controllers. The prefix automaton $\mathcal{A}_{\mathrm{prefix}}$ is realized as a DFA (for regular constraints) or a bounded unfolding of the PDA (for recursive structures) to enable unified token masking.
Theorem 7 (Conditional Soundness and Completeness).
Fix an unfolding bound $d$ for recursive structures. Let $\kappa_{\text{struct}}$ map to prefix automata $\mathcal{A}_{\text{prefix}}$ (e.g., DFA), $\kappa_{\text{sem}}$ map to semantic validators $\mathcal{V}_{\text{sem}}$ (e.g., SHACL), and $\kappa_{\text{logic}}$ map to logic validators $\mathcal{V}_{\text{log}}$ (e.g., SMT). Write $\kappa$ for their union on $\mathcal{C}$ with cross-layer deduplication. Then:
  1. (i)
    Soundness. For all $c\in\mathcal{C}$ and $a\in\mathbb{A}$:  $a\models c\Rightarrow a\models\kappa(c)$.
  2. (ii)
    Conditional completeness. For all $c\in\mathcal{C}$ and $a\in\mathbb{A}$ with $\text{maxDepth}(a)\leq d$:  $a\models\kappa(c)\Rightarrow a\models c$.
Proof sketch.
(i) By construction: DFA acceptance implies satisfaction of the structural fragment; Semantic validators encode graph-structural constraints (e.g., SHACL shapes for first-order relations); logic validators capture temporal and numeric properties (e.g., SMT formulas for bounded arithmetic). Each $\kappa_{\bullet}$ is a semantics-preserving encoding.
(ii) For recursive structures, finite unfolding to depth $d$ yields an automaton whose accepted language over-approximates the target up to depth $d$. Any artifact $a$ with $\text{maxDepth}(a)\leq d$ that satisfies $\kappa(c)$ admits a witness (accepting run / model) reconstructing a proof that $a\models c$. Cross-layer deduplication ensures that constraints assigned to the most specific executable layer remain entailed in stronger layers. ∎
Together, these results guarantee that constraints admitted into the ICM can be enforced in two complementary ways: L1 constraints prevent structurally invalid outputs during generation, and L2 validators certify semantic and logical correctness after generation. The evidence produced by these validators is later combined with audit trails to drive Audit-Guided Repair (Section 3.6).
3.4. Unified Automaton Execution with Theoretical Guarantees
Role in CVG.
Constraint-Guided Verifiable Generation (CVG) enforces Layer-1 (L1) structural constraints during decoding and Layer-2 (L2) semantic/logic constraints after decoding (Section 3.3). This subsection explains how L1 constraints are compiled into a unified automaton that steers the Large Language Model (LLM) token-by-token, how this execution is audited, and why L1 and L2 must remain distinct from a computability perspective. We further show how these mechanisms provide prefix safety, bounded structural repair, and machine-checkable evidence for certification and Audit-Guided Repair (AGR) (Section 3.6).
Figure 7. Unified automaton execution for L1 enforcement during LLM decoding. Input constraints: L1 structural constraints ($\mathcal{C}_{\mathrm{struct}}$ from the ICM; Section 3.3) are compiled into executable automata. Finite-state fragments (JSON Schema / Regex / FSM) are determinized to prefix-closed DFAs. GBNF/CFG fragments are executed via PDA/LR; for unified masking and logging, bounded-depth unfolding (depth $d$) produces an equivalent DFA $\mathcal{A}_{d}$. Runtime decoding loop: At decoding step $t$, the executor computes $M_{t}=\{y\in\Sigma\mid\delta(s_{t},y)\neq\bot\}$, masks invalid tokens in the LLM logits, samples a valid token $y_{t}$, and transitions to $s_{t+1}=\delta(s_{t},y_{t})$. Audit trail: Each step records $\tau_{t}=\langle s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t\rangle$ into the Evidence Registry. Post-generation, L2 validators (semantic checkers for graph constraints and logic checkers for temporal/numeric properties) certify the completed artifact, yielding composite evidence $\Pi=(\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}})\oplus\tau$ (Section 3.5).
Diagram of unified automaton execution for Layer 1 (L1) enforcement during large language model decoding. First, structural constraints from the Integrated Constraint Model (ICM) are compiled into executable automata: JSON Schema, Regex, and FSM fragments become deterministic finite automata, while GBNF and CFG fragments are executed via PDA/LR and optionally unfolded to bounded-depth DFAs for unified masking and logging. During decoding at each step t, the executor computes the set of valid next tokens, masks invalid tokens in the model logits, samples a valid token, and advances the automaton state. Each decoding step logs an audit tuple including states, mask size, chosen token, and timing. After generation finishes, Layer 2 (L2) validators for semantic constraints (e.g., SHACL) and logical constraints (e.g., SMT) and SMT (logical and temporal constraints) certify the final artifact and combine the audit trail into composite evidence.
Why Layering is Necessary: Decidability and Auditability.
Existing grammar-aligned decoding methods enforce a single global grammar or schema during LLM decoding and can align external grammars with subword vocabularies, using speculative decoding to keep runtime overhead negligible while preserving grammar validity (scholakPICARDParsingIncrementally2021; parkGrammaralignedDecoding2025; deljouyiUnderstandableTestGeneration2024; zotero-item-1597; Park2025; Lu2025b). However, large-scale evaluations (e.g., SchemaBench / JSONSchemaBench) report two persistent failure modes: (i) over-constraining, where grammars block structurally unusual but semantically valid outputs, and (ii) under-constraining, where syntactically valid outputs still violate cross-reference integrity, safety rules, or timing constraints (zotero-item-1624). The root cause is computability: not all constraints are prefix-decidable.
Layer-1 constraints (L1). Structural and lexical invariants—“element A must contain child B,” “attribute X must be an integer,” “balanced tags,” etc.—are local and prefix-decidable. Their satisfiability at step $t$ depends only on the partially generated artifact plus bounded lookahead. Such constraints can be compiled into finite automata (DFA/PDA) that, at each decoding step, expose the set of tokens that keep the output on a valid path. They are therefore enforceable during generation.
Layer-2 constraints (L2). Semantic and logic invariants—“every OperationInvokedEvent must reference a defined Operation,” “mode transitions must form an acyclic graph,” “all port connections must be type-compatible”—are global. Their satisfiability depends on full-artifact knowledge (cross-file symbol tables, temporal relations, numerical bounds). Deciding these properties while decoding is infeasible because it would require reasoning about content not yet generated. They must therefore be checked after generation using semantic and logic validators (Section 3.3).
This is not an engineering convenience but a theoretical split. L1 maximizes what can be proven safe online, and L2 certifies global semantics offline. The separation underpins ATLAS’s claim of constraint-guided, auditable generation in safety-critical domains.
Definition 3.0 (Constraint Enforcement Hierarchy).
Let $\mathcal{H}$ be a family of enforcement strategies, ordered by $\preceq$. For $h_{i},h_{j}\in\mathcal{H}$ we write $h_{i}\preceq h_{j}$ iff: (i) $\mathcal{L}(h_{i})\supseteq\mathcal{L}(h_{j})$, i.e., $h_{i}$ permits a superset of sequences accepted by $h_{j}$, and (ii) $\mathcal{P}(h_{i})\leq\mathcal{P}(h_{j})$, where $\mathcal{P}(\cdot)$ orders correctness guarantees. In ATLAS:
  • Layer-1 (Generation-Time Enforcement). Prefix-decidable structural constraints $\mathcal{C}_{\mathrm{struct}}$ are compiled into DFA/PDA automata for token-level masking. L1 enforces structural and lexical correctness through reachability-preserving transitions.
  • Layer-2 (Validation-Time Certification). Global semantic/logic constraints $\mathcal{C}_{\mathrm{sem}}$ and $\mathcal{C}_{\mathrm{log}}$ are compiled into L2 semantic and logic validators and run post-generation. L2 emits machine-checkable certificates for semantic and logical correctness.
Under this hierarchy, L1 has stronger online enforceability guarantees but weaker expressiveness; L2 has broader expressiveness but applies post hoc. Both are required for safety-critical assurance.
Unified Automaton Execution: Formal Semantics.
All L1 constraints are compiled (Section 3.3) into an automaton that provides generation-time guidance and an auditable trace. Let $s_{t}$ denote the automaton configuration (state or stack encoding) at decoding step $t$. We define the allowed token set by
(16) $M_{t}=\mathrm{Allow}(s_{t})=\{y\in\Sigma\mid\delta(s_{t},y)\neq\bot\},$
where $\Sigma$ is the token vocabulary and $\delta$ is the automaton transition function.
Compilation proceeds as follows:
  • Finite-state fragments (JSON Schema / Regex / FSM). These are determinized into a prefix-closed deterministic finite automaton (DFA) $\mathcal{A}=(Q,\Sigma,\delta,q_{0},F)$. For this pure DFA case, we write $q_{t}\in Q$ for the automaton state at decoding step $t$ (so here $s_{t}=q_{t}$), and every reachable non-final state $q_{t}\notin F$ still satisfies $\exists w:\delta^{*}(q_{t},w)\in F$ (accepting reachability).
  • Context-free fragments (GBNF / CFG). These are enforced via a PDA/LR-style automaton that maintains a parsing stack. For unified token masking and audit logging, we apply bounded-depth unfolding to depth $d$ (the same depth bound $d$ used in Theorem 7) to obtain an equivalent DFA $\mathcal{A}_{d}$ whose states encode stack configurations up to depth $d$. Depth $d$ trades off structural expressiveness, automaton size, and runtime cost.
The generation loop is then:
  1. 1.
    The executor queries $s_{t}$ to compute $M_{t}$.
  2. 2.
    The LLM produces logits $z_{t}\in\mathbb{R}^{|\Sigma|}$.
  3. 3.
    Tokens $y\notin M_{t}$ are masked by setting $z^{\prime}_{t}[y]=$-$\infty$. A valid token $y_{t}$ is sampled from $\mathrm{softmax}(z^{\prime}_{t})$.
  4. 4.
    The automaton transitions to $s_{t+1}=\delta(s_{t},y_{t})$. We record an audit tuple $\tau_{t}=\langle s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t\rangle$, where $s_{t+1}=\delta(s_{t},y_{t})$. Optionally a debug variant $\widetilde{\tau}_{t}$ serializes $M_{t}$ itself when needed.
The trace $\tau=\{\tau_{1},\ldots,\tau_{T}\}$ is streamed to the Evidence Registry. From $\tau$ we extract a DFA acceptance path $\rho$ and generate $\pi_{\text{struct}}=\langle\rho,\text{cert}_{L1}\rangle$. Each run is bound to a specific compiled automaton via automaton_id =H(Schema/GBNF source)=H($\text${Schema/GBNF source}), so that external auditors can re-validate L1 conformance using only the artifact, $\pi_{\text{struct}}$, and the automaton definition—without trusting the LLM itself. After generation, L2 validators (semantic validators for $\mathcal{C}_{\mathrm{sem}}$, logic validators for $\mathcal{C}_{\mathrm{log}}$) operate independently on the completed artifact, producing $\pi_{\text{sem}}$ and $\pi_{\text{logic}}$ that, together with $\tau$, form the composite evidence $\Pi$ (Section 3.5).
Expressiveness Baseline and Controlled Restriction.
Before analyzing quality trade-offs, we state an upper bound on expressiveness for unconstrained decoding.
Theorem 9 (Maximum Expressiveness of Unconstrained Strategy).
Let $h_{\mathrm{free}}$ denote unconstrained LLM decoding. Then $\mathcal{L}(h_{\mathrm{free}})=\mathcal{L}_{\text{LLM}}$, the full language model distribution over token sequences.
Proof sketch.
$h_{\mathrm{free}}$ samples directly from $P_{\theta}(y_{t}\mid y_{<t})$ with no masking. Any constrained strategy prunes tokens not in $M_{t}$, restricting the reachable sequences to a (possibly strict) subset. Thus $\mathcal{L}(h_{\mathrm{free}})\supseteq\mathcal{L}(h_{\text{constrained}})$ for any constrained strategy. ∎
Unconstrained decoding maximizes linguistic fluency and variety but does not guarantee structural, semantic, or logic correctness. Our L1/L2 split deliberately trades some expressiveness for verifiability and auditability (cf. the repair convergence analysis in Section 3.6).
Generation Quality vs. Structural Guarantees.
Enforcing DFA-based masking at L1 alters the model’s token distribution, effectively pruning the probability mass of structurally invalid tokens to zero. While ensuring correctness, this strict truncation can exacerbate distribution distortion: the model may be biased toward the shortest valid path (e.g., skipping optional fields or closing lists early) to maximize per-step probability, resulting in syntactically correct but semantically impoverished artifacts (parkGrammaralignedDecoding2025).
To mitigate this, ATLAS employs three runtime strategies tailored to safety-critical model-driven engineering, applicable across both DFA and PDA execution modes:
  1. (i)
    Minimum structural coverage. Instead of purely relying on the automaton’s accepting states ($F$) to allow termination, the executor enforces configurable completeness thresholds. For example, specific transitions leading to the End-of-Sequence (EOS) token are masked until minimum element counts are met or mandatory ”optional” fields (e.g., unique IDs) are generated. This forces the model to explore deeper structural paths.
  2. (ii)
    Two-stage generation with targeted refinement. Phase 1 produces a structurally valid draft under tight constraints. Phase 2 applies targeted refinement using AGR (Section 3.6), relaxing only semantic/logic aspects to enrich detail while preserving the structural invariants already certified by L1.
  3. (iii)
    Calibrated unfolding depth. We tune the unfolding bound $d$ from Theorem 7. Shallower unfolding reduces automaton size while still permitting diverse valid continuations, preventing the ”over-pruning” that often leads to mode collapse in deeply nested recursive grammars.
In practice, these heuristics balance the rigorous requirements of formal verification with the linguistic diversity of LLMs.
Verification-Guided DFA Masking and Prefix Safety.
L1 enforcement guarantees that every prefix remains correctable to an L1-valid artifact. We compile all structural constraints in $\mathcal{C}_{\mathrm{struct}}$ into a single prefix-closed DFA $\mathcal{A}=(Q,\Sigma,\delta,q_{0},F)$ (or its bounded unfolding $\mathcal{A}_{d}$), where every non-final state $q\notin F$ has a path $q\leadsto F$ to acceptance. During generation, after emitting prefix $y_{1..t}$ reaching $q_{t}$, the executor tests each candidate token $y\in\Sigma$ by computing $q^{\prime}=\delta(q_{t},y)$ and admits $y$ into $M_{t}$ iff $q^{\prime}\leadsto F$. If EOS occurs in non-final $q_{t}\notin F$, the executor computes the shortest accepting suffix $\mathcal{T}^{\star}(q_{t})$ via breadth-first search and appends it automatically, yielding structural repair bound $\lambda_{\text{struct}}\leq 1$.
Theorem 10 (Prefix Safety and Bounded Structural Repair).
Under the unified automaton execution framework, for any prefix $y_{1..t}$ produced by the verification-guided decoder over a prefix-closed DFA $\mathcal{A}$ (including bounded-unfolded $\mathcal{A}_{d}$), the current state $q_{t}$ satisfies: (i) $q_{t}\leadsto F$(reachability to acceptance is maintained); (ii) there exists a unique shortest completion suffix $\mathcal{T}^{\star}(q_{t})$; (iii) applying a single closure operation (appending $\mathcal{T}^{\star}(q_{t})$) yields an artifact that satisfies all L1 structural constraints.
Proof sketch.
(i) and (ii) follow from the reachability-based masking policy. For (iii), if generation halts at state $q_{t}\notin F$, the masking guarantee $q_{t}\leadsto F$ ensures a path exists. $\mathcal{T}^{\star}(q_{t})$ represents the minimal sequence of tokens required to traverse from $q_{t}$ to the nearest state in $F$. Appending this suffix constitutes a single deterministic repair action ($\lambda_{\text{struct}}\leq 1$), ensuring the final artifact is structurally valid by construction. ∎
Runtime Overhead.
The unified automaton execution imposes negligible latency. Since the allowed token set $M_{t}$ depends only on the current state’s outgoing transitions, the per-step masking overhead is minimal (<1<1ms on commodity hardware), scaling efficiently with the vocabulary size $|\Sigma|$ and the complexity of the constraint schema.
3.5. Constraint-Guided Verifiable Generation
Role in ATLAS.
Constraint-Guided Verifiable Generation (CVG) is the final stage of the ATLAS pipeline. Given an artifact $a$ generated under Layer-1 (L1) structural guidance (Section 3.4) and validated under Layer-2 (L2) semantic/logic certification (Section 3.3), CVG assembles machine-checkable evidence that proves conformance to domain constraints and binds that evidence to the artifact. The result is a verifiable artifact that can be audited by an external authority without rerunning the (possibly proprietary) generation process. Unlike purely post-hoc verification approaches (misuAIAssistedSynthesisVerified2024; parthasarathygauravTrustworthyAutomatedProgram2024), ATLAS records generation-time traces and post-generation validation results and fuses them into a single composite certificate.
Theoretical Foundation: Map of This Section.
For clarity, we organize the formal development as follows:
  1. (i)
    Verifiable artifacts and composite evidence (Definition 11): how we package an artifact $a$ with its evidence $\Pi$ and a verifier $\varphi$.
  2. (ii)
    Evidence composition (Definitions 1214): how structural, semantic, and logic traces are combined using $\otimes$ (sequential composition) and $\oplus$ (audit enrichment).
  3. (iii)
    Sound compilation (Theorem 7, Section 3.3): why L1/L2 validators preserve the semantics of source constraints $\mathcal{C}_{\mathrm{struct}}$, $\mathcal{C}_{\mathrm{sem}}$, $\mathcal{C}_{\mathrm{log}}$ under bounded depth $d$.
  4. (iv)
    End-to-end correctness: Section 3.4 establishes prefix safety and bounded structural repair (Theorem 10); later, Theorem 16 (Section 3.5) establishes incremental trace soundness for whole-artifact guarantees.
Together, these elements ensure that each generated artifact is paired with independently checkable, regulator-ready evidence $\Pi$, not just with an informal “the model said so” claim.
Definition 3.0 (Verifiable Artifact with Composite Evidence).
A verifiable artifact is a tuple $\mathfrak{A}=(a,\Pi,\varphi)$ where:
  • $a\in\mathcal{L}$ is the generated artifact, formally a structured instance model conforming to the M2-level UMM (realized as ARXML, JSON, or source code AST depending on the domain);
  • $\Pi$ is a composite evidence object aggregating layer-specific traces and runtime audits;
  • $\varphi:\Pi\to\{\top,\bot\}$ is a verifier that evaluates $\Pi$ and returns whether the evidence proves that $a$ satisfies the required constraints.
Evidence Composition Operators.
We next formalize how $\Pi$ is assembled from per-layer traces.
Let $\mathcal{T}$ denote the space of validation traces with temporal ordering $\preceq_{t}$. A validation trace $\pi=(T,V,\sigma)$ comprises timestamps $T$, validator outcomes $V$, and a satisfaction mapping $\sigma$ from constraints to validation status.
Sequential Composition ($\otimes$). For traces $\pi_{1}=(T_{1},V_{1},\sigma_{1})$ and $\pi_{2}=(T_{2},V_{2},\sigma_{2})$:
(17) $\pi_{1}\otimes\pi_{2}$ $=(T_{1}\cup T_{2},\;V_{1}\cup V_{2},\;\sigma_{1}\cup\sigma_{2})$
$\quad\text{subject to }\max(T_{1})\preceq_{t}\min(T_{2}),$
i.e., later-stage validation (e.g., L2 semantic checks) must occur after earlier-stage validation (L1 structural checks).
Audit Enrichment ($\oplus$). Given a validation evidence object $\Pi_{0}$ and a runtime audit trail $\tau=\{\tau_{1},\ldots,\tau_{T}\}$ (Definition 15):
$\Pi_{0}\oplus\tau\;=\;\Pi_{0}\cup\{\text{metadata}(\tau)\},$
where $\text{metadata}(\tau)$ captures process-level observability (e.g., automaton state transitions, masking statistics, timing), without modifying Boolean pass/fail states.
Verifier composition. The verifier $\varphi$ is required to satisfy:
$\varphi(\pi_{1}\otimes\pi_{2})=\varphi(\pi_{1})\wedge\varphi(\pi_{2}),\qquad\varphi(\Pi_{0}\oplus\tau)=\varphi(\Pi_{0}).$
Thus, $\oplus$ enriches evidence with audit metadata without changing correctness, while $\otimes$ enforces conjunctive validity across layers.
Definition 3.0 (Composite Evidence and Composition Operators).
Let $\pi_{\text{struct}}$, $\pi_{\text{sem}}$, $\pi_{\text{logic}}$ be the per-layer validation traces for structural, semantic, and logic constraints, and let $\tau$ be the runtime audit trail.
For notational alignment with the system diagrams, define $\pi_{1}:=\pi_{\text{struct}}$ and $\pi_{2}:=\pi_{\text{sem}}\otimes\pi_{\text{logic}}$.
We then write the composite evidence as
$\Pi=(\pi_{1}\otimes\pi_{2})\oplus\tau=\big(\pi_{\text{struct}}\otimes\pi_{\text{sem}}\otimes\pi_{\text{logic}}\big)\oplus\tau.$
Operationally:
  • $\otimes$ (sequential composition) appends validator outputs in causal order (L1 $\rightarrow$ L2), preserving timestamps and dependency structure.
  • $\oplus$ (audit enrichment) injects runtime audit metadata $\tau$ into the composed traces without altering acceptance.
Finally, $\varphi(\Pi)=\top$ if and only if $\varphi(\pi_{\text{struct}})=\varphi(\pi_{\text{sem}})=\varphi(\pi_{\text{logic}})=\top$. When $\varphi(\Pi)=\top$, we seal the result using a temporal hash $H(a\,\|\,\Pi\,\|\,t)$ to bind artifact $a$, evidence $\Pi$, and timestamp $t$ into an immutable record.
Lemma 13 (Audit Enrichment Conservativity).
Let $\Pi_{0}=\pi_{\text{struct}}\otimes\pi_{\text{sem}}\otimes\pi_{\text{logic}}$ and $\Pi=\Pi_{0}\oplus\tau$. If $\varphi(\Pi_{0})=b$ for $b\in\{\top,\bot\}$, then $\varphi(\Pi)=b$.
Proof sketch.
By Definition 12, $\oplus$ appends audit metadata $\tau$ but does not alter the Boolean outcomes of $\pi_{\text{struct}},\pi_{\text{sem}},\pi_{\text{logic}}$. Since $\varphi$ is defined as the conjunction of those outcomes, $\varphi(\Pi)=\varphi(\Pi_{0})$. ∎
Lemma 13 guarantees that attaching process-level provenance (forensics, performance metrics) cannot “flip” a pass into a fail or vice versa. Evidence is extensible but judgment is stable.
Compositional Verification Architecture.
Figure 8. ATLAS’s validation and evidence composition. The generated artifact and the decode-time audit trail are checked by structural, semantic, and logic validators derived from the ICM. Each stage emits a machine-checkable trace. These traces are bundled into a single evidence package that is used both to accept or reject the artifact and to drive targeted repair when checks fail.
Workflow for validation and evidence composition in ATLAS. Inputs include the generated artifact, the runtime audit trail from Layer 1 guided decoding, and validators compiled from the Integrated Constraint Model (ICM): Layer 1 structural validators such as DFA, JSON Schema, and GBNF, and Layer 2 semantic and logical validators such as SHACL and SMT, including constraints produced with LLM assistance. Validation executes in structural, then semantic, then logic order. Each stage produces a machine-checkable trace that is appended to an Evidence Registry. These traces are combined into a composite evidence package Pi together with the audit trail. A Trace Verifier evaluates this package; if it passes, the artifact plus evidence is sealed and exposed as a verifiable artifact, and if it fails, an Audit-Guided Repair process uses the evidence to generate targeted fixes.
Figure 8 shows that verification is modular. Each validator operates on the completed artifact $a$ but targets one constraint family: L1 structural correctness ($\mathcal{C}_{\mathrm{struct}}$), L2 semantic conformance ($\mathcal{C}_{\mathrm{sem}}$), and L2 logic/temporal consistency ($\mathcal{C}_{\mathrm{log}}$). This modularity yields three critical properties: (i) independent auditability: third parties can re-check a single layer using only $a$ and the corresponding $\pi_{\bullet}$; (ii) incremental repair: a semantic or logic violation at L2 does not erase the structural guarantees at L1, enabling targeted AGR; (iii) composable evidence: $\Pi$ is literally the conjunction (via $\otimes$) of per-layer certificates.
Layer-Specific Traces.
Definition 3.0 (Compositional Verification Trace).
Given constraint families $\mathcal{C}=\mathcal{C}_{\mathrm{struct}}\cup\mathcal{C}_{\mathrm{sem}}\cup\mathcal{C}_{\mathrm{log}}$, the composite verification trace for artifact $a$ is:
$\mathcal{T}_{a}\;=\;\bigotimes_{i\in\{\text{struct},\text{sem},\text{logic}\}}\mathcal{T}_{i}\;\oplus\;\tau,$
where:
  • $\mathcal{T}_{\text{struct}}=\langle Q,\delta,q_{0},F,\rho\rangle$: DFA execution trace for L1 structural constraints. $Q$ is the DFA state set, $\delta$ the transition function, $q_{0}$ the initial state, $F$ the accepting states, and $\rho:\{1,\ldots,T\}\to Q$ the run induced by generation. Acceptance requires $\rho(T)\in F$.
  • $\mathcal{T}_{\text{sem}}=\langle\mathcal{S},\mathcal{V},\sigma\rangle$: Semantic validation report for L2 semantic constraints. $\mathcal{S}=\{s_{1},\ldots,s_{n}\}$ are semantic constraint specifications (e.g., SHACL shapes), $\mathcal{V}\subseteq\mathcal{S}$ are violated shapes, and $\sigma:\mathcal{S}\to\{\text{conforms},\text{violates}\}$. Acceptance requires $\mathcal{V}=\emptyset$.
  • $\mathcal{T}_{\text{logic}}=\langle\Phi,\mathcal{M},\mu\rangle$: Logic validation certificate for L2 logic/temporal/numeric constraints. $\Phi$ are logic constraint formulas (e.g., SMT-LIB2), $\mathcal{M}$ is a satisfying model (if $\Phi$ is satisfiable) or an unsat-core $\mathcal{U}\subseteq\Phi$, and $\mu$ is the solver proof object. Acceptance requires $\Phi$ to be satisfiable.
  • $\tau$: runtime audit trail (Definition 15) capturing generation-time decisions and constraint enforcement, sealed by a temporal hash $H(a\,\|\,\Pi\,\|\,t)$ for tamper evidence.
Runtime Audit Trail: Capturing Generation Decisions.
The runtime audit trail $\tau$ is the process-level provenance of how the LLM was guided by L1 constraints (Section 3.4). It complements the post-hoc validators by documenting how the artifact was produced.
Definition 3.0 (Audit Tuple Structure).
At decoding step $t$, CVG records:
$\tau_{t}\;=\;\langle s_{t},\;|M_{t}|,\;y_{t},\;s_{t+1},\;\Delta t\rangle$
where:
  • $s_{t}$ is the automaton configuration (state $q\in Q$ for DFA or stack configuration for PDA) before emitting the token, aligning with the execution semantics in Section 3.4;
  • $|M_{t}|\in\mathbb{N}$ is the size of the allowed token set $M_{t}=\mathrm{Allow}(s_{t})=\{y\in\Sigma\mid\delta(s_{t},y)\neq\bot\}$, which quantifies how restrictive the mask was;
  • $y_{t}\in\Sigma$ is the token actually selected after masking;
  • $s_{t+1}=\delta(s_{t},y_{t})$ is the successor configuration;
  • $\Delta t\in\mathbb{R}^{+}$ is per-step latency for performance profiling.
The full audit trail is the sequence $\tau=\{\tau_{1},\ldots,\tau_{T}\}$.
Memory-Aware Audit Recording.
To keep runtime overhead tractable, the default mode records only
$(s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t)$ per step. Storing $|M_{t}|$ instead of the full $M_{t}$ keeps per-step cost $\mathcal{O}(1)$ rather than $\mathcal{O}(|\Sigma|)$. When full forensic mode is enabled, and when $|M_{t}|<1000$, the framework may also log the entire $M_{t}$ for that step. The transition $\delta(s_{t},y_{t})$ is stored implicitly via $(s_{t},s_{t+1})$, reducing storage from $\mathcal{O}(|Q|\cdot|\Sigma|)$ to $\mathcal{O}(T)$.
Automaton Identity Binding.
Each audit trail $\tau$ is bound to a specific compiled automaton via
$\textit{automaton\_id}=H(\text{Schema/GBNF source}),$
where $H$ is a cryptographic hash. This lets an auditor reconstruct the DFA/PDA from the same source constraints (ICM-derived $\mathcal{C}_{\mathrm{struct}}$) and check that $\tau$ is a valid accepting run without trusting the LLM or rerunning generation. In production, automaton_id can simply be a version-control commit hash for the constraint schema.
Evidence Registry: Append-Only Storage.
All evidence—including $\tau$, $\pi_{\text{struct}}$, $\pi_{\text{sem}}$, $\pi_{\text{logic}}$, and the composed $\Pi$—is stored in an append-only Evidence Registry. To balance auditability with throughput, the registry uses a two-tier strategy: (i) an in-memory bounded LRU cache to keep active trails for near-term refinement and AGR; and (ii) asynchronous persistence to append-only durable storage. Entries are content-addressed by cryptographic hash, enabling deduplication when multiple artifacts share constraint subsets. A $\Pi$-manifest records $\Pi=(\pi_{\text{struct}}\otimes\pi_{\text{sem}}\otimes\pi_{\text{logic}})\oplus\tau$ together with version metadata (v1, v2, final), supporting iterative repair cycles.
Incrementality and Layer Scope.
L1 validation is prefix-decidable (Section 3.4): at each decoding step, the DFA run $\rho$ can be extended and checked. By contrast, L2 semantic and logic validators require the complete artifact to build global symbol tables, evaluate cross-file references, or solve temporal/logical constraints. This asymmetry yields the following guarantee.
Theorem 16 (Incremental Trace Soundness).
Let $a^{(t)}=y_{1}y_{2}\cdots y_{t}$ be a partial artifact at decoding step $t$. If the incremental structural trace $\mathcal{T}^{(t)}_{\text{struct}}$ satisfies $\varphi(\mathcal{T}^{(t)}_{\text{struct}})$$\top$ (i.e., the DFA run $\rho^{(t)}$ remains reachable to an accepting state), then for any structurally valid completion $a=a^{(t)}\cdot w$, there exists a trace extension $\Delta\Pi$ such that $\varphi(\mathcal{T}^{(t)}_{\text{struct}}\oplus\Delta\Pi)=\top$.
Proof sketch.
By Theorem 10, if the current DFA state $q_{t}$ is still reachable to some accepting state $q_{f}\in F$, there exists a shortest accepting suffix $\mathcal{T}^{\star}(q_{t})$. Appending this suffix yields a structurally valid completion. For conjunctions of multiple structural constraints, the sequential composition operator $\otimes$ preserves conjunctive satisfaction. Thus any prefix that has not violated L1 constraints can be extended to a fully L1-valid artifact. ∎
Theorem 16 formalizes prefix safety for evidence: if we have not violated L1 yet, we can finish in compliance. L2 (semantic/logic) evidence, however, is appended only once the artifact is complete and global context is known.
Engineering Considerations for Deployment.
The CVG runtime and Evidence Registry must serve both high-throughput generation and high-assurance auditing. To that end:
  • Configurable audit granularity. Deployments can choose: full logging (record every $\tau_{t}$ and, when $|M_{t}|$ is small, full $M_{t}$), summary mode (record only aggregate statistics and final DFA acceptance), or key-event logging (log only structural boundaries, cross-file references, and validator feedback points). This tunes overhead versus forensic richness.
  • Backend abstraction. Different grammar-constrained decoders (e.g., LR/PDA backends, DFA masking engines) expose different internal states. The audit layer normalizes these into a canonical tuple $(s_{t},|M_{t}|,y_{t},s_{t+1},\Delta t)$ so that evidence is portable across backends.
  • Deferred context binding. In high-throughput settings with shared schemas, automaton_id binding can be deferred until evidence composition, avoiding expensive schema hashing on the hot path and attaching it only when $\Pi$ is finalized.
Table 1. Constraint hierarchy and evidence mapping. Each constraint family is enforced at its natural layer and emits a corresponding audit/validation artifact.
Type Layer-1 (Structural) Layer-2 (Semantic / Logic) Evidence Artifact
Structural JSON Schema / GBNF $\to$ DFA masking Cross-file reference integrity (deferred) $|M_{t}|$, rejection traces, $\pi_{\text{struct}}$
Semantic Enumerations / local vocab SHACL (hierarchy, reference consistency) SHACL reports in $\pi_{\text{sem}}$
Logic Local numeric guards SMT (temporal, numeric, safety logic) SMT models / unsat-cores in $\pi_{\text{logic}}$
Table 1 summarizes how constraint families map to enforcement layers and to evidence artifacts. Some constraints (e.g., enumerations, basic numeric guards) appear in both layers: L1 enforces the prefix-decidable fragment via DFA masking (preventing illegal tokens), while L2 re-checks global consistency via SHACL/SMT. This cross-layer assignment matches the compilation strategy in Section 3.3 and preserves both preventive guarantees during generation and certifying guarantees after generation.
3.6. Audit-Guided Intelligent Repair
Role in ATLAS.
By the time an artifact reaches Audit-Guided Repair (AGR), Layer-1 structural constraints have already been enforced. Prefix-closed DFA/PDA masking plus one-step completion (Theorem 10) guarantee that the draft artifact satisfies $\mathcal{C}_{\mathrm{struct}}$. Residual violations are therefore almost entirely Layer-2 failures: semantic graph obligations ($\mathcal{C}_{\mathrm{sem}}$) and logical/temporal/numeric obligations ($\mathcal{C}_{\mathrm{log}}$). While recent iterative refinement strategies leverage feedback loops (e.g., compiler logs or test outputs) (shinn2023reflexion; chen2024selfdebug), they often lack the formal provenance required to link failures back to normative standards. AGR advances this by employing an evidence-driven, localized repair strategy backed by formal or verifiable programmatic validators.
Formally, AGR consumes the composite evidence
[…]
and turns validator diagnostics into two coordinated repair routes:
  1. (i)
    Automated semantic/logic repair. The framework invokes an LLM in a tightly constrained mode to synthesize minimal edits that resolve the specific SHACL/SMT violation, using ICM-grounded context.
  2. (ii)
    Model-driven human repair. The artifact is also materialized into a meta-model-aware configuration editor (“Human Review Layer / MDE Editor” in Figure 1) built from the UMM (Path S1, Section 3.2). A domain engineer applies guided fixes in a structured UI that directly edits typed model classes, not raw text.
Both routes are driven by the same evidence $\Pi$, and both feed successful fixes back into ATLAS as reusable constraints (“constraint promotion”), so future generations avoid repeating the violation. The two routes differ only in who performs the edit (LLM vs. human) and how the edit is applied (direct text patch vs. meta-model–aware configuration update). This “graduated automation” design is explicitly supported in the ATLAS architecture: the Validation & Repair Layer and the Human Review Layer exchange artifacts, evidence, and compiled validators through the Evidence Registry and Configuration Editor.:contentReference[oaicite:1]index=1
Evidence-Driven Violation Localization.
Given a failed artifact $(a,\Pi,\varphi)$ with $\varphi(\Pi)=\bot$, AGR decomposes $\Pi$ to localize violations with Layer-2 precision:
  • Semantic violations ($\pi_{\text{sem}}$). The semantic validator (instantiated formally or programmatically) returns, for each violation: (i) a violation path (an address into $a$ such as an entity ID or XPath), (ii) the expected condition (e.g., “must reference an existing Operation”), and (iii) a constraint identifier. During ICM compilation (Section 3.3), each constraint is annotated with ICM provenance: which ICM entry generated this rule, and which UMM entity it is anchored to. AGR uses this to (a) locate the exact locus of failure in $a$, and (b) retrieve the authoritative rule and valid target types from the UMM.
  • Logic violations ($\pi_{\text{logic}}$). The logic validator identifies a minimal conflict set (e.g., an SMT unsat-core or a failed assertion trace) $\mathcal{U}\subseteq\Phi$ that pinpoints the inconsistent subset of numeric/temporal predicates. Each predicate in $\mathcal{U}$ is back-linked to concrete fields in $a$ and to its source obligation in the ICM (e.g., timing bounds, resource-usage limits). AGR can therefore highlight which values conflict and under which domain rule.
  • Structural trace ($\pi_{\text{struct}}$). The DFA run $\rho$ and runtime audit $\tau$ remain attached for auditability, but structural incompleteness has already been resolved before AGR via shortest completion $\mathcal{T}^{\star}$ (Theorem 10). AGR typically does not need further structural edits.
In other words, AGR does not have to “search for the bug”: the L2 validators produce machine-readable failure tickets that name the failing node, the violated obligation, and its ICM/UMM provenance. This is crucial for both automated and human-assisted repair.
Two Repair Routes: Automated Patch vs. MDE-Guided Edit.
The localized violations are then repaired through one of two routes.
Route A: Automated validation-guided regeneration. Unlike blind retry strategies, AGR utilizes the violation reports (e.g., SHACL violation paths or SMT unsat cores) to construct a constraint-aware repair prompt. This prompt explicitly injects the violated rules and the relevant UMM schema context, guiding the LLM to regenerate the specific artifact file (or affected block) with corrected logic. This represents a diagnostic-driven regeneration at the file granularity, which effectively narrows the search space without requiring token-level surgical editing.
Route B: Human-in-the-loop MDE editing. ATLAS acknowledges that for complex, global semantic violations (the “last 20%”), fully automated repair may not converge efficiently. In these cases, ATLAS projects the artifact into a context-aware configuration editor generated via Path S1. Crucially, AGR attaches the machine-readable validator diagnostics directly to the editor’s UI. This transforms the human role from manual authoring to evidence-based decision making, resolving ambiguities that are computationally undecidable for the LLM while leveraging the UMM structure to ensure valid edits.
This route is not an afterthought: it is the fallback path in ATLAS’s “graduated automation” model. Routine violations are auto-repaired, but high-stakes or ambiguous cases go through a configuration workflow that is already accepted by certification processes in safety-critical engineering: configure $\rightarrow$ validate $\rightarrow$ repair in an MDE tool.
Constraint Promotion (Feedback to Generation).
Whether a violation is fixed automatically or via the Configuration Editor, AGR performs constraint promotion: it records the violated SHACL shape or SMT clause (together with its ICM provenance and UMM anchor) and feeds it forward into future generation contexts. Concretely, the promoted constraint can: (i) be injected into retrieval-augmented prompts provided to the LLM during Instance Model Generation, (ii) tighten decoder-side guards for specific entity types, or (iii) be persisted in the ICM as a reusable domain rule if it was previously missing. As a result, the system does not merely “patch and forget”; it accumulates domain obligations and prevents recurring violations on subsequent artifacts.
Dependency-Aware Scheduling and Convergence.
AGR orders fixes using the constraint lattice $\Gamma$ (Definition 4), which encodes precedence between obligations (e.g., well-formed reference bindings take priority over numeric tuning). Algorithm 4 summarises the workflow, now with two explicit repair routes and promotion:
Algorithm 4 Layer-2-Guided Repair with Human/LLM Routes and Constraint Promotion
1:Artifact $a$, Composite evidence $\Pi$, Constraint set $\mathcal{C}$
2:Repaired artifact $a^{*}$ or MANUAL_REVIEW; updated generation context
[…]
8:for $v\in\mathcal{V}_{\text{sorted}}$ do
9:  $\text{prov}\leftarrow\text{ICMProvenance}(v)$ $\triangleright$ Back-pointer to ICM entry and UMM anchor
10:  $\text{ctx}\leftarrow\text{RetrieveFromKG}(v.\text{targetEntity},\text{prov})$
11:  if isAutoRepairable(v) then
12:    $\text{prompt}\leftarrow\text{BuildRepairPrompt}(v,\text{ctx},\Pi)$
13:    $a\leftarrow\text{LLM}(\text{prompt},T{=}0.2)$ $\triangleright$ Route A: constrained LLM patch
14:  else
15:    $a\leftarrow\text{ApplyEditorFix}(a,v,\text{prov},\text{ctx})$ $\triangleright$ Route B: human edit in UMM-derived Configuration Editor
16:  end if
17:  $\text{PromoteConstraintToGeneration}(\text{prov})$ $\triangleright$ Persist rule / tighten future prompts and guards
18:  $\mathcal{R}\leftarrow\text{LocalRevalidate}(a,v,\mathcal{C})$
19:  if $\mathcal{R}\neq\emptyset$ then
20:    return MANUAL_REVIEW
21:  end if
22:end for
23:return $a$ as $a^{*}$
The repair convergence guarantee still holds under two observations: (i) Layer-1 structural conformance is already guaranteed and is not re-broken by AGR; and (ii) each remaining violation is either auto-fixable with probability >0.5>0.5 using ICM-grounded prompts, or is surfaced to a domain engineer through a configuration workflow that encodes the UMM and ICM constraints and is already standard practice in model-driven engineering toolchains (configure $\rightarrow$ validate $\rightarrow$ repair).
4. Evaluation
4.1. Overview and Research Questions
The evaluation validates the ATLAS framework through three complementary research questions that progress from single-artifact generation with comprehensive constraint enforcement to multi-artifact system generation under scalability stress, culminating in cross-domain transferability assessment where explicit meta-models are unavailable. This progression addresses three fundamental challenges in LLM-driven artifact generation for regulated domains: establishing formal correctness guarantees for individual artifacts, maintaining consistency across interdependent artifacts at system scale, and demonstrating architectural generalizability beyond domains with rich formal specifications.
We structure the evaluation around the following research questions:
RQ1 (Single-File Generation): Can layered constraint enforcement with evidence-carrying generation reliably achieve structural correctness on single-file AUTOSAR components while keeping decoding latency and token cost within practical bounds?
RQ2 (Multi-File System Generation): Can the ATLAS framework scale from single-file to multi-file AUTOSAR systems, preserving Layer-1 structural correctness while assessing how well Layer-2 semantic and Layer-3 logical constraints can be satisfied under realistic cross-file dependencies?
RQ3 (Cross-Domain Transferability): Can the UMM-ICM-CVG architecture, instantiated via the UMM S2 path, transfer to a legal domain without an explicit meta-model and improve legal correctness, citation quality, and rule compliance beyond LLM-only and RAG baselines while maintaining layered constraint enforcement and evidence-carrying generation?
The primary evaluation domain is AUTOSAR (AUTomotive Open System ARchitecture), a mature model-driven engineering standard for automotive software that provides well-defined meta-models, comprehensive constraint specifications, and industrial validation tooling. AUTOSAR serves as an ideal testbed for RQ1 and RQ2 because it exhibits the complexity characteristics of safety-critical artifact generation—hierarchical component structures, cross-file reference integrity requirements, and rich semantic constraints—while offering authoritative schemas for validation. For RQ3, we evaluate the framework on jurisdiction determination under the Brussels I bis Regulation in Private International Law, a domain lacking explicit meta-models but requiring formal constraint reasoning over precedence hierarchies and prerequisite conditions. This cross-domain probe assesses whether the framework’s modular architecture—separating meta-model construction pathways (S1 deductive transformation versus S2 inductive extraction), constraint compilation (ICM), and constraint enforcement (CVG)—supports deployment beyond engineering domains with established formal artifacts.
Evaluation methodology employs quantitative metrics for structural correctness (XSD validation), semantic consistency (SHACL validation for graph constraints, SMT validation for logic constraints), cross-file reference integrity, repair convergence efficiency, and computational cost. We complement automated metrics with expert review for system-level AUTOSAR configurations, assessing requirements traceability, architectural quality, engineering usability, and toolchain integration outcomes. All experiments use reproducible configurations with fixed random seeds, documented model versions, and archived generation artifacts to support external validation.
4.2. AUTOSAR Domain Preparation
To ensure reproducibility and external validation of the AUTOSAR experiments, we establish a three-layer domain representation that serves as the foundation for all subsequent generation, validation, and repair procedures. This representation comprises the Unified Meta-Model (UMM), the Instance-Constraint Mapping (ICM), and an executable Knowledge Graph (KG), each constructed through systematic transformation and extraction processes.
4.2.1. Unified Meta-Model Construction
The UMM provides a canonical, machine-readable view of AUTOSAR structure. We derive the UMM exclusively through deterministic model transformation from official AUTOSAR metamodel artifacts in XSD (XML Schema Definition) and XMI (XML Metadata Interchange) formats, without relying on unconstrained large language model inference or speculative schema induction. This corresponds to the S1 pathway described in Section 3, where authoritative platform metamodels exist and are lifted into normalized representations.
The transformation process flattens and reconciles the AUTOSAR metamodel into an explicit graph of core types—including software components, ports, interfaces, runnables, signals, and timing attributes—together with their structural relations such as containment, reference targets, allowed cardinalities, and typed links. The UMM thus acts as a canonical vocabulary that names each concept, fixes its attributes, and records which other concepts it may reference. This vocabulary consistency is essential for downstream constraint compilation and validation.
4.2.2. Instance-Constraint Mapping Construction
Building upon the UMM, we construct an ICM that enriches the meta-model with structured constraints extracted from AUTOSAR PDF specifications and related normative documents. The ICM extraction process follows the dual-channel approach described in Section 3.3, combining deductive extraction from formal schemas with inductive extraction from natural-language requirements.
The extracted constraints encompass four primary categories: mandatory element presence requirements, allowed value ranges for attributes, reference integrity obligations that specify compatible target types, and timing or behavioral expectations expressed in machine-checkable form. Each constraint is linked back to UMM types and relations rather than concrete instances, ensuring reusability. For example, the ICM does not specify that a particular component instance must have field $f$; instead, it declares that any element of metamodel type $T$ must expose role $r$ with property $p$. This type-level linkage enables the same constraint to be applied wherever an element of type $T$ is instantiated or validated.
The resulting ICM transforms natural-language obligations from specification documents into structured, typed constraints aligned with the UMM vocabulary, bridging the gap between informal domain knowledge and formal verification requirements. This ensures that constraints are not merely text, but executable logic anchored to the domain model.
Constraint Scope and Mapping Statistics.
Applying this dual-channel process to the AUTOSAR Software Component Template specification yielded a total of 1,161 normative constraints. The entity linking mechanism successfully anchored 1,045 of these to verified UMM entities (90% precision), with the remainder flagged for review. For evaluation, we map these constraints into executable L2 validators based on their logical characteristics: 794 SHACL node shapes target graph-structural invariants, and 1,005 SMT assertions target logic and value constraints. Crucially, a significant subset of these constraints governs reference integrity and interaction patterns that, while locally defined, effectively manifest as cross-file dependencies in distributed system architectures. Consequently, satisfying these constraints requires not just local syntactic correctness but global semantic consistency. A subset of constraints unsuitable for formal validation (e.g., purely informational hints) was excluded from the automated pipeline.
4.2.3. Knowledge Graph Materialization
The final preparation step materializes a navigable, queryable KG that fuses UMM types and relations with ICM typed constraints. Nodes in the KG correspond to metamodel entities such as software component types, port types, and interface types, while edges capture both structural relations—containment hierarchies, reference targets, cardinalities—and attached constraints, including requirements like ”must provide signal ss” or ”must reference interface of category cc”.
This KG serves dual roles in the experimental environment. First, as a retrieval substrate for prompt construction, the KG enables targeted extraction of relevant metamodel subgraphs rather than arbitrary prose. When generating a specific component, the system retrieves KG subgraphs exposing expected ports, interfaces, and required attributes in a form already aligned with the AUTOSAR metamodel, serving as structural blueprints for generation. Second, as a dynamic schema assembly engine, the KG supports on-demand composition of structural schemas and constraint snippets—such as required fields, allowed references, and value obligations—for downstream generation and validation. Because ICM constraints are linked at the metamodel level, the system can construct appropriate schemas for any candidate component configuration without manual per-instance rule rewriting.
In summary, the AUTOSAR domain preparation yields a three-layer knowledge substrate: the UMM obtained through deterministic transformation from official artifacts, the ICM formed through dual-channel constraint extraction and entity linking, and the KG that operationalizes both structure and constraints for retrieval and dynamic schema construction. This layered representation, containing 1,161 structured constraints over more than 1,000 metamodel entities and 3,049 edges, underpins all AUTOSAR experiments reported in this section.
4.3. RQ1: Single-File AUTOSAR Component Generation
RQ1 evaluates the complete ATLAS pipeline on single-component AUTOSAR generation , examining whether UMM-ICM-guided retrieval, layered constraint enforcement, an d audit instrumentation deliver structural correctness with tractable decoding l atency and token overhead.
4.3.1. Experimental Setup
Research Question.
RQ1 follows §4.1: singl e-file AUTOSAR generation, asking whether Layer-1 constraint enforcement alone c an guarantee structural validity while maintaining acceptable decoding latency a nd token cost.
Baseline Comparisons.
We evaluate four pipeline configurations that represent progressively structured approaches from the recent LLM literature, each implemented using the same base model to isolate the impact of constraint enforcement strategies:
  1. (1)
    vLLM (pure prompting baseline): Direct generation without retrieval augmentation or Layer-1 enforcement, mirroring common practice in LLM-based information extraction systems where pretrained models are directly prompted to emit structured records (Dagdelen2024; Lu2022).
  2. (2)
    vLLM+RAG: Augments prompts with facts retrieved from the AUTOSAR UMM and KG, analogous to instruction-following text-to-database agents that ground extraction in existing schemas (Jiao2024).
  3. (3)
    vLLM+RAG+JSON Schema: Enforces Layer-1 constraints via intermediate JSON instances deterministically projected to ARXML, approximating schema-aligned decoding (Lu2025b).
  4. (4)
    vLLM+RAG+GBNF: Applies grammar-constrained decoding using GBNF (GBNF Notation Format) derived from AUTOSAR specifications, masking invalid continuations at each generation step to allow only grammar-admissible prefixes, following the constrained decoding paradigm (scholakPICARDParsingIncrementally2021; zotero-item-1597; Park2025).
All pipelines use the same backend model (DeepSeek-R1-Distill-Qwen-32B served through vLLM) with fixed random seeds (42, 1001, 20250701), temperature 0.7, and top-p sampling at 0.9 to ensure reproducibility. We evaluate 60 representative AUTOSAR components under three prompt regimes that vary the amount of natural-language specification context provided: Min (minimal requirements), Std (standard documentation), and Full (comprehensive specifications including timing and behavioral constraints).
Evaluation Metrics.
For this structural feasibility study, we focus on three dimensions:
  • Structural correctness: XSD validation pass rate indicating syntactic compliance with AUTOSAR schemas.
  • Evidence coverage: Audit trail completeness measuring the percentage of generation steps with recorded structural proofs $\pi_{\text{struct}}$.
  • Computational cost: End-to-end latency in seconds and token counts for input and output.
Note: Semantic consistency (SHACL) and logic verification (SMT) metrics are reserved for the system-level evaluation in RQ2, as single-file generation lacks the necessary cross-reference context.
4.3.2. Results
Computational Cost.
Table 2 presents end-to-end latency and token consumption across pipelines and prompt regimes. The pure vLLM baseline exhibits moderate latency and minimal input tokens due to lack of retrieval augmentation, but generates longer outputs containing structural errors. Adding RAG increases input context substantially (from 154–227 tokens to 992–6,649 tokens depending on regime) while improving output quality. Constrained decoding pipelines (GBNF and JSON Schema) reduce output tokens by enforcing compact, schema-compliant generation, with JSON Schema achieving the lowest latency due to deterministic intermediate representation.
Table 2. Computational cost for RQ1: end-to-end latency and token counts by pipeline and prompt regime. All measurements use DeepSeek-R1-Distill-Qwen-32B on the same hardware.
Pipeline Configuration Latency (seconds) Input Tokens Output Tokens
vLLM (Min) 21.87 154 812
vLLM (Std) 29.41 188 1,279
vLLM (Full) 46.31 227 2,013
vLLM+RAG (Min) 17.51 992 743
vLLM+RAG (Std) 33.19 2,534 1,397
vLLM+RAG (Full) 85.60 4,493 3,610
vLLM+RAG+GBNF (Min) 8.01 828 155
vLLM+RAG+GBNF (Std) 59.94 2,896 1,446
vLLM+RAG+GBNF (Full) 123.74 6,649 3,118
vLLM+RAG+JSON Schema (Min) 3.83 848 119
vLLM+RAG+JSON Schema (Std) 13.38 2,605 523
vLLM+RAG+JSON Schema (Full) 31.56 2,915 1,291
Evidence Coverage.
For the vLLM+RAG+JSON Schema pipeline, the structured-output path records complete audit coverage with per-token events and Layer-1 candidate sets, producing structural proof $\pi_{\text{struct}}$ for all 60 test cases across all three prompt regimes. Table 3 summarizes audit statistics. The allowed token set size averages 156 tokens per generation step with minimum of 1 (fully constrained choices) and maximum of 650 (unconstrained text fields), demonstrating effective constraint propagation through the DFA. Audit trace density approximates three events per accepted token, providing fine-grained provenance for verification. The vLLM+RAG+GBNF pipeline was not instrumented with the same audit recorder in our current implementation, representing an engineering limitation rather than fundamental incompatibility with GBNF-based constraint enforcement.
Table 3. Layer-1 auditing statistics for vLLM+RAG+JSON Schema pipeline, showing audit coverage and constraint effectiveness across prompt regimes.
Prompt Regime Allowed Token Set (median / min) Generation Steps Audit Coverage Structural Proof $\pi_{\text{struct}}$ Availability
Min 156 / 1 (max $\leq$ 650) 119 100% 100%
Std 156 / 1 (max $\leq$ 650) 524 100% 100%
Full 156 / 1 (max $\leq$ 650) 1,292 100% 100%
Structural Correctness (Syntax Feasibility).
Since single-component generation in isolation lacks the system-level context required to satisfy global semantic constraints (e.g., references to external interfaces), RQ1 focuses strictly on structural validity—the prerequisite for any downstream engineering utility. Table 4 presents the XSD validation pass rates. The pure vLLM baseline failed to produce valid XML (0% pass rate), and even the frontier LLM-API baseline (GPT-5) achieved only a 50% pass rate, struggling with the strict nesting and typing rules of ARXML in a zero-shot setting. In contrast, all pipelines equipped with Layer-1 enforcement (JSON Schema or GBNF) achieved 100% structural compliance. This result answers RQ1 affirmatively regarding structural correctness: the proposed L1 masking is sufficient to guarantee that the LLM outputs syntactically valid artifacts, serving as a reliable foundation for the system-level semantic assembly in RQ2.
Table 4. Structural correctness for RQ1: XSD pass rate per pipeline.
Method XSD Pass Rate (%)
vLLM 0
vLLM+RAG 100
vLLM+RAG+GBNF 100
vLLM+RAG+JSON Schema 100
LLM-API 50
4.3.3. Analysis and Discussion
Validation of Layered Constraint Architecture.
The stark contrast between the 0% XSD pass rate of the unconstrained vLLM baseline and the 100% pass rate of all constrained pipelines (JSON Schema, GBNF) empirically validates the necessity of Layer-1 enforcement. Notably, even the frontier LLM-API baseline, representing state-of-the-art unconstrained capability, only achieved a 50% pass rate, confirming that probabilistic sampling alone cannot guarantee the strict nesting and typing rules required by ARXML. This finding establishes that Layer-1 structural constraints are not optional optimizations but fundamental prerequisites for reliable artifact generation. By enforcing prefix safety during decoding, ATLAS transforms the LLM from a text generator into a valid model emitter, ensuring that every output is structurally ingestible by downstream engineering tools.
Evidence-Carrying Generation and Audit Guarantees.
The 100% audit coverage and structural proof availability for constrained pipelines demonstrate that evidence-carrying generation is practically achievable without prohibitive overhead. The audit trail density provides sufficient granularity to trace every structural decision back to the enforcing automaton state. This recorded trace enables deterministic reconstruction of the generation process, satisfying the reproducibility and auditability requirements of regulated domains—a capability entirely absent in standard API-based generation.
Computational Trade-offs.
The latency and token cost results (Table 2) reveal the inherent trade-offs in constrained generation. Constraint-aware retrieval substantially increases input tokens (up to 6,649 tokens in the Full regime) to provide necessary context, but effectively reduces output tokens through compact, compliant generation. The JSON Schema pipeline achieves the lowest latency by exploiting deterministic intermediate representations, while GBNF exhibits higher latency due to the overhead of per-token automaton state transitions. These metrics suggest that JSON Schema-based enforcement offers the optimal balance between performance and structural guarantees for this domain.
Synthesis: Structural Reliability as a Foundation.
RQ1 validates that ATLAS’s Layer-1 enforcement reliably solves the “syntax problem” of LLM-based generation. While single-file generation inherently precludes the verification of global cross-file semantics (the subject of RQ2), the achievement of 100% structural validity is critical. It ensures that the artifacts passed to the subsequent semantic validation stages are well-formed and parseable. Without this L1 guarantee, any attempt at higher-level semantic reasoning or repair would be thwarted by basic parsing failures, as evidenced by the unconstrained baselines. Thus, RQ1 confirms that the UMM-ICM-CVG architecture provides the necessary structural substrate upon which the complex semantic assembly of RQ2 can be built.
4.4. RQ2: Multi-File System Generation at Scale
The second research question extends the evaluation from the single-component setting of RQ1 to system-level, multi-file AUTOSAR configurations. Here we stress-test ATLAS under realistic cross-file dependencies, asking whether the UMM-ICM-CVG architecture can maintain the strong Layer-1 guarantees observed in RQ1 while surfacing the semantic and logical failure modes that only emerge at system scale.
4.4.1. Experimental Setup
Figure 9. Two-phase AUTOSAR generation, validation, and repair pipeline. Phase 1 synthesizes a high-level blueprint from NLP requirements and AUTOSAR standards, populating the UMM/ICM database. Phase 2 performs constrained LLM decoding to generate component-specific ARXML with JSON Schema enforcement. Post-generation validation applies XSD, SHACL, and SMT checks, with Audit-Guided Repair generating targeted fixes for any violations to produce a verified Decision Package.
System architecture for the AUTOSAR experiment shown as a two-phase pipeline. Phase 1 (AUTOSAR Blueprint) consumes NLP requirements, AUTOSAR standards/PNFs, and the AUTOSAR meta-model (XSD/XMI) to synthesize a high-level component and type layout, populating the UMM/ICM database and the AUTOSAR knowledge graph. Phase 2 (Component-wise Generation) assembles a dynamic JSON Schema for each component, performs constrained LLM decoding to emit per-component JSON, and converts that JSON to ARXML. All intermediate artifacts are logged in the Evidence Registry. After generation, only the produced ARXML is validated: it is checked by XSD validation, SHACL semantic validation, and SMT-based logical checks. A Failure Mode Logger and Audit-Guided Repair module analyze any violations to generate targeted fixes, yielding a verified Decision Package for the AUTOSAR Configuration Tool.
Research Question and Motivation.
RQ2 tests whether ATLAS can scale to multi-file AUTOSAR systems while preserving Layer-1 structural correctness, and to what extent Layer-2 semantic and Layer-3 logical constraints can be satisfied in the presence of thousands of cross-file references and global invariants that do not arise in RQ1.
Dual-Phase Pipeline Architecture.
To address these challenges, we adopt a blueprint-guided dual-phase generation strategy that decomposes system-level generation into two stages, each with distinct objectives and constraint enforcement mechanisms:
  1. (1)
    Phase 1 (Blueprint Synthesis): The system first generates a JSON-structured blueprint specifying the component inventory, dependency graph, file layout, and cross-component interface contracts. This blueprint serves as a global coordination artifact that establishes naming conventions, reference targets, and architectural topology before detailed component generation begins. The blueprint generation employs JSON Schema constraints to ensure structural validity and completeness of the system plan.
  2. (2)
    Phase 2 (Component Assembly): Guided by the blueprint, the system generates individual ARXML files for each component, with interfaces emitted as separate files referencing a shared basic-types definition. Each generation step uses the blueprint as context to resolve cross-component dependencies and maintain reference consistency. Component-level generation employs the same JSON Schema-constrained decoding used in Phase 1, ensuring structural compliance at the artifact level.
This dual-phase decomposition addresses scalability through separation of concerns: Phase 1 handles global architectural decisions that require system-wide reasoning, while Phase 2 focuses on component-level detail generation constrained by the established blueprint. The blueprint acts as a contract between phases, reducing the context window requirements for Phase 2 generation while providing explicit targets for cross-file references.
Model Selection and Constraint Enforcement.
Unlike RQ1, which evaluated local vLLM deployments with comprehensive audit trail recording, RQ2 employs a frontier API-based model (GPT-5) for both phases. This choice reflects three considerations:
  • Architectural reasoning capabilities: System-level blueprint synthesis requires stronger planning and reasoning abilities to maintain global invariants across complex dependency graphs. Frontier models demonstrate superior performance on multi-step reasoning tasks that involve decomposing high-level requirements into structured component inventories and dependency relationships.
  • Practical deployment scenarios: Many organizations deploy LLM-based tooling through API access rather than self-hosted inference, particularly for occasional system generation tasks where the overhead of local infrastructure deployment is unjustified. RQ2 evaluates a representative deployment scenario complementary to the local vLLM evaluation in RQ1.
  • Constraint enforcement preservation: Although API-based generation precludes the fine-grained audit trail recording demonstrated in RQ1, both phases employ JSON Schema-constrained decoding through the API provider’s structured output capabilities. Phase 1 blueprint generation uses a JSON Schema specifying component inventory structure, dependency graph format, and file layout constraints. Phase 2 component generation uses JSON Schemas derived from AUTOSAR metamodel fragments, ensuring that individual ARXML files satisfy structural requirements before deterministic XML serialization. Thus, Layer-1 constraint enforcement remains active, though without the token-level audit provenance available in the vLLM pipeline.
To balance creative architectural exploration with structural conformance, we adopt stage-specific temperature settings: Phase 1 blueprint synthesis uses temperature 0.7 to encourage diverse architectural solutions, while Phase 2 detailed generation uses temperature 0.3 for more deterministic output, and interface file emission uses temperature 0.2 for maximum consistency. Each system is generated once without seed variation, reflecting single-shot engineering workflows common in industrial practice.
Dataset and Evaluation Metrics.
We curate 20 AUTOSAR systems across three complexity tiers designed to stress different aspects of cross-file generation. The dataset spans three distinct architectural configurations that progressively challenge the framework’s ability to maintain cross-file consistency and semantic correctness.
The Simple tier comprises 7 cases with 27 components total, featuring chain topologies suitable for sensor-processor-actuator pipelines. Representative scenarios include data collection systems, control loops, and monitoring applications. Each system contains an average of 3 to 4 components, generating 52 ARXML files across the tier. The Middle tier consists of 7 cases with 45 components total, employing multi-branch tree topologies that represent sensor fusion architectures and multi-domain controllers integrating powertrain, body, and ADAS subsystems. Systems in this tier average 6 to 7 components each, producing 108 ARXML files. The Complex tier includes 6 cases with 54 components total, utilizing mesh topologies with bidirectional dependencies that model sophisticated scenarios such as autonomous driving stacks, by-wire chassis control, and vehicle-cloud coordination. These systems average 8 to 10 components each and generate 124 ARXML files.
We evaluate generated systems across six primary dimensions. File completeness measures whether all blueprint-declared files are produced without truncation, assessed post-generation for all 284 files across the 20 systems as a binary complete-or-incomplete determination. Structural correctness quantifies syntactic compliance with AUTOSAR XSD schemas through xmllint validation, reporting the XSD validation pass rate for all generated ARXML files. Cross-file reference integrity calculates the resolution rate as the proportion of resolvable references over total <REF> elements, measured by parsing all reference targets and verifying their existence in the generated artifact set. Toolchain usability evaluates whether systems successfully import into commercial AUTOSAR configuration tools, specifically Etas (ISOLAR-A), yielding a binary pass-or-fail outcome per system. Generation cost captures both token counts and wall-clock time, measured separately for Phase 1 blueprint synthesis and Phase 2 component assembly across all 20 systems.
Expert review follows a structured protocol evaluating four dimensions: requirements alignment assesses whether generated systems satisfy functional specifications and maintain traceability to original requirements; architectural quality evaluates dependency consistency and type-level correctness; engineering quality measures naming conventions, structural modularity, and estimated repair effort; toolchain integration validates whether systems import successfully into commercial AUTOSAR configuration environments.
4.4.2. Results
Structural Success vs. Semantic Failure.
Table 5 (updated with semantic metrics) illustrates a distinct degradation in model performance as constraints shift from syntactic to semantic and from local to global. Layer-1 (Structure): All 20 systems across all complexity tiers achieved 100% XSD pass rates, confirming that the blueprint-guided constrained decoding effectively solves the syntax generation problem. Layer-2 (Local Semantics via SHACL): When evaluating semantic constraints at the single-file level (checking roughly 794 graph shapes), the pass rates drop precipitously: Simple: 2/57 ($\sim$3.5%), Middle: 16/113 ($\sim$14.2%), and Complex: 6/63 ($\sim$9.5%). These failures largely stem from missing mandatory attributes or incorrect local configurations that are structurally valid but semantically vacuous. Layer-2 (Global Logic via SMT): When imposing SMT constraints that verify cross-file consistency (e.g., end-to-end signal flow, exclusive jurisdiction of components) and stacking them with SHACL, the system-level pass rate was 0% across all tiers. While the model often generated references that resolved to existing IDs (high reference resolution rate, see Table 5), the SMT validation revealed that these connections were frequently semantically invalid (e.g., connecting incompatible ports or creating circular dependencies). This 0% baseline confirms that large-scale system consistency is beyond the zero-shot capability of current LLMs, necessitating the iterative repair and human-in-the-loop workflows ATLAS provides.
Table 5. Structural correctness and cross-file reference integrity by complexity tier. All systems achieve perfect file completeness and XSD compliance, while reference resolution degrades with increasing system complexity.
Complexity Systems Total Files Completeness XSD Pass Rate Total References Resolved References Resolution Rate
Simple 7 52 52/52 (100%) 52/52 (100%) 605 605 100.0%
Middle 7 108 108/108 (100%) 108/108 (100%) 1,754 1,744 99.4%
Complex 6 124 124/124 (100%) 124/124 (100%) 2,086 1,936 92.8%
Total 20 284 284/284 (100%) 284/284 (100%) 4,445 4,285 96.4%
Table 6. Expert review findings aggregated by complexity tier. Scores use 5-point scales for quantitative dimensions; C3 uses categorical assessment (minor/moderate/extensive); D1 reports binary import status. Findings illustrate representative failure patterns observed within each tier rather than universal characteristics.
Dimension Scores Representative Findings
Simple Tier (7 systems, 27 components)
Requirements Alignment A1: 100%; A2: 100%; A3: 5/5 End-to-end functional intent preserved across chain topology. All interfaces emitted as standalone files. Dataflow matches blueprint specifications without omissions.
Architectural Quality B1: 5/5; B2: 5/5 Linear pipeline topology with correctly paired Provided/Required ports. Sender-Receiver and Client-Server interface kinds consistent with specified types.
Engineering Quality C1: 5/5; C2: 5/5; C3: minor edits Consistent naming conventions and modular file separation. No architectural changes required.
Toolchain Integration D1: success All systems import cleanly into commercial AUTOSAR configuration tooling.
Middle Tier (7 systems, 45 components)
Requirements Alignment A1: 100%; A2: 100%; A3: 3/5 Representative issue in one system: Diagnostic coordination module both provides and requires the same Client-Server interface, violating client-server separation. UDS service provided-ports absent on coordination/router components where requirements specified them.
Architectural Quality B1: 2/5; B2: 5/5 Representative issue: Client-Server role confusion in diagnostic subsystem. Local type matching preserved, but global coordination pattern violated.
Engineering Quality C1: 5/5; C2: 5/5; C3: moderate edits Representative remedy: Remove spurious provided services on diagnostic modules; add missing UDS exposure ports.
Toolchain Integration D1: success All systems import despite semantic inconsistencies.
Complex Tier (6 systems, 54 components)
Requirements Alignment A1: 100%; A2: 100%; A3: 2/5 Representative issues: SecOC chain lacks complete tag generation/verification steps. Health module architecture contradicts observer pattern—pulls business data via required-ports rather than consuming pushed health events.
Architectural Quality B1: 1/5; B2: 5/5 Representative issue: Health module pulls business data via required-ports to infer health, contradicting intended observer pattern where producers push health states via provided-ports.
Engineering Quality C1: 5/5; C2: 5/5; C3: partial rewrite Representative remedy: Refactor health monitoring to pure consumer of health events; reestablish SecOC end-to-end tag flow.
Toolchain Integration D1: success All systems import successfully. Structural validity does not guarantee semantic correctness.
Generation Cost and Scalability.
Blueprint synthesis (Phase 1) incurs moderate cost averaging 17,548 to 26,617 input tokens and 127 to 275 seconds across complexity tiers. Component assembly (Phase 2) dominates computational cost, scaling super-linearly with system complexity: Simple systems average 63,959 output tokens and 195 seconds, Middle systems rise to 184,605 tokens and 530 seconds, and Complex systems reach 280,038 tokens and 717 seconds. The approximate 1.5× token increase and 1.35× time increase from Middle to Complex tiers reflect the combinatorial explosion of cross-file reference formation and context integration requirements as dependency graph density increases.
Expert Review Findings.
Table 6 presents aggregated expert assessment across the four evaluation dimensions. For Simple systems, all dimensions score maximally: requirements alignment achieves 100% functional coverage and interface completeness with full traceability (A3: 5/5), architectural quality maintains perfect dependency consistency and type matching (B1: 5/5, B2: 5/5), engineering quality exhibits consistent naming and modular structure requiring only minor edits (C1: 5/5, C2: 5/5, C3: minor), and all systems import successfully into commercial tooling (D1: success).
Middle systems maintain perfect functional coverage and interface completeness but exhibit reduced traceability (A3: 3/5). A representative failure pattern observed in one Middle-tier system involves diagnostic coordination components: a diagnostic module incorrectly both provides and requires the same Client-Server interface, violating the client-server separation principle fundamental to AUTOSAR diagnostics architecture. Additionally, Unified Diagnostic Services (UDS) exposure is incomplete, with diagnostic provided-ports absent on coordination and router components where the requirements specified them. Architectural quality assessment reflects these issues with dependency consistency degrading to B1: 2/5 while local type matching remains correct (B2: 5/5). Engineering quality scores remain high (C1: 5/5, C2: 5/5), with moderate repair effort (C3: moderate edits) required to remove spurious provided services and add missing UDS ports. All Middle systems import successfully despite these semantic inconsistencies.
Complex systems encounter more severe semantic deviations. A representative Complex-tier system exhibits two major issues: the SecOC (Secure Onboard Communication) authentication chain lacks complete tag generation and verification steps, breaking the end-to-end security guarantee; and health monitoring components contradict the intended observer pattern by pulling business data via required-ports to infer health status, rather than consuming health state events pushed via provided-ports as specified. Requirements traceability declines to A3: 2/5, and dependency consistency reaches B1: 1/5, indicating fundamental architectural misalignment. The health monitoring pattern violation exemplifies architectural role drift: the generated component topology is locally type-correct (B2: 5/5) but violates the global observer discipline specified in system requirements. Repair effort escalates to C3: partial rewrite, requiring refactoring of health monitoring subsystems and re-threading of SecOC tag flows. Despite these issues, all Complex systems import into tooling, demonstrating that structural validity alone does not guarantee semantic correctness for system-level configurations.
4.4.3. Analysis and Discussion
4.4.4. Analysis and Discussion
Structural Scaling Success and the Blueprint Role.
The universal 100% XSD pass rates across all complexity tiers (Table 5) confirm that blueprint-guided constrained decoding effectively solves the syntax generation problem. The blueprint successfully stabilizes file organization and local type matching (B2 consistently scores 5/5). However, its effectiveness is limited to local properties: expert reviews reveal that dependency consistency degrades significantly in Complex systems (B1: 1/5). This indicates that while blueprints capture explicit interface declarations, they fail to enforce implicit global invariants such as client-server role separation or observer-style propagation.
Fundamentally, the observed 100% structural validity is not merely empirical but theoretical. The Layer-1 constrained decoding effectively functions as a deterministic model transformation that maps a subset of the XSD schema into a strictly enforced finite automaton. Since the decoding process is mathematically bound to valid transitions within this automaton, any structural error would imply a defect in the constraint compilation rules ($\kappa_{\mathrm{struct}}$) rather than a stochastic failure of the LLM. Thus, for syntactic properties, ATLAS transforms the probabilistic generation task into a controllable, correct-by-construction derivation process, leaving only semantic ambiguities to the probabilistic nature of the model.
The Gap Between Reference Existence and Semantic Correctness.
The discrepancy between high reference resolution rates (>>92.8%) and the 0% SMT pass rate highlights a critical hallucination mode: the LLM minimizes “syntax loss” (generating valid-looking IDs) but ignores “logic loss” (semantic coherence). For instance, a SenderPort might syntactically reference a ClientServerInterface, satisfying XML schema rules but violating domain logic. This confirms that reference resolution is a necessary but insufficient proxy for correctness, necessitating the post-hoc L2 verification layer.
Graduated Automation: Design Philosophy.
The observed failure modes validate the “graduated automation” design philosophy. Rather than striving for theoretical 100% zero-shot automation—which proves brittle in complex semantic environments—ATLAS adopts a strategic division of labor. The framework reliably automates the 80% of routine structural construction (validated by 100% XSD compliance), while identifying and surfacing the remaining 20% of complex semantic ambiguities (validated by L2 diagnostics) for targeted repair. The one-iteration convergence of AGR for local issues (SHACL) and the fallback to the Configuration Editor (Route B) for global conflicts (SMT) demonstrate a practical workflow that aligns with industrial safety standards, where human oversight is mandatory but optimized.
Complexity-Dependent Failure Progression and Manual Repair.
The tier-specific expert review findings reveal that failure modes intensify with system complexity. Middle systems encounter isolated role confusion (e.g., client-server violations) that are amenable to automated repair. However, Complex systems exhibit cascading semantic violations—such as broken SecOC authentication chains or inverted health monitoring patterns—that span multiple subsystems. Expert analysis confirms that while automated repair mechanisms (Route A) efficiently resolve local constraint violations, these systemic architectural incoherencies often require human-in-the-loop intervention (Route B). For instance, re-threading a security chain requires high-level architectural intent that current LLMs struggle to infer from local constraints alone. This validates ATLAS’s graduated automation model: the system produces a structurally sound scaffolding, but relies on domain experts to perform the final ”last-mile” semantic repair via the MDE editor, shifting their role from ground-up authoring to high-value architectural correction.
Engineering Implications and Continuous Optimization.
The observed failure modes inform practical deployment strategies. It is important to note that our current experimental setup excluded explicit constraint descriptions from the RAG retrieval context due to the prohibitive volume of constraints linked to individual files. However, ATLAS is designed as a continuous integration system capable of long-term optimization.
First, audit-driven prompt enhancement can be deployed: by analyzing the Evidence Registry to identify frequently violated constraints, the system can selectively inject high-risk constraints into the generation prompt based on file complexity. This dynamic adjustment—analogous to verbal reinforcement learning loops (shinn2023reflexion)—balances constraint adherence against the risks of hallucination and attention drift caused by excessive context, a trade-off that requires long-term iterative tuning (chen2024selfdebug).
Second, learning from golden data offers a pathway to reduce common errors. As the system accumulates verified artifacts (golden data) from successful repairs, fine-tuning the base model can internalize domain invariants that are currently missed. Recent findings suggest that even a small set of high-quality examples can significantly align model behavior with complex domain rules (zhou2023lima).
Nevertheless, our results suggest that even with these enhancements, achieving zero-shot, perfect generation of complex multi-file systems that satisfy all global constraints remains infeasible. The inherent conflict between local generation contexts and global system invariants ensures that the verify-repair loop—and specifically the integration of human expertise for complex ambiguity resolution—will remain an essential architectural component for safety-critical generation.
Model Capabilities and API-Based Generation.
The structural correctness achieved through API-based generation with JSON Schema constraint enforcement provides guarantees comparable to vLLM-based constrained decoding demonstrated in RQ1, despite lacking token-level audit trails. However, the cross-file semantic failures observed in Middle and Complex tiers indicate that stronger base models, while improving blueprint quality and component coherence, do not alone solve the global consistency problem. Even frontier models operating under structural constraints struggle with system-level reasoning when context limitations prevent full-system visibility during component generation.
This finding implies that addressing the multi-file scalability boundary requires architectural innovations—enhanced blueprint specifications, incremental validation, hybrid human-machine workflows—rather than solely relying on improved base model capabilities. While stronger models may reduce the frequency or severity of semantic violations, the fundamental challenge of maintaining global invariants under local generation constraints persists across model families.
Graduated Automation: Design Philosophy and Validation.
The observed interaction between structural success and semantic refinement empirically validates the “graduated automation” model. Rather than striving for a theoretical 100% zero-shot automation—which proves brittle in complex semantic environments—ATLAS adopts a strategic division of labor. These results confirm the design goal stated in Section 3: the system successfully automates the routine 80% of configuration tasks (as seen in the structural completeness metrics), allowing engineers to focus their expertise on the remaining 20% of complex semantic disambiguation (as captured in the refinement effort). This split aligns with industrial safety standards, where human oversight is mandatory; ATLAS optimizes this oversight by converting it from manual authoring to high-level evidence review.
4.5. RQ3: Cross-Domain Transferability via UMM S2 Path
We evaluate the framework in a domain without an explicit meta-model: cross-border civil and commercial jurisdiction under the Brussels I bis Regulation. This setting probes whether the UMM S2 path can induce a usable meta-model and constraint graph directly from statutes and cases, and whether the ATLAS architecture validated in AUTOSAR can be transferred into a high-stakes legal reasoning setting that demands doctrinal correctness and faithful citation rather than domain-specific engineering artifacts.
4.5.1. Experimental Setup
We examine a legal domain without a formal schema (Brussels I bis). Prior work notes domain hallucination risks (Magesh2025); we therefore use the UMM S2 path to induce entities/relations/constraints from text and test whether layered enforcement reduces doctrinal errors (Magesh2025).
UMM S2 Pipeline Architecture.
The evaluation employs the S2 pathway described in Section 3, comprising four stages:
  1. (1)
    LLM-induced meta-model instantiation: Given a fact pattern describing a cross-border dispute, the system induces a UMM mini-instance capturing entities (parties, contracts, properties), roles (plaintiff, defendant, contracting parties), and connecting factors (domicile, contract performance location, property situs) together with their applicability preconditions.
  2. (2)
    Provision linking (ICM construction): The induced factors are mapped to candidate Brussels I bis provisions. For example, immovable property location links to Article 24(1) (exclusive jurisdiction), contract performance location links to Article 7(1)(b) (special jurisdiction), and a choice-of-court clause links to Article 25 (prorogation), optionally paired with Article 31 in parallel-proceedings / first-seised scenarios.
  3. (3)
    Promote-aware constrained decoding (Layer-1): JSON Schema-constrained generation enforces a fixed priority hierarchy when selecting a jurisdictional basis: exclusive jurisdiction (Article 24) outranks agreement-based jurisdiction (Articles 25 and 31), which outranks lis pendens stays (Article 31 alone), which outranks special jurisdiction (Article 7), which in turn outranks general jurisdiction (Article 4) and appearance-based jurisdiction (Article 26). The system commits to one promoted basis before emitting natural-language justification.
  4. (4)
    Post-hoc verification and repair (Layers 2–3): Deterministic auditing checks legal constraints: Article 24 overrides lower bases; appearance (Article 26) cannot displace exclusive jurisdiction; Article 25 invocation requires proper formality and is paired with Article 31 analysis in parallel proceedings; and missing formal prerequisites (e.g., insufficiently supported choice-of-court agreement) force abstention instead of overconfident assignment.
The terminal output is a Decision Package containing (i) an Evidence Registry, which records applicable provisions, connecting factors, and any missing prerequisites, and (ii) a Structured Decision JSON, which encodes the jurisdictional conclusion, the selected forum type, and cited provisions in machine-auditable form.
Figure 10. UMM S2 pipeline architecture for cross-domain Private International Law evaluation. The pipeline instantiates a UMM mini-instance from fact patterns, links induced factors to Brussels I bis provisions through ICM construction, applies promote-aware constrained decoding with JSON Schema enforcement, and performs multi-layer verification with deterministic repair. Output comprises Evidence Registry and Structured Decision JSON forming a complete Decision Package.
Architecture of the UMM S2 pipeline for evaluating cross-domain Private International Law scenarios. Fact patterns are used to instantiate a UMM mini-instance. Induced factors are linked to Brussels I bis provisions via Integrated Constraint Model (ICM) construction. The system then applies promote-aware constrained decoding with JSON Schema enforcement, followed by multi-layer verification and deterministic repair. The final outputs are an Evidence Registry and a Structured Decision JSON, which together form a complete Decision Package.
Comparative Configurations.
We compare three configurations using the same API-served model to isolate architectural effects:
  • Baseline (LLM-only): Single-pass prompting without retrieval, schema constraints, or audit. The model receives only the fact pattern and produces free-form jurisdiction analysis.
  • RAG: Retrieval-augmented prompting in the style of commercial legal assistants studied in (Magesh2025). The model receives the fact pattern plus top-$k$ Brussels I bis provisions retrieved for that scenario, but generation remains unconstrained: it may assert a jurisdictional basis even if statutory preconditions are not actually satisfied.
  • ATLAS: Full UMM $\rightarrow$ ICM $\rightarrow$ CVG S2 pipeline. The system first induces the structured mini-instance (UMM), links factors to provisions (ICM), then performs promote-aware JSON Schema-constrained decoding followed by deterministic verification and repair. This configuration enforces precedence, exclusivity, pairing requirements, and formality gates before finalizing the decision.
Dataset and Evaluation Metrics.
The evaluation uses 60 expert-authored Brussels I bis jurisdiction cases spanning the full hierarchy of bases of jurisdiction: exclusive (Article 24), agreement-based (Articles 25 and 31), lis pendens / first-seised stay (Article 31), special jurisdiction (Article 7), general jurisdiction (Article 4), and appearance (Article 26). Each case was drafted and doctrinally validated by a subject-matter expert in European civil procedure. As noted above, this dataset is intended to probe architectural transfer, not to claim full coverage of real-world PIL litigation.
[…]
4.5.2. Results
Table 7 summarizes results across the three configurations. ATLAS improves correctness, citation faithfulness, structural compliance, and priority alignment relative to both Baseline and RAG. RAG improves over the Baseline in several dimensions but still fails to guarantee constraint compliance.
Table 7. Cross-domain evaluation on 60 Brussels I bis jurisdiction cases comparing Baseline (LLM-only), RAG retrieval augmentation, and the full ATLAS pipeline.
Metric Baseline RAG ATLAS
Legal Correctness 0.200 0.300 0.467
Citation Precision 0.250 0.428 0.778
Abstention Quality 0.500 0.800 0.800
Promotion Accuracy 0.367 0.567 0.933
Schema Compliance 1.000
Rule Satisfaction 1.000
Legal Correctness and Citation Quality.
[…]
Constraint Compliance, Abstention, and Priority Alignment.
ATLAS achieves perfect Schema-Pass (1.000) and Rule-OK (1.000), meaning every final output is syntactically valid and satisfies precedence, exclusivity, and pairing constraints. Promotion Accuracy (alignment between the claimed forum type and the strongest legally available basis) increases from 0.367 (Baseline) to 0.567 (RAG) and 0.933 (ATLAS), indicating that ATLAS’s promote-aware decoding and deterministic repair suppress “everything applies” answers and force adherence to the Brussels I bis priority ladder.
Abstention Quality—the ability to defer judgment when required information is missing—is 0.500 for Baseline, 0.800 for RAG, and 0.800 for ATLAS. RAG benefits here because retrieved provisions explicitly enumerate preconditions; the model can recognize that those preconditions are not satisfied and refuse to commit. ATLAS preserves that epistemic caution while additionally enforcing that abstention is required when, for example, Article 25 formality prerequisites are underspecified.
Computational Cost.
ATLAS incurs higher cost due to multi-stage generation, schema-constrained decoding, and post-hoc verification. This overhead is the direct cost of guaranteeing structured legality.
4.5.3. Analysis and Discussion
Architectural Transferability.
These results show that the UMM–ICM–CVG architecture transfers to a non-engineering, text-governed domain (PIL jurisdiction) via the S2 inductive pathway. Without any pre-existing metamodel, the system can induce a structured mini-instance, align facts to candidate provisions, enforce hierarchical precedence, and output auditable decisions. ATLAS delivers higher Legal-Correct@1 than both Baseline and RAG (0.467 vs. 0.200 / 0.300), far higher Promotion Accuracy, and perfect Schema-Pass / Rule-OK. This indicates that layered constraint enforcement meaningfully outperforms unconstrained prompting and improves upon retrieval augmentation alone (Magesh2025).
Discussion.
RAG improves correctness (0.300 vs. 0.200 Baseline) and abstention (0.800 vs. 0.500), echoing (Magesh2025), but does not enforce doctrinal validity. ATLAS’s UMM $\rightarrow$ ICM $\rightarrow$ CVG pipeline adds precedence rules, exclusivity constraints, and formality gates, yielding 0.933 Promotion Accuracy, 0.778 Citation Precision, and perfect structural legality (Schema-Pass = Rule-OK = 1.000). Each case logs UMM instance, factor-provision mappings, decoding trace, and repair actions, providing end-to-end provenance for practitioner review.
This 60-case Brussels I bis evaluation is a feasibility probe testing architectural transfer, not claiming exhaustive legal coverage or deployment readiness. Scaling requires broader doctrinal coverage, multi-jurisdictional support, and human oversight in professional workflows.
5. Conclusion and Future Work
This paper presented ATLAS, a high-assurance generation framework integrating LLMs with model-driven engineering and formal methods to produce verifiable, auditable artifacts in safety- and regulation-critical domains. ATLAS combines a Unified Meta-Model (UMM) for typed domain semantics, an Integrated Constraint Model (ICM) for constraint aggregation, and Constraint-guided Verifiable Generation (CVG) with Audit-Guided Repair (AGR) to produce verifiable artifacts $\mathfrak{A}=(a,\Pi,\varphi)$. Constraints are stratified across two layers: L1 enforces structural rules during decoding (Theorem 10), while L2 validates semantic/logic constraints post-generation with machine-checkable certificates.
Positioning.
[…]
Limitations.
  • Incremental semantic/logic validation. Our current L2 semantic and logic validators (instantiated as SHACL and SMT) operate on completed artifacts. While L1 is prefix-decidable and enforces $\mathcal{C}_{\mathrm{struct}}$ online, incremental L2 checking of $\mathcal{C}_{\mathrm{sem}}$ and $\mathcal{C}_{\mathrm{log}}$ during long generation (e.g., multi-file systems) remains future work.
  • System-level semantics at scale. In RQ2, architectural drift and dangling cross-file references appear in more complex systems. ATLAS can surface these violations and drive AGR, but does not yet guarantee global semantic invariants for arbitrarily dense dependency graphs.
  • From formal to programmatic scale. While SHACL and SMT provide rigorous proofs for safety-critical properties, scaling them to thousands of routine business rules introduces significant maintenance overhead and performance bottlenecks. Future industrial implementations of ATLAS may adopt a hybrid validation strategy: employing SMT only for critical safety invariants (e.g., deadlock freedom, timing bounds) while utilizing LLMs to synthesize highly optimized programmatic validators (e.g., Java/C++ assertions) from the ICM. By applying symbol-constrained decoding over the Path S1 APIs, this approach will combine the flexibility of LLM code generation with the type-safety of MDE, resolving the scalability challenges of pure formal verification.
  • Human review remains essential. ATLAS narrows the surface that humans must review and provides $\Pi$ and $\tau$ to justify remaining issues, but it does not claim unsupervised certification. Certified deployment still requires domain experts and regulatory sign-off.
Future Work.
We outline several directions for advancing ATLAS and the broader UMM–ICM–CVG paradigm:
  1. 1)
    Incremental semantic and logic validation. Beyond the post-artifact L2 validation described in Limitations, we aim to integrate streaming SHACL engines (yeSatLMSatisfiabilityAidedLanguage2023) and incremental SMT solvers accepting partial formulas, enabling Layer-2 feedback during generation. A key challenge is exposing partial, request-scoped constraints without exponential solver call blow-up.
  2. 2)
    Fine-grained constraint strategies. Our current L1 execution uses a unified automaton that encodes JSON Schema / Regex / FSM fragments and bounded-unfolded GBNF. Future variants will mix enforcement modes at block granularity: for critical subtrees we can require fully deterministic automata and one-step closure; for descriptive regions we can allow looser local grammars. This opens the door to adaptive enforcement policies that trade completeness of $\pi_{\text{struct}}$ against linguistic naturalness, guided by live feedback from AGR.
  3. 3)
    Audit-guided fine-tuning and distribution-aware decoding. Although ATLAS is model-agnostic and achieved the reported results without domain-specific fine-tuning, the framework exposes systematic opportunities to leverage audit trails for targeted model improvement. The runtime audit logs $\tau$ and AGR repair traces capture precise failure modes and correction patterns that reflect domain-specific constraint violations. These audit-derived signals can guide efficient fine-tuning in vertical domains by identifying high-value constraint patterns where model alignment would reduce repair overhead. Pre-training or fine-tuning on (UMM, ICM, $\Pi$)-annotated corpora could bias $p_{\text{LLM}}$ toward emitting structures that already satisfy $\mathcal{C}_{\mathrm{struct}}$ and respect high-value semantic invariants, reducing AGR workload and improving first-pass success rates, particularly for the $\sim$20% of architecturally dense cases that currently require iterative repair.
    A critical consideration is that grammar-constrained decoding (GCD) may induce distribution distortion by aggressively masking tokens to enforce structural constraints. Recent grammar-aligned decoding (GAD) (parkGrammaralignedDecoding2025) methods aim to respect grammar constraints while preserving the base model’s token distribution. Future work should explore hybrid approaches that combine audit-guided fine-tuning with distribution-aware sampling techniques. Selectively applying stricter GCD enforcement only to safety-critical structural regions while employing GAD-style preservation in descriptive regions could balance formal correctness guarantees with linguistic naturalness. The audit trail $\tau$ provides direct empirical evidence of where enforcement overhead is justified versus where relaxed sampling suffices, enabling data-driven calibration of this enforcement/naturalness trade-off.
    Practitioners must balance infrastructure cost (e.g., vLLM with full per-token auditing) against fine-tuning investment. A systematic decision framework is needed: for a given domain and deployment context, should one invest in targeted fine-tuning of $p_{\text{LLM}}$, or rely on stronger automaton enforcement plus richer KG/RAG retrieval, or adopt a hybrid? This trade-off depends on artifact volume, update frequency of $\mathcal{C}_{\mathrm{sem}}$$\mathcal{C}_{\mathrm{log}}$, and the acceptable amount of human-in-the-loop review for residual hard cases. The layered constraint architecture (UMM–ICM–CVG) is orthogonal to and synergistic with fine-tuning: fine-tuning attempts to align $p_{\text{LLM}}$ with $\mathcal{C}$ before decoding, while CVG guarantees prefix safety, bounded closure, and machine-checkable $\Pi$ during and after decoding. Quantifying this synergy and automating the enforcement/tuning mix is an important next step.
  4. 4)
    Domain applicability and scenario recommendations. The UMM–ICM–CVG architecture exhibits varying degrees of suitability across different LLM application scenarios. For transformations from natural language or multimodal inputs to structured artifacts—the focus of this work—ATLAS provides formal correctness guarantees and audit trails that address the trustworthiness requirements of regulated domains. The framework is particularly well-suited to domains characterized by explicit meta-models (or meta-models that can be inductively extracted from specifications), comprehensive constraint specifications spanning structural, semantic, and logical dimensions, and validation requirements that demand machine-checkable evidence of conformance.
    Conversely, direct structured-to-structured transformations that demand perfect consistency may not benefit from LLM-based generation, as deterministic rule-based transformations typically suffice and avoid probabilistic uncertainty. However, structured-to-natural-language generation represents a promising application domain where UMM-based abstraction can strengthen LLM comprehension of source structures and reduce semantic drift. By materializing persistent UMM representations of structured data sources, the framework enables LLM-based analysis and explanation tasks to retrieve accurate semantic metadata and constraint context via RAG mechanisms. This UMM-guided comprehension pathway prevents the model from hallucinating incorrect structural interpretations and ensures that natural language outputs remain grounded in verified domain semantics. Such scenarios include generating documentation from AUTOSAR configurations, producing compliance reports from legal decision structures, or synthesizing explanatory narratives from complex engineering artifacts where the UMM serves as a semantic anchor preventing drift between the structured source and natural language output.
  5. 5)
    Hardware-aware embedded code generation. Beyond software architecture (AUTOSAR), ATLAS is intrinsically suited for low-level embedded firmware generation. The UMM can be instantiated from hardware description standards (e.g., IP-XACT, CMSIS-SVD) to model chip peripherals and registers. The ICM can capture hardware invariants often ignored by general LLMs, such as read/write access policies, atomic bit-field manipulations, and strict timing constraints. Applying CVG in this domain could produce ”correct-by-construction” device drivers that are mathematically guaranteed to respect the hardware specification, significantly reducing the risk of bricking devices or causing physical safety hazards.
Take-away.
ATLAS operationalizes evidence-carrying artifact generation for regulated domains through layered constraint enforcement. Our evaluation on industrial AUTOSAR systems empirically validates the 80/20 graduated automation principle: By extending the deterministic model transformation principles of S2D2 (DualStageFramework) into the realm of generative AI, ATLAS reliably offloads the bulk of routine structural construction to LLMs, while elevating human experts from manual authors to audit-enabled reviewers for the critical “last mile” of semantic refinement. By coupling the flexibility of LLMs with the accountability of MDE, ATLAS offers a practical path toward trustworthy AI-assisted engineering where complete automation is neither feasible nor desirable.
Acknowledgements.
We thank Xinyue Yang (School of Law, Wuhan University) for helpful discussions on Brussels I bis, legal analysis regarding exclusive jurisdiction, choice-of-court formalities, and for reviewing preliminary examples and gold labels. All remaining errors are our own. This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant 32427801 (National Major Scientific Research Instrument Development Project). We thank the anonymous reviewers for their valuable feedback.